### v2

# Learning not cheating: AI assistance can enhance rather than hinder skill development

Abstract
It is widely believed that outsourcing cognitive work to AI boosts immediate productivity at the expense of long-term human capital development. An overlooked possibility is that AI tools can support skill development by providing just-in-time, high-quality, personalized examples. In this investigation, lay forecasters predicted that practicing writing cover letters with an AI tool would impair learning compared to practicing writing letters without the tool. However, in a highly-powered pre-registered experiment, participants randomly assigned to practice writing with AI improved more on a writing test one day later compared to writers assigned to practice without AI. Notably, writers given access to the AI tool improved more despite exerting less effort, whether measured by time on task, keystrokes, or subjective ratings. We replicated and extended these results in a second pre-registered experiment, showing that writers given access to the AI tool again outperformed those who practiced on their own—but performed no better than writers merely shown an AI-generated cover letter that they could not edit. Collectively, these findings constitute an existence proof that by providing personalized examples of high-quality work, AI tools can improve, rather than undermine, learning.
keywords:
[…]
Generative AI (henceforth AI) tools are increasingly powerful and prevalent (1), and there is mounting evidence that they can dramatically boost performance. For example, working side-by-side with AI as a copilot has been shown to increase both quality and speed in a variety of professional writing tasks (e.g., emails, memos, short reports) (2, 3, 4).
Nevertheless, there is growing concern that AI tools will be used as a crutch, providing immediate gains in performance at the expense of long-run development of human capital (5, 6, 7). For instance, in a 2024 poll, 62% of surveyed adults predicted that Generative AI will “lead to humans becoming less intelligent” (8). In January 2023, New York City public schools banned ChatGPT, citing “concerns about negative impacts on student learning (9).” When this ban was lifted three months later, it was not because of the potential of AI to scaffold learning, but instead because of the “reality that students are participating in and will work in a world where understanding Generative AI is crucial (10).” The sentiment behind the initial ban aligns with teacher perceptions: in a nationally representative poll in May 2024, four times as many K-12 educators judged the use of AI tools as net harmful (24%) than net beneficial (6%) (11).
Concerns that using AI tools hinders learning (while increasing short-term performance) are justified for at least three reasons. First, AI systems based on large language models like GPT-4 have been shown to confidently assert erroneous facts (i.e., hallucinations (12)), make reasoning and arithmetic errors (13), and complete other tasks with varying degrees of accuracy.
Second, regardless of accuracy, the fluent and instantaneous solutions AI tools generate may contribute to an illusion of mastery. To the extent users conflate the skills of an AI tool with their own, they may be less likely to seek feedback and improve. Prior research has found that searching for information on the Internet, for example, creates an illusion whereby people conflate knowledge outside their heads with what they personally know themselves (14).
Third, technological tools reduce the need for the learner to be cognitively engaged with the task at hand. For instance, knowing that we will be able to search for a fact on a computer has been shown to reduce memory for that fact, instead encouraging recall of how to search for it (15). And drivers who use GPS tend to have worse hippocampal-dependent spatial memory, both cross-sectionally and longitudinally (16, 17). To the extent that tools powered by generative AI instantaneously produce turn-key solutions for complex cognitive tasks, they may be especially detrimental to learning. It is, after all, tempting to copy and paste the output of an AI tool without even laying eyes on it.
There is a plausible, albeit less obvious, alternate hypothesis. Using AI tools to help do our work could help us develop our own skills. In particular, the current generation of AI tools may teach by example, offering high-quality and personally tailored demonstrations of abstract principles that are otherwise difficult to grasp and apply. Classic research shows that worked examples of math problems (i.e., not just answers but the step-by-step process by which problems are solved) scaffold learning more effectively than explanations alone (18, 19). Compared to textbooks, conventional computer tutoring programs (e.g., (20)), and even human teachers, today’s AI tools are unprecedented in their ability to deliver hyper-personalized high-quality examples on command. Thus, AI tools may improve skill development if the benefits of exposure to just-in-time examples tailored to learners’ needs outweigh the costs of diminished engagement.
There is little research on how AI tools influence skill development. In working papers, results have been mixed. Some studies have found that interacting with AI tools improves skill on subsequent tests in which AI tools are not available (21, 22), while others have shown null or even negative effects (23, 24, 22). Notably, these studies examine AI tutors, chatbots, or explanations explicitly designed to support learning, rather than simply providing solutions as is typical in real-world use. Further, they focus exclusively on mathematics and computer programming.
In this investigation, we ask whether AI tools can increase intermediate-term skill development, above and beyond just improving performance while using the AI tool. We focus on writing—the most common use of AI at work, as ranked in a nationally representative survey of American adults in August 2024 (25). Participants in our three pre-registered studies were American adults on the survey platform Prolific.
To differentiate the effects of AI use on learning versus performance, we developed a paradigm in which all participants were given a baseline writing test (i.e., revising a poorly written cover letter). They were then introduced to evidence-based strategies for professional writing, with descriptions and examples for each one (26, 27). Next, participants were randomly assigned to one of three conditions: practicing rewriting a different cover letter with access to an AI tool based on the same writing principles that users had just learned (See Figure 1), practicing without this tool, or a no-practice control group. To assess gains in writing skill, participants completed an incentivized test in which they rewrote yet another cover letter without access to AI, with a cash bonus guaranteed for submissions ranked in the top 10 percent. To assess the persistence of skill improvement, all participants were invited to complete a similar incentivized test of writing skill one day later. See Figure 2. We used GPT4o to rate each cover letter for each of the five writing principles introduced in this experiment. We averaged these ratings to produce summary scores of writing skill and, in a random subsample (n = 30), validated these scores using trained human raters (r = .83, p < .001). Additionally, we asked a separate sample of participants to read pairs of test-phase cover letters randomly selected from different conditions, and to indicate which letter would be more likely to secure a job interview. The relative likelihood of a cover letter securing an interview correlated positively with AI ratings of writing skill (rs = .29 and .28, ps < .001, for Study 2 and 3, respectively).
In Study 1, forecasters presented with this design were twice as likely to predict that practicing writing with the assistance of the AI tool would impair learning compared to practicing without the AI tool. In Study 2, however, participants who had practiced with the AI tool learned more (i.e., wrote better cover letters during the test phase) compared to either comparison group—an advantage that persisted in a one-day follow-up test. Finally, in Study 3, we explored the mechanism for these learning gains by introducing an example-only condition. Participants who had merely seen an AI-generated example (but did not have an opportunity to practice) improved in writing skill as much as participants who had practiced with an AI tool; benefits again persisted in a one-day follow-up test. Test-phase cover letters written by participants who had practiced with AI (Studies 2 and 3) or had seen an AI example (Study 3) were more likely to secure hypothetical job interviews compared to cover letters written by participants who had not practiced (Study 2) or had practiced without AI (Studies 2 and 3).
Study 1
Lay forecasters predicted that practicing with AI would hinder learning
We showed N = 150 participants screenshots of a random assignment study with three conditions. We asked them to rank-order these conditions according to how much they predicted future participants would learn in each. Confirming our pre-registered hypothesis, nearly twice as many forecasters (64.7%) ranked practicing alone above practicing writing with access to an AI tool as the converse (35.3%, (1) = 12.9, p < .001), see Figure 3. Participants made this prediction regardless of self-reported experience with AI (OR = 0.83, p = .239) or any other measured demographic characteristic (ps > .05).
In open-ended responses, forecasters who were pessimistic about the effect of the AI tool on learning speculated that it would crowd out effort (e.g., “Practicing alone would force more recall and problem-solving skills, while AI essentially gives the answer for them.”, “I think oftentimes using AI impedes the learning process because it’s the ‘easy way.”’). Those with positive views, on the other hand, cited the possibility of AI providing insights or examples that would be otherwise unavailable (“As much as I hate AI, I do not believe you can improve in any manner if you do not have examples or ways of learning, and AI can provide this.”)
Study 2
Study 2 tested whether the predictions of Study 1 forecasters were accurate. Specifically, N = 2,238 participants completed a baseline questionnaire and pretest (rewriting a poorly written cover letter), followed by a lesson introducing five principles of effective writing (i.e., less is more, make reading easy, design for easy navigation, use enough formatting but no more, make responding easy) (26). Next, participants were randomly assigned to one of three practice conditions: (1) rewriting a new cover letter with an AI writing tool that revised text instantly based on these principles, (2) rewriting the new cover letter without the AI tool, or (3) a no practice control. At the end of the session, all participants completed a test of writing skill (rewriting a yet another cover letter without access to the AI writing tool) and an exit questionnaire. One day later, all participants were invited to complete a similar incentivized test of writing skill.
AI practice improved writing skill
Consistent with other studies demonstrating the productivity benefits of AI tools (3, 2), participants given access to the AI writing tool produced cover letters during the practice phase that were dramatically higher in quality than participants who were not (d = 1.01, p <.001).
The learning advantage of having practiced with AI was evident in the test phase: Consistent with our pre-registered hypothesis, participants who had practiced with the AI tool produced higher-quality writing than did participants who either had practiced without the AI tool (d = .38, p <.001) or who had not practiced at all (d = .47, p <.001). See Figure S4. Likewise, cover letters written by participants who had practiced with AI were more likely to secure a hypothetical job interview than cover letters by participants who had practiced without AI (.54 vs. .50, p = .002) or had not practiced at all (.54 vs. .47, p < .001). See Figure 5. Notably, the effect of practicing with AI was not limited to stronger or weaker writers—participants across all skill levels benefited equally from practicing with AI.
AI practice was less effortful
The learning benefits of using an AI writing tool were evident despite reduced effort during the practice phase. Compared to participants who practiced alone, participants who practiced with the AI tool spent 0.44 fewer minutes during the practice phase (3.73 vs. 4.17; d = -.12, p = .025), logged roughly a quarter as many keystrokes (26% d = -.44, p <.001) and self-reported expending less effort during practice (d = -.31, p <.001).
Nevertheless, it would be inaccurate to label writers practicing with AI as entirely disengaged. Copying, pasting, and submitting the AI tool’s output could be accomplished almost instantly. Yet, the majority of participants chose to interact with the task for over 3 minutes, and 95% made at least one edit to the AI-generated output before final submission. See Figure S5 in Supplementary Information for details.
Differential effort during practice raised the possibility that participants who had practiced with the AI tool outperformed those who had practiced on their own because they were less fatigued during the test phase. However, participants who had practiced with AI did not spend more time as those who had practiced without AI (d = .06, p = .235), but logged more keystrokes (d = .12, p = .026), and self-reported expending less effort (d = -.12, p = .023) during the test phase.
AI practice did not create the illusion of mastery
Following the test phase, there were no differences by condition on self-reported knowledge or motivation to improve. Despite improving more in objectively assessed writing skill, participants who had practiced with AI reported having learned no more than those who had either practiced alone or done no practice at all (ps >.05). Self-ratings of writing skill after the practice phase were also indistinguishable between participants who practiced with AI and those who practiced without the AI tool, or in the no-practice control group (ps >.05), but participants who had practiced without AI rated their skill more highly than those who did not practice (d = .14 p = .008). Finally, compared to participants who did not practice, participants who had practiced with AI were slightly less likely to request feedback after the test phase (.65 vs. .60, p = .039), but just as likely as participants who had practiced without AI (.64 vs. .60, p = .167). See Section B4 in the Supplementary Information for details.
The effects of practicing with AI persist
To examine whether the treatment effects persisted over time, we re-contacted all participants one day later. The majority of participants responded (87%), and attrition rates did not differ by condition (13% to 14% = .68, p .710). Confirming our pre-registered hypothesis, participants who had practiced with the AI tool to practice the previous day continued to outperform those who had practiced without the tool (d = .41, p <.001) as well as those who had not practiced at all (d = .46, p <.001). Participants who had practiced without the AI tool performed no better than those who did not practice (d = .05, p = .331). See Figure 4 and Section B5 of Supplementary Information for details.
None of the findings above were moderated by individual difference variables, including past experience with AI, age, gender, race, education, motivation to learn, and baseline writing skill, BH-corrected p-values >.05. See Table S12 in the Supplementary Information for details.
Study 3
To gain insight into what might be driving the benefit of practicing with AI, in Study 3 (N = 2,003), we preregistered a replication and extension in which we replaced the no-practice condition of Study 2 with an example-only condition. In this condition, participants were shown an AI-generated writing example that they could not edit. To the extent that the benefit of practicing with AI was driven by exposure to examples, the example-only condition should improve performance in the test phase as much as the practice with AI condition.
Seeing an AI example was as effective as practicing with AI
As in Study 2, participants given access to the AI writing tool dramatically outperformed participants who did not get access to it, both while using it during the practice phase (d = 1.22, p <.001), and during the no-AI test phase (d = .34, p < .001). Their test-phase cover letters were also relatively more likely to secure them hypothetical job interviews (.51 vs. .47, p = .008).
Participants who had merely seen an AI-generated example also improved more in writing skill than those who had practiced without AI (d = .37, p = < .001), and produced letters that were relatively more likely to secure them interviews (.52 vs. .47, p = .007). Notably, they improved as much as participants who had practiced with the AI tool (and could edit its output, d = .03, p = .883), and were offered hypothetical interviews at similar rates as them (.51 vs. .52, p = .561). See Figures S6 and 7. As in Study 2, gains from practicing with AI and seeing AI examples were comparable across all levels of baseline writing skill.
Seeing an AI example was even less effortful than practicing with AI
During the practice phase, participants who saw the AI example spent 2.32 fewer minutes than participants practicing with AI (d = 0.99, p < .001) and 2.96 fewer minutes than participants practicing without AI (d = 1.13 p = < .001), and reported expending less effort than those practicing with an AI tool (d = 0.19, p = .003) and those practicing without AI (d = 0.32, p < .001). As in Study 2, participants who practiced with AI logged 2.83 times fewer keystrokes than did participants who practiced without AI. As expected, participants exposed to the AI example logged 0 keystrokes.
During the test phase, participants who had seen the AI example worked for an additional 56 seconds more than participants who had practiced without AI (d = 0.22, p < .001). They also logged 30% more keystrokes than participants who had practiced with AI (d = 1.86 p <.001) and 60% more than those who practiced without AI (d = 2.39, p < .001). Across conditions, all participants self-reported similar levels of subjective effort (ds < .08, ps > .05).
Seeing an AI example did not create the illusion of mastery
As in Study 2, despite learning more, participants who had practiced with AI or had merely seen an AI-generated example reported learning similar amounts to those who practiced without AI (ps > .05) and rated their writing skill after practice at comparable levels (ps > .05). Moreover, all participants requested feedback at similar rates (proportions ranged from 63% to 67%). See Section C4 in the Supplementary Information for details.
The effects of seeing an AI example persist
When we recontacted a subsample of (n = 800) participants one day later, the majority responded (n = 633, 80%); the attrition rates ranged from 17% to 24% and did not differ by condition ( (2) = 4.56, p = .102). The effect remained robust after 24 hours. Participants who had practiced with the AI-tool (d = .29, p = .006) and participants who had simply seen an AI example (d = .32, p = .003), both continued to outperform those who had practiced without the tool. Participants who had merely seen an AI example performed as well as those who had practiced with AI (d = .02, p = .830). See Figure 6 and Section C5 of Supplementary Information for details.
As in Study 2, the above findings were not moderated by individual differences. See Table S23 in the Supplementary Information for details.
Discussion
Contrary to the expectations of lay forecasters (Study 1), participants who practiced writing cover letters with an AI tool learned more than those who practiced on their own (Studies 2 and 3). Specifically, participants who had practiced with AI wrote (without AI assistance) cover letters that were rated higher in writing quality and were more likely to secure a hypothetical interview—both immediately after practice and one day later. Learning gains were not the result of greater effort; in fact, participants who’d practiced with AI expended less effort during practice than those who’d practiced alone. Instead, these gains can be explained by exposure to a high-quality, just-in-time personalized example: participants who merely viewed an AI-generated example cover letter (without editing it) improved their writing skill as much as those given the option to practice editing the cover letter (Study 3).
While not obvious to forecasters, teachers (11), and the general population (8), the benefits of seeing AI-generated examples are consistent with prior research on how people learn. In addition to the literature on worked examples mentioned earlier, research has shown that humans are especially adept at observing, imitating, and learning from others (28, 29, 30). Our findings also align with the expert performance literature: the most successful learners engage in deliberate practice, which (in addition to concentration, feedback, and repetition) depends upon detailed mental representations of excellent performance (31).
Future Directions
Three promising directions for future research are worth highlighting. First, it remains to be seen whether learning from AI examples occurs in domains other than writing, currently the modal use case for AI tools in the workplace (25) . Writing may be particularly suited to learning by example because at a glance, a single AI-generated example visually communicates the elements of effective professional writing (e.g., short sentences and boldface formatting). In other domains, merely observing a solution may be less informative. For instance, the final answer to a math or computer programming problem does not instantly reveal the procedure that produced it. Studies of AI tutors in these domains have, perhaps not surprisingly, found null or negative effects (23, 24).
Second, additional research is needed to explore moderators of learning from AI tools. Certain strategies for interacting with AI may bolster their effectiveness. For example, experimental research suggests that learners benefit more from AI explanations for math problems if they first try to solve them on their own (21). Similarly, correlational evidence that asking AI for explanations as opposed to answers is associated with more learning in mathematics (23) and computer programming (22). On the other hand, other factors could minimize learning benefits. In our experimental paradigm, participants practiced for as long as they wanted, with the foreknowledge that their skills would subsequently be tested (and rewarded monetarily) without access to AI. During practice, therefore, participants were incentivized to prioritize gains in acquired skill over performance in the moment. In real-world settings, there is often time pressure and competing incentives around performance and learning, which we speculate would reduce the learning gains associated with practicing with AI.
Third, in our experimental paradigm, participants who interacted with the AI tool did so only once. It is common, however, to use AI tools repeatedly. When do repeated interactions lead to diminishing or even negative returns, and in what scenarios might skill development continue over time? Consider, for instance, the game of Go. The introduction of superhuman AI has been associated with an increase in the novelty and quality of decisions made by human Go players over time, with elite players reporting that they have been inspired by AI solutions they’d never seen before (32). Future research, ideally in field settings, is needed to establish the long-term benefits and costs of relying on AI tools.
Conclusion
Our findings should temper widespread concern that AI tools invariably boost momentary productivity at the expense of long-term skill development. Although it reduced the effort users invested in practicing, the AI writing tool nevertheless accelerated skill development. It accomplished this by providing high-quality, just-in-time, personalized examples of excellent writing. The underappreciated efficacy of timely and tailored examples has practical implications for the design of AI tools. Many AI tools designed to support learning are explicitly programmed not to “give away” answers. It may be that in addition to hints, leading questions, and explanations, learners benefit from demonstrations of the principles they are attempting to master.
Decades before the advent of generative AI, the legendary UCLA baseball coach John Wooden declared that the four laws of learning are explanation, demonstration, imitation, and repetition (33). Few learners have access to the best human teachers, coaches, and mentors, but generative AI now makes it possible to learn from personalized, just-in-time demonstrations tailored to any domain. In doing so, AI has the potential not only to boost productivity but also to democratize opportunities to build human capital at scale.
Methods
Ethical Considerations
The study was assessed by the University of Pennsylvania’s IRB, and was approved before implementation (Protocol 853653). All participants completed informed consent.
Pre-registration
All our studies were pre-registered. See pre-registrations for Study 1 at https://aspredicted.org/x9mm-7qwp.pdf.
Study 2 was run twice because of a technical issue that caused imbalanced missing data issue the first time it was run. The findings in this sample were consistent with the ones reported here. Pre-registration is available at https://aspredicted.org/4sw4-mpny.pdf. The version of Study 2 presented above was pre-registered https://aspredicted.org/xyyn-gmzc.pdf.
Pre-registartion for Study 3 is available at https://aspredicted.org/3mty-fcfy.pdf. The one-day follow-up for Study 3 was collected in three batches. We pre-registered batch 2 https://aspredicted.org/5jcx-bhg9.pdf, but report pooled results in the main text. Details on the separate batches are available in Supplementary Information Section C5.
Participants
We sampled participants from Prolific from all our studies. We excluded all Prolific users who participated in one of the earlier studies from participation in subsequent studies.
Participants in Study 1 (N = 150) were predominantly female (n = 93, 62%), and ranged in age from 21 to 81 (M = 38.4, SD = 12.2). They were predominantly white (75%). A small proportion were students (13%), and most were employed (62%)
In Study 2, the sample was more evenly split between men (46%) and women (52%), and ranged in age from 18 to 82 (M = 36.0, SD = 12.5). Over half of the sample (58%) was white, with the rest being comprised by Black (33%), latino (6%), and asian (5%). 77% had college degrees. Most participants (93%) were at least somewhat motivated to improve their writing, and had varying levels of experience with AI writing assistants (36% had tried them, but hardly ever used them, 47% used them at least a few times per week, and 17% had never used AI assistants before).
Study 3 had similar prooportions of men (46%) and women (53%), and participants ranged in age from 18 to 95 (M = 37.9, SD = 12.6). Over half of the sample (64%) was white, with the rest being comprised of Black (24%), Latino (8%), and Asian (6%) participants. Most participants (74%) had college degrees. Most participants (91%) were at least somewhat motivated to improve their writing, and had varying levels of experience with AI writing assistants (40% had tried them, but hardly ever used them, 42% used them at least a few times per week, and 18% had never used AI assistants before).
Procedure
Participants first saw an introductory screen about what they were about to do. They then completed a brief questionnaire where they reported their demographics, their experience with AI writing tools, their motivation to improve their writing, and their perceived writing skill. They then completed a 2-minute pre-test in which they saw a poorly worded cover letter and were asked to improve it. After this, all participants completed a lesson about the five principles of effective writing. They were then randomized to the practice condition, or skipped ahead if assigned to the no-practice control. During the practice phase, participants rewrote a different cover letter, or (in the see example condition) generated an AI rewritten version of that letter. This example was not explicitly labeled as AI-generated. Then, participants saw the text they submitted (or the AI-generated example), and below it, AI-generated feedback highlighting one way in which this letter could be improved. See more information about the feedback procedure in Supplementary Information section A3. Immediately after seeignt this feedback screen, all participants then completed two questions, reporting how much they had learned and how hard they had worked on the task so far. They then edited a new cover letter in an incentivized, 7-minute test without the help of any AI tools. To minimize the possibility of cheating, we used custom JavaScript to restrict copy-pasting functionality. Finally, participants were invited to see optional feedback and were asked if they had used any outside resources during the test.
Measurement
As per our pre-registration, we used OpenAI’s GPT-4o to rate text samples for writing quality. To do this, we independently rated each cover letter and each writing principle. Research has demonstrated that large language models can provide ratings of writing quality that align closely with human judgments, offering reliability and consistency across various evaluation contexts (34, 35). See our prompts in Table S1. We then took the unweighted average of these 5 scores as our main dependent variable. See disaggregated analyses by each writing principle and additional outcomes on Table S6 and S16.
To validate these ratings, the first author and a trained research assistant took a sample of n = 30 cover letters from Study 2, and rated them on the 5 principles. The average of these two ratings correlated more highly with the computer ratings (r = .83, p < .001), than the average interrater reliability (r = .69, p < .001).
To address concerns that particular LLMs might be biased in favor of their own output, we also used Claude to rate the cover letters. We find that GPT ratings correlate highly with Claude ratings (r = .71, p < .001), and that the effects are not attenuated by using different models (See Tables S6 and S16), suggesting that our effects are not explainable by same-model bias.
Statistical analysis
As per our pre-registrations, we fit ANCOVA models predicting outcomes from condition indicators, controlling for pretest score and baseline characteristics (age, gender, race/ethnicity, primary language, education level, motivation to improve writing skills, self-rated writing skill, experience with AI writing assistants, and baseline writing effectiveness). We used logistic regression to predict whether participants chose to see optional feedback for their test from condition, controlling for pretest score and baseline characteristics. Our analyses of the hypothetical hiring situation use beta regression, because the relative likelihood of a cover letter being preferred is bounded between 0 and 1. When correcting for multiple comparisons in exploratory moderation analyses, we use the Benjamini-Hochberg correction (36).
C Results Study 3
- C1 Randomization, Balance, and Missingness
- C2 AI examples improve writing skill
- C3 Seeing AI examples was less effortful
- C4 Seeing an AI example did not discourage motivation for future learning
- C5 The effects of seeing an AI example persist
- C6 Seeing AI examples was equally effective across subgroups
Appendix A Additional methods
A1 AI Ratings
After participants completed the procedure outlined in Figure 2, we had three writing samples for participants who practiced with or without AI (one for each phase of the experiment: pretest, practice, test), and two writing samples for participants who did not practice, or simply saw an AI generated example. We used GPT-4o to rate these texts for five writing principles. Each text and rating was completed independently of each other (i.e., the model had no memory of seeing that text before or of having rated it for any of the other writing principles). For robustness checks, we also used Anthropic’s Claude Haiku.
Table S1 shows the prompts used to have GPT-4o and Claude rate the rewritten cover letters on the five principles. Our pre-registered main outcome is the unweighted mean of these five principles.
| Writing principle | Rating prompt |
| Less is more | On a 0 10 scale, how much does the text follow the less is more principle? The text should use as few words as needed, as few ideas as needed, and make as few requests as needed. |
| Easy reading | On a 0 10 scale, how much does the text make reading easy. The text should use short and common words, use straightforward sentences, and shorter sentences. |
| Easy navigation | On a 0 10 scale, how much does the text make navigation easy. The text should make key information immediately visible, separate distinct ideas, place related ideas together, order ideas by priority, include headings when necessary, and use visuals if needed. |
| Formatting | On a 0 10 scale, how much does the text use appropriate formatting. The text should follow readers expectations regarding formatting, use bolding, italics, underline, or highlight to draw attention to the most important ideas, and should not overdo formatting. |
| Easy responding | On a 0 10 scale, how much does the text make responding easy. The text should simplify the steps required to act, organize the key information needed for action, and minimize the amount of attention required. |
We used a randomly selected sample of 100 cover letters from Study 2 to validate the AI ratings. Two trained raters independently rated each of the 5 principles. The inter-rater correlation was r = .74, p < .001. The human raters correlated with the AI-generated ratings satisfactorily (r = .70, p < .001). See Figure S1.
Claude and GPT-4o ratings were positively correlated, with pretest ratings being less so. In Study 2, the correlations were .38, .78, .65, and .67 for pretest, practice, test, and followup, respectively. All p-values were below .001.
A2 Pairwise Comparisons
Prolific participants were shown pairs of cover letters sampled from different conditions. They were asked to “Imagine you’re hiring a social media manager for your company; which cover letter would make you more likely to offer an interview to the candidate? Choose one.”. Each cover letter was compared to at least three other letters, sampled uniformly at random from the other two conditions. Most letters were compared against 3 or 4 other letters, min, max, med (plot?). For each cover letter we calculated the relative likelihood of it securing a hypothetical interview, defined as the total number of times that letter was preferred, divided by the total number of contests for that cover letter.
As shown in Figure S2, the relative likelihood of an interview correlated positively with the AI-generated measure of writing skill.
Table S2 shows the beta regressions for the relative likelihood by condition.
| (1) | (2) | (3) | (4) | |
| (Intercept) | *** | ** | ||
| No practice | *** | *** | ||
| Practice wo AI | ** | |||
| Practice w AI | ** | |||
| See AI example | *** | |||
| Precision () | *** | *** | ||
| Symmertry (Log()) | *** | |||
| Demographic and baseline performance covariates | No | Yes | No | Yes |
| Num.Obs. | ||||
| AIC | ||||
| BIC | ||||
| Log.Lik. | ||||
| RMSE |
A3 Feedback
In Studies 2 and 3, participants received feedback for their submissions. The feedback was displayed immediately after the practice cover letter submission. The feedback page read: “Here is the email”, then reproduced the participants submission verbatim, then read “Here is one way in which it could be made better.” The feedback was personalized and created by GPT-4o model.
The feedack prompt is shown in Figure S3.
Appendix B Results Study 2
B1 Randomization, Balance, and Missingness
To allow users to format their responses flexibly, we used the TinyMCE rich text editor, which is interfaced with Qualtrics. While this allowed users to use bolding, lists, and italicizing, a small percentage of users experienced technical issues that resulted in their text data not being recorded (3.31%). These users did type in the box, as evidenced by their time and keystroke data, and completed the experiment.
There was also attrition in the follow-up sample. While most people responded, 13.45% of recontacted participants did not respond. This attrition was not selective by condition. As shown in Table S3, missingness and attrition rates were low for the main and followup samples, and did not differ by condition.
| Condition | Main Sample | Followup Sample |
| No practice | ||
| Practice wo AI | ||
| Practice w AI | ||
| Overall | ||
| p-value |
Pre-treatment variables were balanced across experimental conditions, ensuring that random assignment was successful. To assess balance, we conducted a series of one-way ANOVAs for continuous variables and chi-square tests for categorical variables. Given the multiple comparisons, we applied the Benjamini-Hochberg (BH) procedure to control the false discovery rate. All statistical tests confirmed that none of the pre-treatment variables differed significantly across conditions. See Table S4.
| Overall | No practice | Practice wo AI | Practice w AI | p | SMD | |
| n | 2238 | 755 | 752 | 731 | ||
| Age (mean (SD)) | 36.22 (12.71) | 35.99 (13.01) | 36.39 (12.70) | 36.29 (12.42) | .923 | 0.021 |
| Gender (%) | .636 | 0.077 | ||||
| Female | 1189 (53.1) | 421 (55.8) | 394 (52.4) | 374 (51.2) | ||
| Male | 1027 (45.9) | 328 (43.4) | 348 (46.3) | 351 (48.0) | ||
| Other | 22 (1.0) | 6 (0.8) | 10 (1.3) | 6 (0.8) | ||
| Race/Ethnicity | ||||||
| White (%) | 1288 (57.6) | 419 (55.5) | 458 (60.9) | 411 (56.2) | .213 | 0.073 |
| Black (%) | 745 (33.3) | 262 (34.7) | 223 (29.7) | 260 (35.6) | .192 | 0.084 |
| Asian (%) | 134 (6.0) | 49 (6.5) | 42 (5.6) | 43 (5.9) | .923 | 0.025 |
| Latino (%) | 155 (6.9) | 48 (6.4) | 65 (8.6) | 42 (5.7) | .213 | 0.075 |
| Other (%) | 62 (2.8) | 23 (3.0) | 22 (2.9) | 17 (2.3) | .923 | 0.030 |
| Education Level (%) | .923 | 0.099 | ||||
| Less than high school degree | 14 (0.6) | 5 (0.7) | 5 (0.7) | 4 (0.5) | ||
| High school graduate (high school diploma or equivalent including GED) | 207 (9.2) | 68 (9.0) | 72 (9.6) | 67 (9.2) | ||
| Some college but no degree | 321 (14.3) | 117 (15.5) | 109 (14.5) | 95 (13.0) | ||
| Associate degree in college (2-year) | 168 (7.5) | 61 (8.1) | 60 (8.0) | 47 (6.4) | ||
| Bachelor’s degree in college (4-year) | 984 (44.0) | 314 (41.6) | 334 (44.4) | 336 (46.0) | ||
| Master’s degree | 474 (21.2) | 165 (21.9) | 153 (20.3) | 156 (21.3) | ||
| Doctoral degree (PhD) | 44 (2.0) | 16 (2.1) | 11 (1.5) | 17 (2.3) | ||
| Non-PhD Professional degree (JD, MD) | 26 (1.2) | 9 (1.2) | 8 (1.1) | 9 (1.2) | ||
| Perceived Writing Skill (mean (SD)) | 6.70 (1.70) | 6.77 (1.67) | 6.56 (1.72) | 6.76 (1.69) | .192 | 0.083 |
| Motivation to improve writing (%) | .410 | 0.126 | ||||
| Not at all motivated | 33 (1.5) | 6 (0.8) | 15 (2.0) | 12 (1.6) | ||
| Hardly motivated | 106 (4.7) | 38 (5.0) | 34 (4.5) | 34 (4.7) | ||
| Somewhat motivated | 644 (28.8) | 218 (28.9) | 236 (31.4) | 190 (26.0) | ||
| Very motivated | 932 (41.6) | 311 (41.2) | 311 (41.4) | 310 (42.4) | ||
| Extremely motivated | 523 (23.4) | 182 (24.1) | 156 (20.7) | 185 (25.3) | ||
| Experience with AI writing assistants (%) | .701 | 0.094 | ||||
| I have never tried any AI writing assistant | 354 (15.8) | 109 (14.4) | 139 (18.5) | 106 (14.5) | ||
| I have tried AI writing assistant(s) but hardly ever use them | 859 (38.4) | 291 (38.5) | 288 (38.3) | 280 (38.3) | ||
| I use AI writing assistant(s) a few times per week | 477 (21.3) | 164 (21.7) | 147 (19.5) | 166 (22.7) | ||
| I use AI writing assistant(s) about once a week | 395 (17.6) | 136 (18.0) | 133 (17.7) | 126 (17.2) | ||
| I use AI writing assistant(s) every day | 153 (6.8) | 55 (7.3) | 45 (6.0) | 53 (7.3) | ||
| Pretest Writing Skill (mean (SD)) | 3.32 (0.78) | 3.31 (0.73) | 3.32 (0.78) | 3.32 (0.83) | .923 | 0.013 |
B2 AI practice improved writing skill
The AI tool improved performance while participants used it. Table S5 shows means and standardized differences for different measures of writing skill during the practice phase. The robustness checks included after the main specification, show that results are similar when using a different language model (Column 2), when not including control variables (Column 3), when excluding participants who admitted to cheating in the test phase (Column 4), for the subset of non-attriting participants to the follow-up phase (Column 5), and for each of the 5 principles separately (Columns 6 - 10).
| GPT-4o | Claude | Ex. Controls | Ex. Cheaters | Followup | LM | ER | EN | F | ER | |
| Means — (SE) | ||||||||||
| Practice wo AI | 4.58 | 6.52 | 4.41 | 4.42 | 4.27 | 4.29 | 6.31 | 5.65 | 3.38 | 4.26 |
| (.222) | (.091) | (.054) | (.055) | (.281) | (.230) | (.150) | (.258) | (.480) | (.268) | |
| Practice w AI | 6.05 | 7.01 | 5.86 | 5.89 | 5.78 | 5.56 | 7.11 | 6.75 | 6.30 | 5.54 |
| (.224) | (.092) | (.055) | (.055) | (.286) | (.232) | (.151) | (.260) | (.484) | (.271) | |
| Effect Sizes (d) — (SE) | ||||||||||
| Practice wo AI vs. Practice w AI | 1.01*** | .81*** | .98*** | 1.00*** | 1.05*** | .84*** | .82*** | .65*** | .93*** | .73*** |
| (.056) | (.055) | (.055) | (.056) | (.061) | (.055) | (.055) | (.054) | (.056) | (.055) | |
| Note. GPT-4o is the main specification. Ex. Controls is the main specification, unadjusted for demographic and pretreatment variables, Ex. Cheaters excludes the 3% of participants who admitted to cheating on the test phase. Followup is the subsample of non-attriting participants who returned to the one-day followup. LM to ER are disagregated scores for each of the five principles. LM = Less is More, ER = Easy Reading, EN = Easy Navigation, F = Formatting, ER = Easy Responding. *** p < .001, ** p < .01, * p < .05. | ||||||||||
During the test phase, when participants had to rewrite a cover letter without the help of the AI tool, participants who had practiced with AI outperformed participants who had not practiced, or had practiced without the AI tool. Again, the learning gains are robust to different specifications, subsamples, and measures or writing quality. See Table S6. For participants assigned to practice with the AI tool, The quality of AI rewrites did not correlate with participants’ final submissions, r = .06, p = .25.
| GPT-4o | Claude | Ex. Controls | Ex. Cheaters | Followup | LM | ER | EN | F | ER | |
| Means — (SE) | ||||||||||
| No practice | 4.41 | 6.70 | 4.39 | 4.39 | 4.52 | 3.71 | 5.90 | 5.55 | 2.47 | 4.44 |
| (.161) | (.072) | (.047) | (.048) | (.200) | (.155) | (.143) | (.192) | (.394) | (.202) | |
| Practice wo AI | 4.53 | 6.74 | 4.51 | 4.52 | 4.63 | 3.84 | 6.03 | 5.56 | 2.76 | 4.45 |
| (.160) | (.071) | (.048) | (.048) | (.199) | (.154) | (.142) | (.190) | (.392) | (.200) | |
| Practice w AI | 5.01 | 6.90 | 4.98 | 4.99 | 5.11 | 4.12 | 6.21 | 6.17 | 3.86 | 4.66 |
| (.161) | (.072) | (.049) | (.049) | (.202) | (.155) | (.143) | (.192) | (.394) | (.202) | |
| Effect Sizes (d) — (SE) | ||||||||||
| No practice vs. Practice wo AI | .09 | .09 | .09 | .10 | .09 | .10* | .11* | .01 | .09 | .01 |
| (.053) | (.053) | (.052) | (.053) | (.056) | (.053) | (.053) | (.053) | (.053) | (.053) | |
| No practice vs. Practice w AI | .47*** | .36*** | .46*** | .46*** | .48*** | .34*** | .28*** | .42*** | .46*** | .14** |
| (.054) | (.053) | (.053) | (.054) | (.057) | (.053) | (.053) | (.053) | (.054) | (.053) | |
| Practice wo AI vs. Practice w AI | .38*** | .28*** | .36*** | .36*** | .39*** | .23*** | .17** | .41*** | .36*** | .13* |
| (.054) | (.053) | (.053) | (.054) | (.057) | (.054) | (.053) | (.054) | (.054) | (.053) | |
| Note. GPT-4o is the main specification. Ex. Controls is the main specification, unadjusted for demographic and pretreatment variables, Ex. Cheaters excludes the 3% of participants who admitted to cheating on the test phase. Followup is the subsample of non-attriting participants who returned to the one-day followup. LM to ER are disagregated scores for each of the five principles. LM = Less is More, ER = Easy Reading, EN = Easy Navigation, F = Formatting, ER = Easy Responding. *** p < .001, ** p < .01, * p < .05. | ||||||||||
B3 AI practice was less effortful
Table S7 shows OLS models predicting practice effort metrics from practice condition. Results show that participants practicing without AI expended more effort, measured subjectively or objectively, through keystrokes or practice time. As pre-registered, time is square-root-transformed, and keystrokes are log-transformed. Differences are slightly smaller when using untransformed variables.
| sqrt(Time) | log(Keystrokes) | Subjective Rating (0 - 10) | Time | Keystrokes | |
| Means — (SE) | |||||
| Practice wo AI | 2.37 | 4.31 | 6.52 | 6.76 | 430.95 |
| (.152) | (.322) | (.291) | (.913) | (57.349) | |
| Practice w AI | 2.30 | 3.36 | 5.93 | 6.62 | 383.38 |
| (.153) | (.325) | (.293) | (.919) | (57.887) | |
| Effect Sizes (d) — (SE) | |||||
| Practice wo AI vs. Practice w AI | -.07 | -.45*** | -.31*** | -.02 | -.13* |
| (.053) | (.054) | (.053) | (.053) | (.053) | |
| Note. *** p < .001, ** p < .01, * p < .05. | |||||
Table S8 shows OLS models predicting test effort metrics from practice condition. Results show some differences: participants who practiced with AI pressed more keys but reported less subjective effort.
| sqrt(Time) | log(Keystrokes) | Subjective Rating (0 - 10) | Time | Keystrokes | |
| Means — (SE) | |||||
| No practice | 2.32 | 5.01 | 6.55 | 5.62 | 399.10 |
| (.069) | (.213) | (.266) | (.257) | (41.391) | |
| Practice wo AI | 2.15 | 4.87 | 6.91 | 4.98 | 409.09 |
| (.068) | (.211) | (.265) | (.255) | (41.143) | |
| Practice w AI | 2.19 | 5.05 | 6.69 | 5.14 | 446.59 |
| (.069) | (.213) | (.267) | (.257) | (41.408) | |
| Effect Sizes (d) — (SE) | |||||
| No practice vs. Practice wo AI | -.31*** | -.09 | .18*** | -.32*** | .03 |
| (.053) | (.053) | (.053) | (.053) | (.053) | |
| No practice vs. Practice w AI | -.24*** | .02 | .07 | -.24*** | .15** |
| (.053) | (.053) | (.053) | (.053) | (.053) | |
| Practice wo AI vs. Practice w AI | .07 | .11* | -.11* | .08 | .12* |
| (.053) | (.053) | (.053) | (.053) | (.053) | |
| Note. *** p < .001, ** p < .01, * p < .05. | |||||
Table S9 shows OLS models predicting learning rate metrics from practice condition. Learning rate is defined as the difference between test and pretest, divided by the effort metric. It shows how many points (10 point scale) the participant improved per unit effort (e.g., per minute spent practicing). Participants who practiced with AI improved their skill more efficiently.
| sqrt(Time) | log(Keystrokes) | Subjective Rating (0 - 10) | Time | Keystrokes | |
| Means — (SE) | |||||
| Practice wo AI | .28 | .12 | .27 | .32 | .43 |
| (.062) | (.094) | (.038) | (.096) | (.073) | |
| Practice w AI | .43 | .36 | .37 | .61 | .62 |
| (.062) | (.094) | (.038) | (.097) | (.073) | |
| Effect Sizes (d) — (SE) | |||||
| Practice wo AI vs. Practice w AI | .39*** | .40*** | .41*** | .47*** | .40*** |
| (.054) | (.054) | (.054) | (.054) | (.054) | |
| Note. *** p < .001, ** p < .01, * p < .05. | |||||
Most participants did not engage passively with the AI tool. As shown in Figure S5, an overwhelming majority of participants changed the AI tool’s output text before submitting it as their answer. A smaller proportion of participants even edited the cover letter email before passing it along to the AI tool.
B4 Seeing an AI example did not discourage motivation for future learning
Table S10 presents differences in perceived learning, perceived writing skill, and the likelihood of asking for feedback across conditions, with effect sizes and means reported for each comparison. Despite objectively learning more, participants who practiced with AI perceived their learning and skill levels to be similar to those who practiced without AI and asked for feedback at comparable rates.
| Perceived learning | Perceived writing skill | Asked for feedback | |
| Means/Proportions | |||
| No practice | 5.91 | 6.40 | .64 |
| (.209) | (.208) | (.064) | |
| Practice wo AI | 5.90 | 6.61 | .62 |
| (.207) | (.206) | (.066) | |
| Practice w AI | 6.03 | 6.56 | .58 |
| (.209) | (.208) | (.068 ) | |
| Effect Sizes (ds/odds ratios) | |||
| No practice vs. Practice wo AI | -.00 | .13* | 1.10 |
| (.053) | (.053) | (.128) | |
| No practice vs. Practice w AI | .08 | .10 | 1.30* |
| (.053) | (.053) | (.149) | |
| Practice wo AI vs. Practice w AI | .08 | -.03 | 1.17 |
| (.053) | (.053) | (.134) | |
| Note. *** p < .001, ** p < .01, * p < .05. | |||
B5 The effects of practicing with AI persist
Table S11 shows means and standardized differences for measures of writing skill and related outcomes during the follow-up phase. The main specification demonstrates that participants who practiced with AI continued to outperform those who did not practice or practiced without AI. Robustness checks, including using a different language model (Column 2), excluding control variables (Column 3), and removing participants who admitted to cheating (Column 4) confirm the consistency of these effects. The results also hold when evaluating each of the five principles separately (Columns 5–9). These findings suggest that the benefits of practicing with AI are durable and persist even after participants stop using the tool.
| GPT-4o | Claude | Ex. Controls | Ex. Cheaters | LM | ER | EN | F | ER | |
| Means — (SE) | |||||||||
| No practice | 4.73 | 6.75 | 4.75 | 4.75 | 4.31 | 6.36 | 5.56 | 2.43 | 5.01 |
| (.212) | (.094) | (.054) | (.054) | (.201) | (.206) | (.235) | (.510) | (.267) | |
| Practice wo AI | 4.79 | 6.78 | 4.84 | 4.86 | 4.44 | 6.45 | 5.52 | 2.59 | 4.96 |
| (.211) | (.093) | (.053) | (.054) | (.200) | (.205) | (.234) | (.507) | (.266) | |
| Practice w AI | 5.34 | 6.95 | 5.35 | 5.37 | 4.72 | 6.67 | 6.14 | 3.85 | 5.30 |
| (.214) | (.094) | (.055) | (.055) | (.203) | (.208) | (.237) | (.515) | (.270) | |
| Effect Sizes (d) — (SE) | |||||||||
| No practice vs. Practice wo AI | .05 | .06 | .07 | .08 | .11 | .08 | -.02 | .05 | -.03 |
| (.056) | (.054) | (.055) | (.056) | (.056) | (.056) | (.056) | (.056) | (.056) | |
| No practice vs. Practice w AI | .46*** | .33*** | .44*** | .45*** | .32*** | .25*** | .40*** | .45*** | .17** |
| (.057) | (.055) | (.056) | (.057) | (.056) | (.056) | (.057) | (.057) | (.056) | |
| Practice wo AI vs. Practice w AI | .41*** | .28*** | .37*** | .37*** | .22*** | .17** | .42*** | .40*** | .20*** |
| (.057) | (.055) | (.056) | (.056) | (.056) | (.056) | (.057) | (.057) | (.056) | |
| Note. GPT-4o is the main specification. Ex. Controls is the main specification, unadjusted for demographic and pretreatment variables, Ex. Cheaters excludes the 3% of participants who admitted to cheating on the test phase. LM to ER are disagregated scores for each of the five principles. LM = Less is More, ER = Easy Reading, EN = Easy Navigation, F = Formatting, ER = Easy Responding. *** p < .001, ** p < .01, * p < .05. | |||||||||
B6 AI practice was equally effective across subgroups
To test for moderation effects of pre-treatment demographic variables, we ran separate linear in which writing skill during the test phase was regressed on condition, the pre-treatment moderator of interest, writing skill at baseline, and an interaction term between the moderator condition. After correcting the p-values for the interaction terms, none were significant at the .05 level, suggesting that practicing with AI was equally effective across groups.
| Test | Follow-Up | Time Practice | Keys Practice | Effort Practice | Per. Learning | Per. Skill | Want Feedback | ||||||
| Level | No AI | With AI | No AI | With AI | With AI | With AI | With AI | No AI | With AI | No AI | With AI | No AI | With AI |
| Continuous Moderators | |||||||||||||
| Pretest | 0.955 | 0.914 | 0.631 | 0.699 | 0.984 | 0.914 | 0.941 | 0.820 | 0.914 | 0.914 | 0.914 | 0.574 | 0.851 |
| Year of Birth | 0.533 | 0.931 | 0.868 | 0.955 | 0.914 | 0.955 | 0.914 | 0.914 | 0.955 | 0.955 | 0.914 | 0.914 | 0.618 |
| Writing Skill | 0.914 | 0.914 | 0.955 | 0.545 | 0.913 | 0.838 | 0.914 | 0.851 | 0.914 | 0.914 | 0.919 | 0.914 | 0.914 |
| Gender (vs. Female) | |||||||||||||
| Male | 0.955 | 0.914 | 0.970 | 0.955 | 0.699 | 0.914 | 0.955 | 0.699 | 0.979 | 0.914 | 0.851 | 0.699 | 0.533 |
| Other | 0.719 | 0.919 | 0.914 | 0.955 | 0.931 | 0.955 | 0.931 | 0.914 | 0.914 | 0.955 | 0.931 | 0.876 | 0.913 |
| Race | |||||||||||||
| White | 0.914 | 0.955 | 0.914 | 0.574 | 0.955 | 0.914 | 0.914 | 0.737 | 0.931 | 0.643 | 0.851 | 0.533 | 0.699 |
| Black | 0.914 | 0.913 | 0.914 | 0.851 | 0.914 | 0.914 | 0.955 | 0.574 | 0.919 | 0.295 | 0.699 | 0.574 | 0.533 |
| Asian | 0.973 | 0.973 | 0.955 | 0.737 | 0.964 | 0.931 | 0.955 | 0.931 | 0.890 | 0.919 | 0.955 | 0.914 | 0.699 |
| Latino | 0.919 | 0.914 | 0.970 | 0.914 | 0.914 | 0.868 | 0.533 | 0.868 | 0.944 | 0.533 | 0.914 | 0.970 | 0.574 |
| Other | 0.973 | 0.914 | 0.914 | 0.973 | 0.914 | 0.931 | 0.643 | 0.914 | 0.914 | 0.955 | 0.914 | 0.533 | 0.533 |
| Motivation (vs. Not at all) | |||||||||||||
| Hardly | 0.914 | 0.914 | 0.955 | 0.955 | 0.914 | 0.955 | 0.964 | 0.914 | 0.914 | 0.955 | 0.931 | 0.699 | 0.533 |
| Somewhat | 0.876 | 0.663 | 0.973 | 0.913 | 0.914 | 0.914 | 0.914 | 0.973 | 0.964 | 0.914 | 0.955 | 0.663 | 0.533 |
| Very | 0.851 | 0.566 | 0.980 | 0.914 | 0.851 | 0.968 | 0.964 | 0.970 | 0.982 | 0.914 | 0.973 | 0.699 | 0.533 |
| Extremely | 0.861 | 0.533 | 0.973 | 0.868 | 0.914 | 0.946 | 0.914 | 0.931 | 0.964 | 0.914 | 0.964 | 0.749 | 0.533 |
| Experience with AI writing assistants (vs. None) | |||||||||||||
| Hardly ever | 0.931 | 0.914 | 0.931 | 0.868 | 0.914 | 0.533 | 0.914 | 0.574 | 0.931 | 0.931 | 0.914 | 0.955 | 0.964 |
| A few times per week | 0.914 | 0.667 | 0.955 | 0.955 | 0.914 | 0.533 | 0.914 | 0.574 | 0.919 | 0.931 | 0.914 | 0.914 | 0.955 |
| About once a week | 0.876 | 0.214 | 0.914 | 0.955 | 0.955 | 0.699 | 0.914 | 0.574 | 0.914 | 0.931 | 0.919 | 0.914 | 0.970 |
| Every day | 0.914 | 0.914 | 0.914 | 0.964 | 0.919 | 0.914 | 0.984 | 0.914 | 0.931 | 0.955 | 0.914 | 0.955 | 0.931 |
| Note. Models for test and follow-up performance, square-root practice time, log keystrokes, subjective effort, perceived learning and perceived writing skill or OLS models. Asking to see feedback was a binary Yes/No variable, and was modelled with logistic regression. Models match the pre-registered main specification, and thus control for all other pre-treatment variables. Per. = Perceived | |||||||||||||
Appendix C Results Study 3
C1 Randomization, Balance, and Missingness
As in Study 2, technical issues caused small amounts of missing data. Overall, 5.64% of data was missing in for the test phase analysis, which was not differentially missing by condition. There was also attrition in the follow-up sample. While most people responded, 13.45% of recontacted participants did not respond. This attrition was not selective by condition. As shown in Table S13, missingness and attrition rates were low for the main and follow-up samples and did not differ by condition.
| Condition | Main Sample | Followup Sample |
| Practice wo AI | ||
| Practice w AI | ||
| See AI example | ||
| Overall | ||
| p-value |
Pre-treatment variables were balanced across experimental conditions, ensuring that random assignment was successful. To assess balance, we conducted a series of one-way ANOVAs for continuous variables and chi-square tests for categorical variables. Given the multiple comparisons, we applied the Benjamini-Hochberg (BH) procedure to control the false discovery rate. All statistical tests confirmed that none of the pre-treatment variables differed significantly across conditions. See Table S14.
| Overall | Practice wo AI | Practice w AI | See AI example | p | SMD | |
| n | 2003 | 672 | 652 | 679 | ||
| Age (mean (SD)) | 37.89 (12.63) | 37.77 (12.37) | 37.87 (12.85) | 38.03 (12.69) | .997 | 0.014 |
| Gender (%) | .822 | 0.055 | ||||
| Female | 1056 (52.7) | 341 (50.7) | 350 (53.7) | 365 (53.8) | ||
| Male | 923 (46.1) | 321 (47.8) | 296 (45.4) | 306 (45.1) | ||
| Other | 24 (1.2) | 10 (1.5) | 6 (0.9) | 8 (1.2) | ||
| Race/Ethnicity (%) | ||||||
| White = 1 | 1287 (64.3) | 419 (62.4) | 430 (66.0) | 438 (64.5) | .655 | 0.050 |
| Black = 1 | 484 (24.2) | 184 (27.4) | 144 (22.1) | 156 (23.0) | .324 | 0.082 |
| Asian = 1 | 127 (6.3) | 37 (5.5) | 43 (6.6) | 47 (6.9) | .715 | 0.039 |
| Latino = 1 | 163 (8.1) | 55 (8.2) | 46 (7.1) | 62 (9.1) | .655 | 0.051 |
| Other = 1 | 3 (0.1) | 1 (0.1) | 2 (0.3) | 0 (0.0) | .655 | 0.055 |
| Education Level (%) | .655 | 0.152 | ||||
| Less than high school degree | 10 (0.5) | 3 (0.4) | 3 (0.5) | 4 (0.6) | ||
| High school graduate | 205 (10.2) | 74 (11.0) | 61 (9.4) | 70 (10.3) | ||
| Some college, no degree | 305 (15.2) | 104 (15.5) | 110 (16.9) | 91 (13.4) | ||
| Associate degree | 169 (8.4) | 68 (10.1) | 45 (6.9) | 56 (8.2) | ||
| Bachelor’s degree | 850 (42.4) | 255 (37.9) | 290 (44.5) | 305 (44.9) | ||
| Master’s degree | 401 (20.0) | 144 (21.4) | 126 (19.3) | 131 (19.3) | ||
| Doctoral degree (PhD) | 36 (1.8) | 14 (2.1) | 11 (1.7) | 11 (1.6) | ||
| Professional degree (JD, MD) | 27 (1.3) | 10 (1.5) | 6 (0.9) | 11 (1.6) | ||
| Writing Skill (mean (SD)) | 6.60 (1.70) | 6.63 (1.67) | 6.71 (1.69) | 6.46 (1.73) | .228 | 0.100 |
| Motivation (%) | .997 | 0.042 | ||||
| Not at all motivated | 28 (1.4) | 9 (1.3) | 10 (1.5) | 9 (1.3) | ||
| Hardly motivated | 154 (7.7) | 50 (7.4) | 53 (8.1) | 51 (7.5) | ||
| Somewhat motivated | 639 (31.9) | 221 (32.9) | 202 (31.0) | 216 (31.8) | ||
| Very motivated | 762 (38.0) | 249 (37.1) | 249 (38.2) | 264 (38.9) | ||
| Extremely motivated | 420 (21.0) | 143 (21.3) | 138 (21.2) | 139 (20.5) | ||
| Experience with AI (%) | .655 | 0.103 | ||||
| Never used AI writing assistant | 351 (17.5) | 128 (19.0) | 105 (16.1) | 118 (17.4) | ||
| Tried AI but hardly use | 807 (40.3) | 267 (39.7) | 269 (41.3) | 271 (39.9) | ||
| Use AI a few times per week | 375 (18.7) | 108 (16.1) | 133 (20.4) | 134 (19.7) | ||
| Use AI about once a week | 343 (17.1) | 127 (18.9) | 102 (15.6) | 114 (16.8) | ||
| Use AI every day | 127 (6.3) | 42 (6.2) | 43 (6.6) | 42 (6.2) | ||
| Pretest Writing Skill (mean (SD)) | 4.21 (0.88) | 4.23 (0.90) | 4.24 (0.90) | 4.17 (0.84) | .655 | 0.051 |
C2 AI examples improve writing skill
The AI tool improved performance while participants used it. Table S15 shows means and standardized differences for different measures of writing skill during the practice phase. The robustness checks included after the main specification, show that results are similar when using a different language model (Column 2), when not including control variables (Column 3), when excluding participants who admitted to cheating in the test phase (Column 4), for the subset of non-attriting participants to the follow-up phase (Column 5), and for each of the 5 principles separately (Columns 6 - 10).
During the test phase, when participants had to rewrite a cover letter without the help of the AI tool, participants who simply had seen an AI example outperformed participants who had practiced without the AI tool, and perfomed comparably to those who had practiced with the AI tool. Replicating Study 2, participants who had practiced with the AI tool performed better than those who had practiced without it. Again, the learning gains are robust to different specifications, subsamples, and measures or writing quality. See Table S16. For participants assigned to practice with the AI tool, The quality of AI rewrites did not correlate with participants’ final submissions, r = .06, p = .25.
| GPT-4o | Claude | Ex. Controls | Ex. Cheaters | Followup | LM | ER | EN | F | ER | |
| Means — (SE) | ||||||||||
| Practice wo AI | 4.72 | 4.96 | 4.62 | 4.61 | 4.64 | 4.33 | 6.61 | 5.43 | 2.93 | 4.31 |
| (.374) | (.379) | (.048) | (.048) | (.651) | (.396) | (.265) | (.438) | (.795) | (.466) | |
| Practice w AI | 6.19 | 6.36 | 6.08 | 6.08 | 5.98 | 5.53 | 7.44 | 6.60 | 5.64 | 5.74 |
| (.373) | (.379) | (.048) | (.049) | (.656) | (.395) | (.265) | (.438) | (.794) | (.466) | |
| See AI example | 7.83 | 8.04 | 7.72 | 7.72 | 7.58 | 7.23 | 8.38 | 8.18 | 8.34 | 7.02 |
| (.375) | (.380) | (.048) | (.049) | (.656) | (.397) | (.266) | (.439) | (.797) | (.467) | |
| Effect Sizes (d) — (SE) | ||||||||||
| Practice wo AI vs. Practice w AI | 1.22*** | 1.15*** | 1.19*** | 1.21*** | 1.14*** | .95*** | .96*** | .83*** | 1.06*** | .95*** |
| (.060) | (.059) | (.059) | (.060) | (.112) | (.059) | (.059) | (.058) | (.059) | (.059) | |
| Practice wo AI vs. See AI example | 2.58*** | 2.52*** | 2.54*** | 2.55*** | 2.50*** | 2.28*** | 2.07*** | 1.95*** | 2.12*** | 1.81*** |
| (.070) | (.070) | (.069) | (.070) | (.132) | (.067) | (.066) | (.065) | (.066) | (.063) | |
| Practice w AI vs. See AI example | 1.36*** | 1.37*** | 1.35*** | 1.35*** | 1.36*** | 1.33*** | 1.10*** | 1.12*** | 1.06*** | .86*** |
| (.061) | (.061) | (.060) | (.061) | (.113) | (.060) | (.059) | (.059) | (.059) | (.058) | |
| Note. GPT-4o is the main specification. Ex. Controls is the main specification, unadjusted for demographic and pretreatment variables, Ex. Cheaters excludes the 3% of participants who admitted to cheating on the test phase. LM to ER are disagregated scores for each of the five principles. LM = Less is More, ER = Easy Reading, EN = Easy Navigation, F = Formatting, ER = Easy Responding. *** p < .001, ** p < .01, * p < .05. | ||||||||||
| GPT-4o | Claude | Ex. Controls | Ex. Cheaters | Followup | LM | ER | EN | F | ER | |
| Means — (SE) | ||||||||||
| Practice wo AI | 5.39 | 5.28 | 4.55 | 4.55 | 4.71 | 4.39 | 7.01 | 6.19 | 4.26 | 5.13 |
| (.411) | (.426) | (.055) | (.055) | (.717) | (.391) | (.367) | (.479) | (.982) | (.496) | |
| Practice w AI | 5.82 | 5.84 | 5.00 | 5.00 | 5.00 | 4.69 | 7.11 | 6.61 | 5.29 | 5.38 |
| (.410) | (.426) | (.056) | (.056) | (.722) | (.390) | (.366) | (.478) | (.981) | (.495) | |
| See AI example | 5.87 | 5.95 | 5.02 | 5.03 | 5.08 | 4.66 | 7.08 | 6.78 | 5.47 | 5.36 |
| (.412) | (.427) | (.054) | (.054) | (.722) | (.392) | (.368) | (.480) | (.985) | (.497) | |
| Effect Sizes (d) — (SE) | ||||||||||
| Practice wo AI vs. Practice w AI | .32*** | .41*** | .32*** | .33*** | .22* | .24*** | .09 | .28*** | .33*** | .16** |
| (.057) | (.057) | (.056) | (.057) | (.106) | (.057) | (.057) | (.057) | (.057) | (.057) | |
| Practice wo AI vs. See AI example | .36*** | .49*** | .34*** | .35*** | .29** | .22*** | .06 | .39*** | .38*** | .14** |
| (.056) | (.056) | (.056) | (.057) | (.106) | (.056) | (.056) | (.056) | (.056) | (.056) | |
| Practice w AI vs. See AI example | .04 | .08 | .01 | .02 | .06 | -.03 | -.02 | .11* | .06 | -.01 |
| (.056) | (.056) | (.056) | (.057) | (.104) | (.056) | (.056) | (.056) | (.056) | (.056) | |
| Note. GPT-4o is the main specification. Ex. Controls is the main specification, unadjusted for demographic and pretreatment variables, Ex. Cheaters excludes the 3% of participants who admitted to cheating on the test phase. LM to ER are disagregated scores for each of the five principles. LM = Less is More, ER = Easy Reading, EN = Easy Navigation, F = Formatting, ER = Easy Responding. *** p < .001, ** p < .01, * p < .05. | ||||||||||
C3 Seeing AI examples was less effortful
Table S17 shows OLS models predicting practice effort metrics from practice condition. Results show that participants seeing an AI example expended considerably less effort, measured subjectively or objectively, through keystrokes or practice time, when compared both to participants who practiced with AI and without it. As in Study 2, participants who practiced with AI still expended less effort than those who practiced without it. As pre-registered, time is square-root-transformed, and keystrokes are log-transformed. Differences are slightly smaller when using untransformed variables.
| sqrt(Time) | log(Keystrokes) | Subjective Rating (0 - 10) | Time | Keystrokes | |
| Means — (SE) | |||||
| Practice wo AI | 2.83 | 5.01 | 6.17 | 9.00 | 259.34 |
| (.270) | (.541) | (.642) | (1.495) | (99.161) | |
| Practice w AI | 2.71 | 4.05 | 5.89 | 8.65 | 228.45 |
| (.270) | (.540) | (.641) | (1.493) | (99.050) | |
| See AI example | 1.85 | .81 | 5.52 | 4.99 | 24.98 |
| (.271) | (.542) | (.643) | (1.499) | (99.392) | |
| Effect Sizes (d) — (SE) | |||||
| Practice wo AI vs. Practice w AI | -.14* | -.55*** | -.14* | -.07 | -.10 |
| (.056) | (.056) | (.057) | (.056) | (.056) | |
| Practice wo AI vs. See AI example | -1.13*** | -2.41*** | -.32*** | -.83*** | -.73*** |
| (.059) | (.067) | (.056) | (.057) | (.056) | |
| Practice w AI vs. See AI example | -.99*** | -1.86*** | -.18** | -.76*** | -.64*** |
| (.058) | (.063) | (.056) | (.057) | (.056) | |
| Note. *** p < .001, ** p < .01, * p < .05. | |||||
Table S18 shows OLS models predicting test effort metrics from practice condition. Results show some differences: participants who had seen the AI example write for longer duing the test, and pressed more keys, however their subjective experience of effort was not different from those who practice with or without the AI tool.
| sqrt(Time) | log(Keystrokes) | Subjective Rating (0 - 10) | Time | Keystrokes | |
| Means — (SE) | |||||
| Practice wo AI | 2.45 | 5.40 | 7.05 | 6.14 | 432.01 |
| (.170) | (.562) | (.623) | (.655) | (109.856) | |
| Practice w AI | 2.50 | 5.52 | 7.19 | 6.35 | 466.90 |
| (.169) | (.561) | (.622) | (.654) | (109.733) | |
| See AI example | 2.57 | 5.86 | 7.21 | 6.63 | 517.26 |
| (.170) | (.563) | (.625) | (.656) | (110.112) | |
| Effect Sizes (d) — (SE) | |||||
| Practice wo AI vs. Practice w AI | .09 | .06 | .07 | .10 | .10 |
| (.057) | (.056) | (.057) | (.057) | (.056) | |
| Practice wo AI vs. See AI example | .22*** | .25*** | .08 | .23*** | .24*** |
| (.056) | (.055) | (.056) | (.056) | (.055) | |
| Practice w AI vs. See AI example | .13* | .19*** | .01 | .13* | .14* |
| (.056) | (.056) | (.056) | (.056) | (.055) | |
| Note. *** p < .001, ** p < .01, * p < .05. | |||||
Table S19 shows OLS models predicting learning rate metrics from practice condition. Learning rate is defined as the difference between test and pretest, divided by the effort metric. It shows how many points (10 point scale) the participant improved per unit effort (e.g., per minute spent practicing). Participants who had seen an AI example improved their skill more efficiently.
| sqrt(Time) | log(Keystrokes) | Subjective Rating (0 - 10) | Time | Keystrokes | |
| Means — (SE) | |||||
| Practice wo AI | .20 | .15 | .18 | .31 | .34 |
| (.155) | (.272) | (.085) | (.278) | (.167) | |
| Practice w AI | .30 | .22 | .25 | .45 | .49 |
| (.155) | (.272) | (.085) | (.277) | (.167) | |
| See AI example | .51 | .97 | .28 | 1.08 | .62 |
| (.155) | (.273) | (.085) | (.278) | (.168) | |
| Effect Sizes (d) — (SE) | |||||
| Practice wo AI vs. Practice w AI | .21*** | .08 | .25*** | .15** | .27*** |
| (.057) | (.057) | (.057) | (.057) | (.057) | |
| Practice wo AI vs. See AI example | .62*** | .94*** | .38*** | .86*** | .53*** |
| (.057) | (.058) | (.056) | (.058) | (.057) | |
| Practice w AI vs. See AI example | .41*** | .86*** | .13* | .71*** | .25*** |
| (.057) | (.058) | (.056) | (.057) | (.056) | |
| Note. *** p < .001, ** p < .01, * p < .05. | |||||
C4 Seeing an AI example did not discourage motivation for future learning
Table S20 presents differences in perceived learning, perceived writing skill, and the likelihood of asking for feedback across conditions, with effect sizes and means reported for each comparison. Despite objectively learning more, participants who practiced with AI and saw an AI example perceived their learning and skill levels to be similar to those who practiced without AI and asked for feedback at comparable rates.
| Perceived learning | Perceived writing skill | Asked for feedback | |
| Means — (SE) | |||
| Practice wo AI | 5.26 | 6.33 | .64 |
| (.549) | (.517) | (.670) | |
| Practice w AI | 5.25 | 6.36 | .46 |
| (.549) | (.516) | (.669) | |
| See AI example | 5.42 | 6.23 | .55 |
| (.551) | (.518) | (.671) | |
| extbfEffect Sizes (d) | |||
| Practice wo AI vs. Practice w AI | -.01 | .02 | 1.19 |
| (.057) | (.057) | (.122) | |
| Practice wo AI vs. See AI example | .09 | -.06 | 1.10 |
| (.056) | (.056) | (.121) | |
| Practice w AI vs. See AI example | .09 | -.08 | 0.921 |
| (.056) | (.056) | (.120) | |
| Note. *** p < .001, ** p < .01, * p < .05. | |||
C5 The effects of seeing an AI example persist
Table S21 shows means and standardized differences for measures of writing skill and related outcomes during the follow-up phase. The main specification demonstrates that participants who practiced with AI continued to outperform those who did not practice or practiced without AI. Robustness checks, including using a different language model (Column 2), excluding control variables (Column 3), and removing participants who admitted to cheating (Column 4) confirm the consistency of these effects. The results also hold when evaluating each of the five principles separately (Columns 5–9). These findings suggest that the benefits of practicing with AI are durable and persist even after participants stop using the tool.
| GPT-4o | Claude | Ex. Controls | Ex. Cheaters | LM | ER | EN | F | ER | |
| Means — (SE) | |||||||||
| Practice wo AI | 4.95 | 5.10 | 4.87 | 4.88 | 5.32 | 6.83 | 5.40 | 1.98 | 5.23 |
| (.776) | (.798) | (.109) | (.110) | (.750) | (.730) | (.847) | (1.829) | (.905) | |
| Practice w AI | 5.37 | 5.67 | 5.34 | 5.38 | 5.57 | 6.98 | 5.76 | 2.91 | 5.61 |
| (.781) | (.804) | (.103) | (.105) | (.756) | (.735) | (.853) | (1.842) | (.912) | |
| See AI example | 5.40 | 5.71 | 5.37 | 5.36 | 5.54 | 6.98 | 5.87 | 3.14 | 5.46 |
| (.781) | (.804) | (.103) | (.105) | (.756) | (.735) | (.854) | (1.843) | (.912) | |
| Effect Sizes (d) — (SE) | |||||||||
| Practice wo AI vs. Practice w AI | .29** | .40*** | .32** | .34** | .18 | .11 | .24* | .28** | .23* |
| (.106) | (.107) | (.104) | (.106) | (.106) | (.106) | (.106) | (.106) | (.106) | |
| Practice wo AI vs. See AI example | .32** | .43*** | .35*** | .33** | .16 | .11 | .31** | .35*** | .14 |
| (.106) | (.107) | (.104) | (.106) | (.106) | (.106) | (.106) | (.107) | (.106) | |
| Practice w AI vs. See AI example | .02 | .03 | .02 | -.01 | -.02 | -.00 | .07 | .07 | -.09 |
| (.104) | (.104) | (.101) | (.103) | (.104) | (.104) | (.104) | (.104) | (.104) | |
| Note. GPT-4o is the main specification. Ex. Controls is the main specification, unadjusted for demographic and pretreatment variables, Ex. Cheaters excludes the 3% of participants who admitted to cheating on the test phase. LM to ER are disagregated scores for each of the five principles. LM = Less is More, ER = Easy Reading, EN = Easy Navigation, F = Formatting, ER = Easy Responding. *** p < .001, ** p < .01, * p < .05. | |||||||||
The follow-up analyses pool three separate follow-up samples collected on consecutive days. Table S22 are the results for each of these samples separately.
| Overall | Batch 1 | Batch 2 | Batch 3 | |
| Means — (SE) | ||||
| Practice wo AI | 4.95 | 6.02 | 5.47 | 4.85 |
| (.776) | (1.539) | (.492) | (.811) | |
| Practice w AI | 5.37 | 6.64 | 5.57 | 5.41 |
| (.781) | (1.446) | (.483) | (.817) | |
| See AI example | 5.40 | 6.51 | 5.88 | 5.32 |
| (.781) | (1.502) | (.472) | (.818) | |
| Effect Sizes (d) — (SE) | ||||
| Practice wo AI vs. Practice w AI | .29** | .43 | .07 | .40** |
| (.106) | (.387) | (.187) | (.147) | |
| Practice wo AI vs. See AI example | .32** | .34 | .29 | .34* |
| (.106) | (.354) | (.190) | (.149) | |
| Practice w AI vs. See AI example | .02 | -.09 | .22 | -.06 |
| (.104) | (.344) | (.177) | (.145) | |
| Note. GPT-4o is the main specification. Ex. Controls is the main specification, unadjusted for demographic and pretreatment variables, Ex. Cheaters excludes the 3% of participants who admitted to cheating on the test phase. LM to ER are disagregated scores for each of the five principles. LM = Less is More, ER = Easy Reading, EN = Easy Navigation, F = Formatting, ER = Easy Responding. *** p < .001, ** p < .01, * p < .05. | ||||
C6 Seeing AI examples was equally effective across subgroups
As in Study 2, we tested whether each of the pretreatment demographic variables moderated the effects of seeing an AI example. To do this, we ran separate linear in which writing skill during the test phase was regressed on condition, the pre-treatment moderator of interest, writing skill at baseline, and an interaction term between the moderator condition. After correcting the p-values for the interaction terms, none were significant at the .05 level, suggesting that seeing AI examples was equally effective across groups.
| Test | Follow-Up | Time Practice | Keys Practice | Effort Practice | Per. Learning | Per. Skill | Want Feedback | |||||||||
| Level | PAI | AIE | PAI | AIE | PAI | AIE | PAI | AIE | PAI | AIE | PAI | AIE | PAI | AIE | PAI | AIE |
| Continuous Moderators | ||||||||||||||||
| Pretest | 0.565 | 0.546 | 0.708 | 0.945 | 0.987 | 0.857 | 0.987 | 0.940 | 0.987 | 0.405 | 0.987 | 0.967 | 0.987 | 0.576 | 0.565 | 0.274 |
| YOB | 0.961 | 0.987 | 0.967 | 0.857 | 0.565 | 0.855 | 0.405 | 0.405 | 0.940 | 0.763 | 0.516 | 0.724 | 0.987 | 0.724 | 0.763 | 0.871 |
| Writing Skill | 0.943 | 1.000 | 0.405 | 0.987 | 0.878 | 0.987 | 0.967 | 0.871 | 0.763 | 0.707 | 0.986 | 0.405 | 0.900 | 0.987 | 0.434 | 0.405 |
| Gender | ||||||||||||||||
| Male | 0.816 | 0.987 | 0.987 | 0.987 | 0.987 | 0.532 | 0.793 | 0.535 | 0.565 | 0.405 | 0.450 | 0.703 | 0.724 | 0.535 | 0.707 | 0.565 |
| Other | 0.446 | 0.565 | 0.900 | 0.987 | 0.405 | 0.724 | 0.842 | 0.728 | 0.426 | 0.791 | 0.655 | 0.987 | 0.987 | 0.763 | 0.793 | 0.791 |
| Race | ||||||||||||||||
| White | 0.822 | 0.724 | 0.499 | 0.605 | 0.811 | 0.987 | 0.967 | 0.855 | 0.405 | 0.937 | 0.987 | 0.763 | 0.987 | 0.533 | 0.987 | 0.797 |
| Black | 0.718 | 0.405 | 0.791 | 0.718 | 0.987 | 0.901 | 0.987 | 0.899 | 0.274 | 0.565 | 0.760 | 0.945 | 0.987 | 0.499 | 0.793 | 0.446 |
| Asian | 0.940 | 0.766 | 0.734 | 0.532 | 0.703 | 0.405 | 0.934 | 0.605 | 0.341 | 0.565 | 0.937 | 0.987 | 0.989 | 0.280 | 0.535 | 0.565 |
| Latino | 0.987 | 0.899 | 0.405 | 0.405 | 0.763 | 0.987 | 0.405 | 0.816 | 0.405 | 0.940 | 1.000 | 0.987 | 0.857 | 0.624 | 0.987 | 0.991 |
| Other | 0.405 | 0.855 | 0.987 | 0.855 | 0.987 | 0.987 | 0.987 | 0.987 | ||||||||
| Education Level | ||||||||||||||||
| High School Graduate | 0.940 | 0.900 | 0.749 | 0.565 | 0.405 | 0.938 | 0.718 | 0.987 | 0.405 | 0.405 | 0.987 | 0.967 | 0.987 | 0.763 | 0.987 | 0.734 |
| Some College | 0.987 | 0.987 | 0.791 | 0.565 | 0.405 | 0.900 | 0.707 | 0.986 | 0.405 | 0.405 | 1.000 | 0.987 | 0.987 | 0.878 | 0.987 | 0.987 |
| Associate Degree | 0.900 | 0.987 | 0.791 | 0.565 | 0.447 | 0.855 | 0.707 | 0.940 | 0.446 | 0.520 | 0.987 | 0.987 | 0.989 | 0.811 | 0.987 | 0.814 |
| Bachelor’s Degree | 0.986 | 0.987 | 0.793 | 0.499 | 0.405 | 0.899 | 0.707 | 0.967 | 0.405 | 0.405 | 1.000 | 0.987 | 0.989 | 0.791 | 0.987 | 0.964 |
| Master’s Degree | 0.986 | 0.987 | 0.724 | 0.405 | 0.405 | 0.940 | 0.724 | 0.964 | 0.405 | 0.434 | 0.987 | 0.987 | 0.987 | 0.763 | 0.987 | 0.987 |
| Doctoral Degree | 0.987 | 0.987 | 0.524 | 0.724 | 0.516 | 0.832 | 0.987 | 0.940 | 0.406 | 0.516 | 0.987 | 0.987 | 0.832 | 0.760 | 0.987 | 0.964 |
| Professional Degree | 0.987 | 0.900 | 0.987 | 0.943 | 0.763 | 0.987 | 0.987 | 0.987 | 0.707 | 0.405 | 0.763 | 0.938 | 0.763 | 0.749 | 0.987 | 0.987 |
| Motivation | ||||||||||||||||
| Hardly Motivated | 0.987 | 0.900 | 0.624 | 0.899 | 0.768 | 0.763 | 0.832 | 0.987 | 0.763 | 0.900 | 0.685 | 0.707 | 0.760 | 0.535 | 0.405 | 0.565 |
| Somewhat Motivated | 0.977 | 0.535 | 0.763 | 0.763 | 0.871 | 0.585 | 0.937 | 0.987 | 0.763 | 0.791 | 0.791 | 0.763 | 0.797 | 0.707 | 0.536 | 0.763 |
| Very Motivated | 0.907 | 0.705 | 0.778 | 0.703 | 0.986 | 0.405 | 0.797 | 0.987 | 0.797 | 0.708 | 0.847 | 0.857 | 0.847 | 0.752 | 0.451 | 0.749 |
| Extremely Motivated | 0.855 | 0.536 | 0.724 | 0.707 | 0.763 | 0.724 | 0.763 | 0.987 | 0.987 | 0.763 | 0.987 | 0.900 | 0.763 | 0.763 | 0.535 | 0.535 |
| Experience with AI Writing Assistants | ||||||||||||||||
| Hardly Ever Use Them | 0.724 | 0.763 | 0.907 | 0.791 | 0.405 | 0.900 | 0.405 | 0.987 | 0.967 | 0.446 | 0.565 | 0.987 | 0.987 | 0.724 | 0.987 | 0.405 |
| Use a Few Times Per Week | 0.576 | 0.987 | 0.763 | 0.763 | 0.766 | 0.766 | 0.763 | 0.987 | 0.565 | 0.152 | 0.903 | 0.763 | 0.783 | 0.405 | 0.945 | 0.405 |
| Use About Once a Week | 0.846 | 0.707 | 0.623 | 0.786 | 0.406 | 0.855 | 0.405 | 0.987 | 0.763 | 0.718 | 0.967 | 0.987 | 0.899 | 0.405 | 0.707 | 0.797 |
| Use Every Day | 0.987 | 0.724 | 0.987 | 0.763 | 0.341 | 0.763 | 0.405 | 0.987 | 0.405 | 0.280 | 0.943 | 0.987 | 0.899 | 0.763 | 0.734 | 0.987 |
| Note. Models for test and follow-up performance, square-root practice time, log keystrokes, subjective effort, perceived learning and perceived writing skill or OLS models. Asking to see feedback was a binary Yes/No variable, and was modelled with logistic regression. Models match the pre-registered main specification, and thus control for all other pre-treatment variables. Per. = Perceived, PAI = Practice with AI, AIE = See AI example. | ||||||||||||||||

### v3

@@
-Learning not cheating: AI assistance can enhance rather than hinder skill development
+Coach not crutch: AI assistance can enhance rather than hinder skill development
 Abstract
-It is widely believed that outsourcing cognitive work to AI boosts immediate productivity at the expense of long-term human capital development.
-An overlooked possibility is that AI tools can support skill development by providing just-in-time, high-quality, personalized examples.
-In this investigation, lay forecasters predicted that practicing writing cover letters with an AI tool would impair learning compared to practicing writing letters without the tool.
-However, in a highly-powered pre-registered experiment, participants randomly assigned to practice writing with AI improved more on a writing test one day later compared to writers assigned to practice without AI.
-Notably, writers given access to the AI tool improved more despite exerting less effort, whether measured by time on task, keystrokes, or subjective ratings.
-We replicated and extended these results in a second pre-registered experiment, showing that writers given access to the AI tool again outperformed those who practiced on their own—but performed no better than writers merely shown an AI-generated cover letter that they could not edit.
-Collectively, these findings constitute an existence proof that by providing personalized examples of high-quality work, AI tools can improve, rather than undermine, learning.
+Does using AI make you stupid?
+Contrary to predictions by lay forecasters, participants randomly assigned to practice writing with access to an AI tool improved more on a writing test one day later compared to writers assigned to practice without AI—despite exerting less effort.
+Likewise, whereas forecasters predicted that AI would be less helpful, and less worth paying for, than personalized feedback from experienced human editors, we found the converse: practicing with AI increased skill more than feedback from human editors (or simply googling examples and tips).
+The results of a third experiment suggest that AI may teach by example: writers merely shown an AI-generated cover letter (without the opportunity to practice) performed as well as writers who practiced writing with the original AI tool.
+Collectively, the findings of these pre-registered studies constitute an existence proof that by providing personalized, high-quality examples, AI can improve, rather than undermine, learning.
 keywords:
@@
 Nevertheless, there is growing concern that AI tools will be used as a crutch, providing immediate gains in performance at the expense of long-run development of human capital (5, 6, 7).
-For instance, in a 2024 poll, 62% of surveyed adults predicted that Generative AI will “lead to humans becoming less intelligent” (8).
-In January 2023, New York City public schools banned ChatGPT, citing “concerns about negative impacts on student learning (9).” When this ban was lifted three months later, it was not because of the potential of AI to scaffold learning, but instead because of the “reality that students are participating in and will work in a world where understanding Generative AI is crucial (10).” The sentiment behind the initial ban aligns with teacher perceptions: in a nationally representative poll in May 2024, four times as many K-12 educators judged the use of AI tools as net harmful (24%) than net beneficial (6%) (11).
+For instance, in a 2024 poll, 62% of surveyed adults predicted that Generative AI would “lead to humans becoming less intelligent” (8).
+In January 2023, New York City public schools banned ChatGPT, citing “concerns about negative impacts on student learning (9).” When this ban was lifted three months later, it was not because of AI’s potential to scaffold learning, but instead because of the “reality that students are participating in and will work in a world where understanding Generative AI is crucial (10).” However, teachers remain worried: in a nationally representative poll in May 2024, four times as many K-12 educators judged the use of AI tools as net harmful (24%) than net beneficial (6%) (11).
 Concerns that using AI tools hinders learning (while increasing short-term performance) are justified for at least three reasons.
 First, AI systems based on large language models like GPT-4 have been shown to confidently assert erroneous facts (i.e., hallucinations (12)), make reasoning and arithmetic errors (13), and complete other tasks with varying degrees of accuracy.
-Second, regardless of accuracy, the fluent and instantaneous solutions AI tools generate may contribute to an illusion of mastery.
+Second, the fluent and instantaneous solutions AI tools generate may contribute to an illusion of mastery.
 To the extent users conflate the skills of an AI tool with their own, they may be less likely to seek feedback and improve.
 Prior research has found that searching for information on the Internet, for example, creates an illusion whereby people conflate knowledge outside their heads with what they personally know themselves (14).
-Third, technological tools reduce the need for the learner to be cognitively engaged with the task at hand.
-For instance, knowing that we will be able to search for a fact on a computer has been shown to reduce memory for that fact, instead encouraging recall of how to search for it (15).
+Third, technological tools reduce the learner’s need to be cognitively engaged with the task at hand.
+For instance, knowing that one can search for a fact on a computer has been shown to reduce memory for that fact, instead encouraging recall of how to search for it (15).
 And drivers who use GPS tend to have worse hippocampal-dependent spatial memory, both cross-sectionally and longitudinally (16, 17).
-To the extent that tools powered by generative AI instantaneously produce turn-key solutions for complex cognitive tasks, they may be especially detrimental to learning.
+When AI-powered tools instantaneously produce turn-key solutions for complex cognitive tasks, they may be especially detrimental to learning.
 It is, after all, tempting to copy and paste the output of an AI tool without even laying eyes on it.
-There is a plausible, albeit less obvious, alternate hypothesis.
-Using AI tools to help do our work could help us develop our own skills.
-In particular, the current generation of AI tools may teach by example, offering high-quality and personally tailored demonstrations of abstract principles that are otherwise difficult to grasp and apply.
+On the other hand, using AI tools to assist us in our work may, in fact, help us develop our own skills.
+In particular, the current generation of AI tools may teach by example, offering high-quality and personally tailored demonstrations of abstract principles that are otherwise difficult to grasp.
 Classic research shows that worked examples of math problems (i.e., not just answers but the step-by-step process by which problems are solved) scaffold learning more effectively than explanations alone (18, 19).
-Compared to textbooks, conventional computer tutoring programs (e.g., (20)), and even human teachers, today’s AI tools are unprecedented in their ability to deliver hyper-personalized high-quality examples on command.
-Thus, AI tools may improve skill development if the benefits of exposure to just-in-time examples tailored to learners’ needs outweigh the costs of diminished engagement.
-There is little research on how AI tools influence skill development.
-In working papers, results have been mixed.
-Some studies have found that interacting with AI tools improves skill on subsequent tests in which AI tools are not available (21, 22), while others have shown null or even negative effects (23, 24, 22).
-Notably, these studies examine AI tutors, chatbots, or explanations explicitly designed to support learning, rather than simply providing solutions as is typical in real-world use.
-Further, they focus exclusively on mathematics and computer programming.
-In this investigation, we ask whether AI tools can increase intermediate-term skill development, above and beyond just improving performance while using the AI tool.
+It seems likely that, compared to textbooks, conventional computer tutoring programs (e.g., (20)), and human teachers, AI tools may provide higher quality, just-in-time examples exquisitely tailored to a learner’s specific needs.
+Indeed, it’s possible that in both quality and didactic utility, AI examples (e.g., of an excellent cover letter) might surpass even those provided by domain experts (e.g., professional writers and editors).
+Thus, if the upside of exposure to excellent examples tailored to the learners’ needs outweighs the downside of diminished engagement, AI tools may support skill development.
+The scant research on how AI tools influence skill development has yielded mixed results.
+Some studies report that interacting with AI tools improves skill on subsequent tests in which AI tools are not available (21, 22), while others have shown null or even negative effects (23, 24, 22).
+Notably, these studies examine AI tutors, chatbots, or explanations explicitly designed to support learning, rather than simply providing solutions, as is typical in real-world use.
+Further, they focused exclusively on mathematics and computer programming.
+In this investigation, we ask whether AI tools can support skill development, above and beyond merely improving performance while using the AI tool.
 We focus on writing—the most common use of AI at work, as ranked in a nationally representative survey of American adults in August 2024 (26).
-Participants in our three pre-registered studies were American adults on the survey platform Prolific.
-To differentiate the effects of AI use on learning versus performance, we developed a paradigm in which all participants were given a baseline writing test (i.e., revising a poorly written cover letter).
-They were then introduced to evidence-based strategies for professional writing, with descriptions and examples for each one (26, 27).
-Next, participants were randomly assigned to one of three conditions: practicing rewriting a different cover letter with access to an AI tool based on the same writing principles that users had just learned (See Figure 1), practicing without this tool, or a no-practice control group.
-To assess gains in writing skill, participants completed an incentivized test in which they rewrote yet another cover letter without access to AI, with a cash bonus guaranteed for submissions ranked in the top 10 percent.
-To assess the persistence of skill improvement, all participants were invited to complete a similar incentivized test of writing skill one day later.
-See Figure 2.
-We used GPT4o to rate each cover letter for each of the five writing principles introduced in this experiment.
-We averaged these ratings to produce summary scores of writing skill and, in a random subsample (n = 30), validated these scores using trained human raters (r = .83, p < .001).
-Additionally, we asked a separate sample of participants to read pairs of test-phase cover letters randomly selected from different conditions, and to indicate which letter would be more likely to secure a job interview.
-The relative likelihood of a cover letter securing an interview correlated positively with AI ratings of writing skill (rs = .29 and .28, ps < .001, for Study 2 and 3, respectively).
-In Study 1, forecasters presented with this design were twice as likely to predict that practicing writing with the assistance of the AI tool would impair learning compared to practicing without the AI tool.
-In Study 2, however, participants who had practiced with the AI tool learned more (i.e., wrote better cover letters during the test phase) compared to either comparison group—an advantage that persisted in a one-day follow-up test.
-Finally, in Study 3, we explored the mechanism for these learning gains by introducing an example-only condition.
-Participants who had merely seen an AI-generated example (but did not have an opportunity to practice) improved in writing skill as much as participants who had practiced with an AI tool; benefits again persisted in a one-day follow-up test.
-Test-phase cover letters written by participants who had practiced with AI (Studies 2 and 3) or had seen an AI example (Study 3) were more likely to secure hypothetical job interviews compared to cover letters written by participants who had not practiced (Study 2) or had practiced without AI (Studies 2 and 3).
-Study 1
-Lay forecasters predicted that practicing with AI would hinder learning
+In our experimental paradigm all participants are given a baseline writing test (i.e., revising a poorly written cover letter), followed by a short lesson on five evidence-based strategies for professional writing (25, 27).
+Next, participants are randomly assigned to different types of practice (e.g., practice with access to an AI tool, practice without access to an AI tool, practice with feedback from professional editors, and practice with Google Search to find examples and tips).
+Finally, all participants complete an incentivized test in which they rewrite cover letters without access to AI.
+In Study 1, forecasters presented with this design were twice as likely to predict that practicing writing with the assistance of AI would impair learning compared to practicing without it.
+In Study 2, however, participants who practiced with AI actually learned more (i.e., wrote better cover letters during the test phase) than participants who practiced without AI—an advantage that persisted in a one-day follow-up test.
+In Study 3, forecasters predicted that feedback from AI would be less helpful than feedback from professional editors and, accordingly, were less willing to pay for it.
+However, in Study 4, participants who practiced with AI learned more than participants who practiced with feedback from professional editors (or those who practiced with Google Search to find examples and tips).
+Finally, in Study 5, we explored the mechanism for these learning gains by introducing an example-only condition.
+Participants who saw an AI-generated example (but did not have an opportunity to practice) improved their writing as much as participants who practiced with AI; benefits again persisted in a one-day follow-up test.
+Study 1: Forecasting the effect of AI on learning
 We showed N = 150 participants screenshots of a random assignment study with three conditions.
 We asked them to rank-order these conditions according to how much they predicted future participants would learn in each.
+Lay forecasters predicted that practicing with AI would hinder learning
 Confirming our pre-registered hypothesis, nearly twice as many forecasters (64.7%) ranked practicing alone above practicing writing with access to an AI tool as the converse (35.3%, (1) = 12.9, p < .001), see Figure 3.
-Participants made this prediction regardless of self-reported experience with AI (OR = 0.83, p = .239) or any other measured demographic characteristic (ps > .05).
-In open-ended responses, forecasters who were pessimistic about the effect of the AI tool on learning speculated that it would crowd out effort (e.g., “Practicing alone would force more recall and problem-solving skills, while AI essentially gives the answer for them.”, “I think oftentimes using AI impedes the learning process because it’s the ‘easy way.”’).
-Those with positive views, on the other hand, cited the possibility of AI providing insights or examples that would be otherwise unavailable (“As much as I hate AI, I do not believe you can improve in any manner if you do not have examples or ways of learning, and AI can provide this.”)
-Study 2
+Participants made this prediction regardless of self-reported experience with AI (OR = 0.72, p = .096) or any other measured demographic characteristic (ps > .05).
+See Table S2
+In open-ended responses, forecasters who were pessimistic about the effect of the AI tool on learning speculated that it would crowd out effort (e.g., “Practicing alone would force more recall and problem-solving skills, while AI essentially just gives them the answer.”, “I think oftentimes using AI impedes the learning process because it’s the ‘easy way.”’).
+Those with positive views, on the other hand, cited the possibility of AI providing insights or examples that would be otherwise unavailable (“As much as I hate AI, I do not believe you can improve in any manner if you do not have examples or other ways of learning, and AI can provide this.”)
+Study 2: Effect of practicing with AI on learning
 Study 2 tested whether the predictions of Study 1 forecasters were accurate.
 Specifically, N = 2,238 participants completed a baseline questionnaire and pretest (rewriting a poorly written cover letter), followed by a lesson introducing five principles of effective writing (i.e., Less is more, Make reading easy, Design for easy navigation, Use enough formatting but no more, Make responding easy) (25).
-Next, participants were randomly assigned to one of three practice conditions: (1) rewriting a new cover letter with an AI writing tool that revised text instantly based on these principles, (2) rewriting the new cover letter without the AI tool, or (3) a no practice control.
-At the end of the session, all participants completed a test of writing skill (rewriting a yet another cover letter without access to the AI writing tool) and an exit questionnaire.
+Next, participants were randomly assigned to one of three practice conditions: (1) rewriting a new cover letter with an AI writing tool that instantly revised text based on these principles, (2) rewriting the new cover letter without the AI tool, or (3) a no-practice control.
+At the end of the session, all participants completed a test of writing skill (rewriting a new cover letter without access to the AI writing tool) and an exit questionnaire.
 One day later, all participants were invited to complete a similar incentivized test of writing skill.
+See Figure 2.
+We assessed writing skill in two ways.
+First, we used GPT-4o to rate cover letters on each of the principles of effective writing and then averaged these ratings to produce a summary score (Cronbach’s = .81).
+This operationalization demonstrated convergent validity with ratings by human research assistants who rated a random subsample of n = 100 cover letters on the same principles (r = .70, p < .001).
+Second, we recruited participants naive to this investigation to read two letters and select the one more likely to secure its writer a job interview.
+These two operationalizations converged: preferred cover letters obtained higher AI ratings of writing skill (rs = .29).
 AI practice improved writing skill
-Consistent with other studies demonstrating the productivity benefits of AI tools (3, 2), participants given access to the AI writing tool produced cover letters during the practice phase that were dramatically higher in quality than participants who were not (d = 1.01, p <.001).
-The learning advantage of having practiced with AI was evident in the test phase: Consistent with our pre-registered hypothesis, participants who had practiced with the AI tool produced higher-quality writing than did participants who either had practiced without the AI tool (d = .38, p <.001) or who had not practiced at all (d = .47, p <.001).
+Consistent with other studies demonstrating the productivity benefits of AI tools (3, 2), participants given access to the AI writing tool produced cover letters during the practice phase that were dramatically higher in quality than participants without access (d = 1.01, p <.001).
+The learning advantage of having practiced with AI was evident in the test phase: Consistent with our preregistered hypothesis, participants who had practiced with the AI tool produced higher-quality writing than did participants who either had practiced without the AI tool (d = .38, p <.001) or who had not practiced at all (d = .47, p <.001).
 See Figure S3.
@@
 .47, p < .001).
-See Figure 5.
-Notably, the effect of practicing with AI was not limited to stronger or weaker writers—participants across all skill levels benefited equally from practicing with AI.
+See Figure 4B.
 AI practice was less effortful
-The learning benefits of using an AI writing tool were evident despite reduced effort during the practice phase.
 Compared to participants who practiced alone, participants who practiced with the AI tool spent 0.44 fewer minutes during the practice phase (3.73 vs.
-4.17; d = -.12, p = .025), logged roughly a quarter as many keystrokes (26% d = -.44, p <.001) and self-reported expending less effort during practice (d = -.31, p <.001).
+4.17; d = .12, p = .025, quantity of effort), logged roughly a quarter as many total keystrokes (26% d = .44, p <.001) and logged 65% fewer keystrokes per minute (d = .22, p <.001, intensity or quality of effort), and self-reported expending less effort during practice (d = .31, p <.001).
 Nevertheless, it would be inaccurate to label writers practicing with AI as entirely disengaged.
-Copying, pasting, and submitting the AI tool’s output could be accomplished almost instantly.
-Yet, the majority of participants chose to interact with the task for over 3 minutes, and 95% made at least one edit to the AI-generated output before final submission.
+Rather than copying and pasting the AI tool’s output, the majority of participants practicing with AI chose to interact with the task for over 3 minutes, and 95% made at least one edit to the AI-generated output before final submission.
 See Figure S4 in Supplementary Information for details.
-Differential effort during practice raised the possibility that participants who had practiced with the AI tool outperformed those who had practiced on their own because they were less fatigued during the test phase.
-However, participants who had practiced with AI did not spend more time as those who had practiced without AI (d = .06, p = .235), but logged more keystrokes (d = .12, p = .026), and self-reported expending less effort (d = -.12, p = .023) during the test phase.
 AI practice did not create the illusion of mastery
-Following the test phase, there were no differences by condition on self-reported knowledge or motivation to improve.
-Despite improving more in objectively assessed writing skill, participants who had practiced with AI reported having learned no more than those who had either practiced alone or done no practice at all (ps >.05).
-Self-ratings of writing skill after the practice phase were also indistinguishable between participants who practiced with AI and those who practiced without the AI tool, or in the no-practice control group (ps >.05), but participants who had practiced without AI rated their skill more highly than those who did not practice (d = .14 p = .008).
-Finally, compared to participants who did not practice, participants who had practiced with AI were slightly less likely to request feedback after the test phase (.65 vs.
-.60, p = .039), but just as likely as participants who had practiced without AI (.64 vs.
-.60, p = .167).
-See Section B4 in the Supplementary Information for details.
-The effects of practicing with AI persist
+In the exit questionnaire, participants who had practiced with AI reported learning about as much as participants who had practiced without it or not practiced.
+Likewise, they judged their writing skill similarly to those who had either practiced alone or done no practice at all (|ds| 0.10, ps >.05).
+When asked if they wanted to see additional feedback, they were about as likely to say yes as participants who had practiced without AI (.64 vs.
+.60, p = .167), and slightly less likely than those who had not practice at all (.65 vs.
+.60, p = .039).
+See Section C4 in the Supplementary Information for details.
+The benefits of practicing with AI were just as large a day later
 To examine whether the treatment effects persisted over time, we re-contacted all participants one day later.
-The majority of participants responded (87%), and attrition rates did not differ by condition (13% to 14% = .68, p .710).
-Confirming our pre-registered hypothesis, participants who had practiced with the AI tool to practice the previous day continued to outperform those who had practiced without the tool (d = .41, p <.001) as well as those who had not practiced at all (d = .46, p <.001).
+Most participants responded (87%), and attrition rates did not differ by condition (13% to 14% = .68, p = .710).
+Confirming our pre-registered hypothesis, participants who had practiced with the AI tool the previous day continued to outperform those who had practiced without the tool (d = .41, p <.001) as well as those who had not practiced at all (d = .46, p < .001).
 Participants who had practiced without the AI tool performed no better than those who did not practice (d = .05, p = .331).
-See Figure 4 and Section B5 of Supplementary Information for details.
+The magnitude of the follow-up effect size was not significantly different from the immediate test phase effect size (condition time interactions were non-significant (ps > .69).
+See Figure 4A and Section C5 of Supplementary Information for details.
+Results were not moderated by individual differences
 None of the findings above were moderated by individual difference variables, including past experience with AI, age, gender, race, education, motivation to learn, and baseline writing skill, BH-corrected p-values >.05.
 See Table S13 in the Supplementary Information for details.
-Study 3
-To gain insight into what might be driving the benefit of practicing with AI, in Study 3 (N = 2,003), we preregistered a replication and extension in which we replaced the no-practice condition of Study 2 with an example-only condition.
-In this condition, participants were shown an AI-generated writing example that they could not edit.
+Study 3: Willingness to pay for AI vs. human editors
+Given unwarranted pessimism about the pedagogical power of AI, in Study 3 we examined whether people underinvest in AI as a learning resource.
+Specifically, we asked N = 150 people to imagine they were participants in Study 2 and asked them how much they would be willing to pay for feedback from the AI tool vs.
+feedback from experienced human editors.
+Participants were willing to pay more for feedback from expert human editors than for AI feedback
+Out of a maximum of a dollar (the total amount of the hypothetical bonus participants could earn), participants on average were willing to pay 28 cents for the AI tool vs.
+38 cents for feedback from an experienced human editor (t(149) = 4.83, p <.001, d = 0.39).
+Participants predicted that editorial feedback would be more beneficial
+Participants predicted that experienced human editors (M = 6.68 on a 0—10 Likert scale) would be more effective in helping them improve their writing, compared to feedback from an AI tool (M = 5.64; t(148) = 4.77, p <.001, d = 0.39).
+Predicted effectiveness mediated underinvestment in AI feedback
+People’s beliefs about how effective AI and human editor feedback would be predicted willingness to pay (b = 0.06, SE = 0.01, p < .001) and accounted for most of the preference for human editors (indirect effect = 0.06, p < .001).
+Once these predictions were accounted for, the remaining preference for human feedback was reduced by 60% (c’ = 0.04, SE = 0.02, p = .054).
+See Figure 5 and Section D1 of Supplementary Information for details.
+Study 4: Practicing with AI vs. with expert feedback (vs. with Google Search)
+Participants (N = 2,997) completed a baseline questionnaire and pretest, followed by the same lesson on effective writing as in Study 2, and then practiced rewriting a new cover letter on their own.
+Next, participants were randomly assigned to one of three conditions: (1) practice with AI, (2) practice with feedback from professional human editors, (3) practice with Google Search.
+Letters written by participants assigned to the second condition were given to professional editors averaging 25 years of professional editorial and writing experience.
+These 49 editors, recruited in partnership with the Journalists’ Resource at Harvard Kennedy School, provided personalized revisions and constructive feedback.
+The following day, all participants attempted to improve the cover letter they had previously submitted, using the resources afforded by their assigned condition.
+Finally, all participants completed a writing test (rewriting a cover letter without access to AI) and an exit questionnaire.
+Practicing with AI improved writing skill more than feedback from professional editors and Google Search
+During the practice phase, participants working with AI outperformed participants who received personalized feedback from professional editors (d = 0.76, p <.001) and those assigned to Google cover letter examples and tips (d = 1.03, p <.001).
+During the test phase, in which participants were asked to write a new cover letter without access to any outside resources, participants who had practiced with the AI tool produced higher-quality cover letters than those who had practiced with Google Search (d = 0.46, p <.001), and even outperformed those who had practiced with personalized feedback from professional editors (d = 0.20, p <.001).
+See Figure 6A.
+Cover letters written by participants who had practiced with AI were as likely to secure a hypothetical job interview as cover letters by participants who had practiced with editor feedback (.51 vs.
+.51, p = .843), and more likely to than those written by participants who had practiced with Google Search (.51 vs.
+.47, p = .024).
+See Figure 6B.
+A possible concern is that reviewing many letters consecutively caused editors to get tired and give worse feedback over time.
+This could suggest that our experiment compared AI against a “strawman” version of human expertise.
+However, analysis of the editors’ performance revealed the opposite pattern: while editors became more efficient, spending less time per letter as the task progressed (b = 0.387, p < .001), the quality of their edits and feedback actually improved (b = 0.013, p < 0.001).
+Thus, the AI tool’s advantage over professional editors cannot be attributed to a decline in editorial performance.
+See Section E2 of Supplementary Information for details.
+Practicing with AI was not more effortful than getting feedback from professional editors and was less effortful than looking for examples online
+During the practice phase, participants with access to AI spent about the same time as those who practiced with editor feedback (difference = 0.43 minutes, d = 0.09, p > .05), but about 1.18 fewer minutes than those who practiced with Google Search (d = 0.32, p < .001).
+Nevertheless, participants across all three groups reported similar levels of subjective effort (ps > .05).
+Likewise, participants who practiced with AI logged about as many keystroles as those who practiced with editor feedback (68% fewer, d = 0.08, p > .05) and 76% fewer keystrokes than those who practiced with Google Search (d = 0.32, p < .001).
+They also typed at a slower rate than participants who practiced with editor feedback (d = 0.41, p < .001) or those who practiced with Google Search (d = 0.34, p < .001).
+No evidence of AI-induced illusion of mastery
+Participants who had practiced with AI did not report learning more than those who had practiced with editor feedback (d = 0.07, p > .05), nor did they consider their writing skills to be better (d = 0.02, p > .05).
+Both groups, however, correctly recognized that they had learned more than those who had practiced with Google Search (ds = .28 and .20, respectively), and evaluated their writing skills accordingly (ds = .11 and .12, respectively).
+Participants who had practiced with AI and those who had practiced with editor feedback requested feedback on their test submissions at similar rates (71% and 70%, respectively).
+Both groups were more likely to request feedback than those who had practiced with Google search (64%, p = .003 and .007, respectively).
+See Section E6 in the Supplementary Information for details.
+Results were not moderated by individual differences
+As in Study 2, the above findings were consistent across all measured individual differences. See Table S22 in the Supplementary Information for details.
+Study 5: Examples as the mechanism for AI learning gains
+To better understand what drives the benefit of AI practice, in Study 5 (N = 2,003), we preregistered a replication and extension in which we replaced the no-practice condition of Study 2 with an example-only condition.
+In this condition, we showed participants an AI-generated writing example that they could not edit.
 To the extent that the benefit of practicing with AI was driven by exposure to examples, the example-only condition should improve performance in the test phase as much as the practice with AI condition.
@@
 As in Study 2, participants given access to the AI writing tool dramatically outperformed participants who did not get access to it, both while using it during the practice phase (d = 1.22, p <.001), and during the no-AI test phase (d = .34, p < .001).
-Their test-phase cover letters were also relatively more likely to secure them hypothetical job interviews (.51 vs.
+Their test phase cover letters were also relatively more likely to secure them hypothetical job interviews (.51 vs.
 .47, p = .008).
-Participants who had merely seen an AI-generated example also improved more in writing skill than those who had practiced without AI (d = .37, p = < .001), and produced letters that were relatively more likely to secure them interviews (.52 vs.
+Participants who had merely seen an AI-generated example also improved more in writing skill than those who had practiced without AI (d = .37, p < .001), and produced letters that were relatively more likely to secure them interviews (.52 vs.
 .47, p = .007).
@@
 .52, p = .561).
-See Figures S6 and 7.
-As in Study 2, gains from practicing with AI and seeing AI examples were comparable across all levels of baseline writing skill.
+See Figures S10 and 7B.
 Seeing an AI example was even less effortful than practicing with AI
 During the practice phase, participants who saw the AI example spent 2.32 fewer minutes than participants practicing with AI (d = 0.99, p < .001) and 2.96 fewer minutes than participants practicing without AI (d = 1.13 p = < .001), and reported expending less effort than those practicing with an AI tool (d = 0.19, p = .003) and those practicing without AI (d = 0.32, p < .001).
-As in Study 2, participants who practiced with AI logged 2.83 times fewer keystrokes than did participants who practiced without AI.
+As in Study 2, participants who practiced with AI logged 74% fewer keystrokes and 68% fewer keystrokes per minute (d = 0.15, p = .007) compared to participants who practiced without AI.
 As expected, participants exposed to the AI example logged 0 keystrokes.
-During the test phase, participants who had seen the AI example worked for an additional 56 seconds more than participants who had practiced without AI (d = 0.22, p < .001).
-They also logged 30% more keystrokes than participants who had practiced with AI (d = 1.86 p <.001) and 60% more than those who practiced without AI (d = 2.39, p < .001).
-Across conditions, all participants self-reported similar levels of subjective effort (ds < .08, ps > .05).
 Seeing an AI example did not create the illusion of mastery
-As in Study 2, despite learning more, participants who had practiced with AI or had merely seen an AI-generated example reported learning similar amounts to those who practiced without AI (ps > .05) and rated their writing skill after practice at comparable levels (ps > .05).
+As in Studies 2 and 4, despite learning more, participants who had practiced with AI or had merely seen an AI-generated example reported learning similar amounts to those who practiced without AI (ps > .05) and rated their writing skill after practice at comparable levels (ps > .05).
 Moreover, all participants requested feedback at similar rates (proportions ranged from 63% to 67%).
-See Section C4 in the Supplementary Information for details.
-The effects of seeing an AI example persist
-When we recontacted a subsample of (n = 800) participants one day later, the majority responded (n = 633, 80%); the attrition rates ranged from 17% to 24% and did not differ by condition ( (2) = 4.56, p = .102).
+See Section F4 in the Supplementary Information for details.
+The benefits of seeing an AI example were just as large a day later
+When we recontacted a subsample of participants (n = 800) one day later, the majority responded (n = 633, 80%); the attrition rates ranged from 17% to 24% and did not differ by condition ( (2) = 4.56, p = .102).
 The effect remained robust after 24 hours.
@@
 Participants who had merely seen an AI example performed as well as those who had practiced with AI (d = .02, p = .830).
-See Figure 6 and Section C5 of Supplementary Information for details.
-As in Study 2, the above findings were not moderated by individual differences. See Table S23 in the Supplementary Information for details.
+As in Study 2, the magnitude of the follow-up effect size was not significantly different from the immediate test phase effect size (condition time interactions were non-significant (ps > .96).
+See Figure 7A and Section F5 of Supplementary Information for details.
+Results were not moderated by individual differences
+As in the previous studies, the above findings were consistent across all measured individual differences. See Table S35 in the Supplementary Information for details.
 Discussion
-Contrary to the expectations of lay forecasters (Study 1), participants who practiced writing cover letters with an AI tool learned more than those who practiced on their own (Studies 2 and 3).
-Specifically, participants who had practiced with AI wrote (without AI assistance) cover letters that were rated higher in writing quality and were more likely to secure a hypothetical interview—both immediately after practice and one day later.
-Learning gains were not the result of greater effort; in fact, participants who’d practiced with AI expended less effort during practice than those who’d practiced alone.
-Instead, these gains can be explained by exposure to a high-quality, just-in-time personalized example: participants who merely viewed an AI-generated example cover letter (without editing it) improved their writing skill as much as those given the option to practice editing the cover letter (Study 3).
-While not obvious to forecasters, teachers (11), and the general population (8), the benefits of seeing AI-generated examples are consistent with prior research on how people learn.
-In addition to the literature on worked examples mentioned earlier, research has shown that humans are especially adept at observing, imitating, and learning from others (28, 29, 30).
+Contrary to lay forecasters’ expectations (Studies 1 and 3), participants who practiced writing cover letters with an AI tool learned more than those who practiced on their own (Studies 2 and 5), more than those who had access to Google cover letter examples, and even more than those who received personalized feedback from professional human editors (Study 4).
+Specifically, participants who had practiced with AI, subsequently wrote cover letters (without AI assistance) that were rated higher in writing quality and were more likely to secure a hypothetical interview—both immediately after practice and one day later.
+Learning gains were not the result of greater effort; in fact, participants who had practiced with AI expended less effort during practice than those who had practiced alone.
+Instead, these gains were at least in part explained by exposure to high-quality, just-in-time personalized examples: participants who merely viewed an AI-generated cover letter (without editing it) improved their writing skill as much as those given the option to practice editing the cover letter (Study 5).
+Study 1 forecasters who believed AI would harm learning cited reduced effort as the mechanism—and prior research supports this concern (28, 29).
+All things equal, the more effort a learner invests in a given practice task, the more they learn.
+So why did practicing with AI improve writing skill more than practicing alone, even while reducing both the quantity and quality of effort?
+Likewise, why did writers who googled how to compose a cover letter work longer and harder than writers who practiced with AI, yet learn less?
+We propose a third mechanism that offsets decrements in human effort: the utility of the practice task itself.
+See Figure 8.
+Perhaps the excellent personalized examples provided by AI are analogous to demonstrations of an Olympic coach.
+The higher the utility of the practice task, the faster the learning.
+While not immediately apparent to forecasters, teachers (11), and the general population (8), the pedagogical benefits of viewing AI-generated examples are consistent with prior research on learning.
+Research has shown that humans are especially adept at observing, imitating, and learning from others (30, 31, 32).
 Our findings also align with the expert performance literature: the most successful learners engage in deliberate practice, which (in addition to concentration, feedback, and repetition) depends upon detailed mental representations of excellent performance (33).
@@
 Three promising directions for future research are worth highlighting.
-First, it remains to be seen whether learning from AI examples occurs in domains other than writing, currently the modal use case for AI tools in the workplace (25) .
-Writing may be particularly suited to learning by example because at a glance, a single AI-generated example visually communicates the elements of effective professional writing (e.g., short sentences and boldface formatting).
-In other domains, merely observing a solution may be less informative.
+First, it is not clear whether the observed benefits of AI examples would be as large in domains other than professional writing.
+At a glance, a single AI-generated example visually communicates the elements of effective professional writing (e.g., short sentences and boldface formatting).
+In other domains, however, merely observing a solution may be less informative.
 For instance, the final answer to a math or computer programming problem does not instantly reveal the procedure that produced it.
-Studies of AI tutors in these domains have, perhaps not surprisingly, found null or negative effects (23, 24).
 Second, additional research is needed to explore moderators of learning from AI tools.
-Certain strategies for interacting with AI may bolster their effectiveness.
+Certain metacognitive strategies may enhance the benefits of interacting with AI.
 For example, experimental research suggests that learners benefit more from AI explanations for math problems if they first try to solve them on their own (21).
-Similarly, correlational evidence that asking AI for explanations as opposed to answers is associated with more learning in mathematics (23) and computer programming (22).
-On the other hand, other factors could minimize learning benefits.
+Similarly, correlational evidence suggests that asking AI for explanations as opposed to answers is associated with more learning in mathematics (23) and computer programming (22).
+Conversely, other factors may minimize learning benefits.
 In our experimental paradigm, participants practiced for as long as they wanted, with the foreknowledge that their skills would subsequently be tested (and rewarded monetarily) without access to AI.
@@
 In real-world settings, there is often time pressure and competing incentives around performance and learning, which we speculate would reduce the learning gains associated with practicing with AI.
-Third, in our experimental paradigm, participants who interacted with the AI tool did so only once.
+More broadly, we believe concerns about AI as a crutch are justified when AI displaces effort without increasing task utility.
+Third, in our experimental paradigm, participants interacted with the AI tool only once.
 It is common, however, to use AI tools repeatedly.
@@
 The introduction of superhuman AI has been associated with an increase in the novelty and quality of decisions made by human Go players over time, with elite players reporting that they have been inspired by AI solutions they’d never seen before (34).
-Future research, ideally in field settings, is needed to establish the long-term benefits and costs of relying on AI tools.
+Future research, ideally in field settings, is needed to establish the long-term benefits and costs of using AI tools.
 Conclusion
 Our findings should temper widespread concern that AI tools invariably boost momentary productivity at the expense of long-term skill development.
-Although it reduced the effort users invested in practicing, the AI writing tool nevertheless accelerated skill development.
-It accomplished this by providing high-quality, just-in-time, personalized examples of excellent writing.
-The underappreciated efficacy of timely and tailored examples has practical implications for the design of AI tools.
-Many AI tools designed to support learning are explicitly programmed not to “give away” answers.
-It may be that in addition to hints, leading questions, and explanations, learners benefit from demonstrations of the principles they are attempting to master.
-Decades before the advent of generative AI, the legendary UCLA baseball coach John Wooden declared that the four laws of learning are explanation, demonstration, imitation, and repetition (33).
+Although the AI writing tool reduced the effort users invested in practicing, it nevertheless accelerated their skill development.
+The underappreciated efficacy of timely and tailored examples carries practical implications: Many AI tools designed to support learning are explicitly programmed not to “give away” answers.
+However, it may be that in addition to hints, leading questions, and explanations, learners benefit from demonstrations of the principles they are attempting to master.
+Decades before the advent of generative AI, the legendary UCLA basketball coach John Wooden declared that the four laws of learning are explanation, demonstration, imitation, and repetition (35).
 Few learners have access to the best human teachers, coaches, and mentors, but generative AI now makes it possible to learn from personalized, just-in-time demonstrations tailored to any domain.
-In doing so, AI has the potential not only to boost productivity but also to democratize opportunities to build human capital at scale.
+In doing so, AI has the potential not only to boost productivity but also to democratize opportunities for building human capital at scale.
 Methods
@@
 Pre-registration
-All our studies were pre-registered. See pre-registrations for Study 1 at https://aspredicted.org/x9mm-7qwp.pdf.
-Study 2 was run twice because of a technical issue that caused imbalanced missing data issue the first time it was run.
-The findings in this sample were consistent with the ones reported here.
-Pre-registration is available at https://aspredicted.org/4sw4-mpny.pdf.
-The version of Study 2 presented above was pre-registered https://aspredicted.org/xyyn-gmzc.pdf.
-Pre-registartion for Study 3 is available at https://aspredicted.org/3mty-fcfy.pdf.
-The one-day follow-up for Study 3 was collected in three batches.
-We pre-registered batch 2 https://aspredicted.org/5jcx-bhg9.pdf, but report pooled results in the main text.
-Details on the separate batches are available in Supplementary Information Section C5.
+All studies were preregistered on https://aspredicted.org. Accordingly, the analyses presented in the main text were also pre-registered, with exceptions noted below.
+Study 1 was preregistered (#191800). The moderation analyses were not pre-registered.
+Study 2 was pre-registered (#201140).
+The pairwise comparisons analysis was pre-registered separately (#205316).
+After preregistration, we considered a keystrokes-per-minute metric of effort intensity rather than duration; thus, this analysis was not preregistered.
+Study 3 was pre-registered (#236487). Analyses on predicted effectiveness of AI vs. human feedback and the mediation analysis were not preregistered.
+Study 4 was pre-registered (#239157). The pairwise comparisons analysis was pre-registered separately (#243115). The keystrokes-per-minute metric of effort was not pre-registered.
+Study 5 was pre-registered (#197704).
+The one-day follow-up was collected in three batches: a pilot, followed by second and third batches.
+We pre-registered batch 2 (#199451) but report pooled results in the main text, with separate batch details available in Supplementary Information Section F5.
+The pairwise comparisons analysis was pre-registered separately (#205315).
+The keystrokes-per-minute metric of effort was not pre-registered.
 Participants
 We sampled participants from Prolific from all our studies. We excluded all Prolific users who participated in one of the earlier studies from participation in subsequent studies.
-Participants in Study 1 (N = 150) were predominantly female (n = 93, 62%), and ranged in age from 21 to 81 (M = 38.4, SD = 12.2).
+Participants in Study 1 were predominantly female (n = 93, 62%), with ages ranging from 21 to 81 (M = 38.4, SD = 12.2).
 They were predominantly white (75%).
-A small proportion were students (13%), and most were employed (62%)
-In Study 2, the sample was more evenly split between men (46%) and women (52%), and ranged in age from 18 to 82 (M = 36.0, SD = 12.5).
-Over half of the sample (58%) was white, with the rest being comprised by Black (33%), latino (6%), and asian (5%).
-77% had college degrees.
+A small proportion were students (13%), and most were employed (62%).
+In Study 2, the sample was more evenly split between men (46%) and women (52%), ranging in age from 18 to 82 (M = 36.0, SD = 12.5).
+Over half of the sample (58%) was White, with the rest comprising Black (33%), Latino (6%), and Asian (5%) participants.
+Most participants (77%) had college degrees.
 Most participants (93%) were at least somewhat motivated to improve their writing, and had varying levels of experience with AI writing assistants (36% had tried them, but hardly ever used them, 47% used them at least a few times per week, and 17% had never used AI assistants before).
-Study 3 had similar prooportions of men (46%) and women (53%), and participants ranged in age from 18 to 95 (M = 37.9, SD = 12.6).
-Over half of the sample (64%) was white, with the rest being comprised of Black (24%), Latino (8%), and Asian (6%) participants.
+Participants in Study 3 were evenly split between men (52%) and women (48%), and their ages ranged from 18 to 75 (M = 42.4, SD = 14.4).
+They were predominantly white (67%).
+A small proportion were students (19%), and most were employed (56%).
+In Study 4, the sample consisted of relatively fewer men (38%) than women (61%), and ranged in age from 19 to 92 (M = 41.5, SD = 13.8).
+Over half of the sample (74%) was White, with the rest comprising Black (16%), Latino (8%), and Asian (7%) participants.
+Most participants (74%) had college degrees.
+Most participants (94%) were at least somewhat motivated to improve their writing, and had varying levels of experience with AI writing assistants: 17% had never used AI assistants, 40% had tried them but hardly ever used them, and 43% used them at least weekly.
+Study 5 had similar proportions of men (46%) and women (53%), and participants ranged in age from 18 to 95 (M = 37.9, SD = 12.6).
+Over half of the sample (64%) was White, with the remainder comprising Black (24%), Latino (8%), and Asian (6%) participants.
 Most participants (74%) had college degrees.
@@
 Procedure
-Participants first saw an introductory screen about what they were about to do.
-They then completed a brief questionnaire where they reported their demographics, their experience with AI writing tools, their motivation to improve their writing, and their perceived writing skill.
-They then completed a 2-minute pre-test in which they saw a poorly worded cover letter and were asked to improve it.
-After this, all participants completed a lesson about the five principles of effective writing.
-They were then randomized to the practice condition, or skipped ahead if assigned to the no-practice control.
-During the practice phase, participants rewrote a different cover letter, or (in the see example condition) generated an AI rewritten version of that letter.
+All studies followed the same general structure with study-specific variations.
+Participants first saw an introductory screen and completed a brief questionnaire reporting demographics, prior experience with AI writing tools, motivation to improve their writing, and perceived writing skill.
+They then completed a short pre-test (2 minutes) in which they edited a poorly written cover letter, followed by a brief lesson on the five principles of effective writing.
+In Studies 2 and 5, participants were then randomized to a practice condition (or skipped practice in the no-practice control).
+During practice, participants rewrote a new cover letter or, in the example-only condition, viewed an AI-rewritten version of the letter.
 This example was not explicitly labeled as AI-generated.
-Then, participants saw the text they submitted (or the AI-generated example), and below it, AI-generated feedback highlighting one way in which this letter could be improved.
-See more information about the feedback procedure in Supplementary Information section A3.
-Immediately after seeignt this feedback screen, all participants then completed two questions, reporting how much they had learned and how hard they had worked on the task so far.
-They then edited a new cover letter in an incentivized, 7-minute test without the help of any AI tools.
+After practice, participants saw either their own rewritten letter or the AI-generated example, along with AI-generated feedback highlighting one suggested improvement (see Supplementary Information, Section A3).
+Immediately after this feedback, participants rated how much they had learned and how hard they had worked so far.
+They then completed a 7-minute incentivized test of writing skill by editing a new cover letter without access to AI tools.
 To minimize the possibility of cheating, we used custom JavaScript to restrict copy-pasting functionality.
-Finally, participants were invited to see optional feedback and were asked if they had used any outside resources during the test.
+Finally, participants could request optional feedback and were asked whether they had used outside resources during the test.
+A small percentage of participants (2.9%) admitted to doing so; as per our preregistration, these participants are included in analyses, though results excluding them are consistent with our main interpretation (see Tables S6, S18, and S27).
+In Study 4, we adapted the procedure to compare AI practice with two ecologically valid human alternatives: feedback from professional editors and guidance from Google Search.
+Participants completed the same baseline questionnaire, pretest, and writing lesson.
+They then edited a new cover letter without any assistance and were randomly assigned to one of three conditions: (1) practice with AI, (2) practice with feedback from professional editors, (3) practice with a Google Search for cover letter examples and tips.
+In the second condition condition, participants’ cover letters were forwarded to editors recruited in partnership with The Journalists’ Resource at Harvard Kennedy School, who provided both written feedback and a rewritten version of the letter.
+The following day, participants were asked to revise their original cover letter using the resources from their assigned condition.
+All participants then completed a final writing test by rewriting a new cover letter without assistance, followed by an exit questionnaire.
 Measurement
@@
 To do this, we independently rated each cover letter and each writing principle.
-Research has demonstrated that large language models can provide ratings of writing quality that align closely with human judgments, offering reliability and consistency across various evaluation contexts (34, 35).
+Research has demonstrated that large language models can provide ratings of writing quality that align closely with human judgments, offering reliability and consistency across various evaluation contexts (36, 37).
 See our prompts in Table S1.
 We then took the unweighted average of these 5 scores as our main dependent variable.
-See disaggregated analyses by each writing principle and additional outcomes on Table S6 and S16.
-To validate these ratings, the first author and a trained research assistant took a sample of n = 30 cover letters from Study 2, and rated them on the 5 principles.
-The average of these two ratings correlated more highly with the computer ratings (r = .83, p < .001), than the average interrater reliability (r = .69, p < .001).
+See disaggregated analyses by each writing principle and additional outcomes on Tables S6, S18, and S27.
+To validate these ratings, the first author and a trained research assistant took a sample of n = 100 cover letters from Study 2, and rated them on the 5 principles.
+The average of these two ratings correlated about as highly with the computer ratings (r = .70, p < .001) as they did with the average interrater reliability (r = .74, p < .001).
+These ratings had high internal consistency (Cronbach’s s = 0.81, 0.78, and 0.83 for studies 2, 4, 5, respectively.)
 To address concerns that particular LLMs might be biased in favor of their own output, we also used Claude to rate the cover letters.
-We find that GPT ratings correlate highly with Claude ratings (r = .71, p < .001), and that the effects are not attenuated by using different models (See Tables S6 and S16), suggesting that our effects are not explainable by same-model bias.
+We find that GPT ratings correlate strongly with Claude ratings (r = .71, p < .001), and that the effects are not attenuated by using different models (See Tables S6, S18, and S27), suggesting that our effects are not explainable by same-model bias.
+Additionally, we asked a separate sample of participants to read pairs of test-phase cover letters randomly selected from different conditions, and to indicate which letter would be more likely to secure a job interview.
+Cover letters more likely to secure an interview obtained higher AI ratings of writing skill (rs = .29, .35, and .28, ps < .001, for Study 2, 4 and 5, respectively).
 Statistical analysis
 As per our pre-registrations, we fit ANCOVA models predicting outcomes from condition indicators, controlling for pretest score and baseline characteristics (age, gender, race/ethnicity, primary language, education level, motivation to improve writing skills, self-rated writing skill, experience with AI writing assistants, and baseline writing effectiveness).
-We used logistic regression to predict whether participants chose to see optional feedback for their test from condition, controlling for pretest score and baseline characteristics.
+We used logistic regression to predict whether participants chose to see optional test feedback from condition, controlling for pretest score and baseline characteristics.
 Our analyses of the hypothetical hiring situation use beta regression, because the relative likelihood of a cover letter being preferred is bounded between 0 and 1.
-When correcting for multiple comparisons in exploratory moderation analyses, we use the Benjamini-Hochberg correction (36).
-C Results Study 3
+When correcting for multiple comparisons in exploratory moderation analyses, we used the Benjamini-Hochberg correction (38).
+C Results Study 2
 - C1 Randomization, Balance, and Missingness
-- C2 AI examples improve writing skill
-- C3 Seeing AI examples was less effortful
-- C4 Seeing an AI example did not discourage motivation for future learning
-- C5 The effects of seeing an AI example persist
-- C6 Seeing AI examples was equally effective across subgroups
+- C2 AI practice improved writing skill
+- C3 AI practice was less effortful
+- C4 AI practice did not discourage motivation for future learning
+- C5 The benefits of practicing with AI were just as large a day later
+- C6 AI practice was equally effective across subgroups
+- C7 Pairwise comparisons
+- D Results Study 3
+E Results Study 4
+- E1 Randomization, Balance, and Missingness
+- E2 Details on professional editors
+- E3 Details on participants’ search behavior
+- E4 AI practice improved writing skill more than getting feedback from professional editors and looking for examples online
+- E5 AI practice wasn’t any more effortful than getting feedback from professional editors and was less effortful than looking for examples online
+- E6 AI practice did not create more of an illusion of mastery than getting feedback from professional editors or looking for examples online
+- E7 Results were not moderated by individual differences
+- E8 Pairwise comparisons
+F Results Study 5
+- F1 Randomization, Balance, and Missingness
+- F2 AI examples improve writing skill
+- F3 Seeing AI examples was less effortful
+- F4 Seeing an AI example did not discourage motivation for future learning
+- F5 The benefits of seeing an AI example were just as large a day later
+- F6 Seeing AI examples was equally effective across subgroups
+- F7 Pairwise Comparisons
 Appendix A Additional methods
 A1 AI Ratings
-After participants completed the procedure outlined in Figure 2, we had three writing samples for participants who practiced with or without AI (one for each phase of the experiment: pretest, practice, test), and two writing samples for participants who did not practice, or simply saw an AI generated example.
+After participants completed the procedure outlined in Figure 2, we had three writing samples for participants who practiced with or without AI (one for each phase of the experiment: pretest, practice, test), and two writing samples for participants who did not practice, or simply saw an AI-generated example.
 We used GPT-4o to rate these texts for five writing principles.
@@
 Claude and GPT-4o ratings were positively correlated, with pretest ratings being less so.
-In Study 2, the correlations were .38, .78, .65, and .67 for pretest, practice, test, and followup, respectively.
+In Study 2, the correlations were .38, .78, .65, and .67 for pretest, practice, test, and follow-up, respectively.
 All p-values were below .001.
@@
 They were asked to “Imagine you’re hiring a social media manager for your company; which cover letter would make you more likely to offer an interview to the candidate?
-Choose one.”.
-Each cover letter was compared to at least three other letters, sampled uniformly at random from the other two conditions.
-Most letters were compared against 3 or 4 other letters, min, max, med (plot?).
-For each cover letter we calculated the relative likelihood of it securing a hypothetical interview, defined as the total number of times that letter was preferred, divided by the total number of contests for that cover letter.
-As shown in Figure S2, the relative likelihood of an interview correlated positively with the AI-generated measure of writing skill.
-Table S2 shows the beta regressions for the relative likelihood by condition.
-| (1) | (2) | (3) | (4) | |
-| (Intercept) | *** | ** | ||
-| No practice | *** | *** | ||
-| Practice wo AI | ** | |||
-| Practice w AI | ** | |||
-| See AI example | *** | |||
-| Precision () | *** | *** | ||
-| Symmertry (Log()) | *** | |||
-| Demographic and baseline performance covariates | No | Yes | No | Yes |
-| Num.Obs. | ||||
-| AIC | ||||
-| BIC | ||||
-| Log.Lik. | ||||
-| RMSE |
+Choose one.” Each cover letter was compared to at least three other letters, sampled uniformly at random from the other two conditions.
+Most letters were compared against 3 or 4 other letters.
+For each cover letter, we calculated the relative likelihood of it securing a hypothetical interview, defined as the total number of times that letter was preferred, divided by the total number of contests for that cover letter.
 A3 Feedback
@@
 The feedback was displayed immediately after the practice cover letter submission.
-The feedback page read: “Here is the email”, then reproduced the participants submission verbatim, then read “Here is one way in which it could be made better.” The feedback was personalized and created by GPT-4o model.
-The feedack prompt is shown in Figure S3.
-Appendix B Results Study 2
-B1 Randomization, Balance, and Missingness
+The feedback page read: “Here is the email, then reproduced the participant’s submission verbatim, then read “Here is one way in which it could be made better.” The feedback was personalized and created by the GPT-4o model.
+The feedback prompt is shown in Figure S2.
+Appendix B Results Study 1
+Table S2 reports results from a logistic regression predicting the probability of answering correctly as a function of demographic and background characteristics.
+Odds ratios (ORs) are displayed along with their standard errors, Wald z statistics, and p-values.
+The model includes age, sex, ethnicity, student status, employment, and prior AI experience as predictors.
+| Predictor | OR | SE | z | p-value |
+| Prior AI experience | 0.718 | 0.199 | -1.664 | 0.096 |
+| Age (years) | 1.014 | 0.021 | 0.675 | 0.500 |
+| Male (vs. Female) | 1.035 | 0.467 | 0.073 | 0.942 |
+| White (vs. Other) | 0.966 | 0.509 | -0.068 | 0.946 |
+| Student (vs. Not) | 1.120 | 0.618 | 0.183 | 0.855 |
+| Employment status | ||||
+| Other | 0.770 | 0.687 | -0.382 | 0.703 |
+| Part-time | 0.663 | 0.601 | -0.684 | 0.494 |
+| Unemployed (seeking) | 0.645 | 0.593 | -0.738 | 0.461 |
+Appendix C Results Study 2
+C1 Randomization, Balance, and Missingness
 To allow users to format their responses flexibly, we used the TinyMCE rich text editor, which is interfaced with Qualtrics.
@@
 This attrition was not selective by condition.
-As shown in Table S3, missingness and attrition rates were low for the main and followup samples, and did not differ by condition.
-| Condition | Main Sample | Followup Sample |
+As shown in Table S3, missingness and attrition rates were low for the main and follow-up samples, and did not differ by condition.
+| Condition | Main Sample | Follow-up Sample |
 | No practice | ||
-| Practice wo AI | ||
-| Practice w AI | ||
+| Practice w/o AI | ||
+| Practice w/ AI | ||
 | Overall | ||
@@
 See Table S4.
-| Overall | No practice | Practice wo AI | Practice w AI | p | SMD | |
+| Overall | No practice | Practice w/o AI | Practice w/ AI | p | SMD | |
 | n | 2238 | 755 | 752 | 731 | ||
@@
 | Pretest Writing Skill (mean (SD)) | 3.32 (0.78) | 3.31 (0.73) | 3.32 (0.78) | 3.32 (0.83) | .923 | 0.013 |
-B2 AI practice improved writing skill
+C2 AI practice improved writing skill
 The AI tool improved performance while participants used it.
@@
 The robustness checks included after the main specification, show that results are similar when using a different language model (Column 2), when not including control variables (Column 3), when excluding participants who admitted to cheating in the test phase (Column 4), for the subset of non-attriting participants to the follow-up phase (Column 5), and for each of the 5 principles separately (Columns 6 - 10).
-| GPT-4o | Claude | Ex. Controls | Ex. Cheaters | Followup | LM | ER | EN | F | ER | |
+| GPT-4o | Claude | Ex. Controls | Ex. Cheaters | Follow-up | LM | ER | EN | F | ER | |
 | Means — (SE) | ||||||||||
-| Practice wo AI | 4.58 | 6.52 | 4.41 | 4.42 | 4.27 | 4.29 | 6.31 | 5.65 | 3.38 | 4.26 |
+| Practice w/o AI | 4.58 | 6.52 | 4.41 | 4.42 | 4.27 | 4.29 | 6.31 | 5.65 | 3.38 | 4.26 |
 | (.222) | (.091) | (.054) | (.055) | (.281) | (.230) | (.150) | (.258) | (.480) | (.268) | |
-| Practice w AI | 6.05 | 7.01 | 5.86 | 5.89 | 5.78 | 5.56 | 7.11 | 6.75 | 6.30 | 5.54 |
+| Practice w/ AI | 6.05 | 7.01 | 5.86 | 5.89 | 5.78 | 5.56 | 7.11 | 6.75 | 6.30 | 5.54 |
 | (.224) | (.092) | (.055) | (.055) | (.286) | (.232) | (.151) | (.260) | (.484) | (.271) | |
 | Effect Sizes (d) — (SE) | ||||||||||
-| Practice wo AI vs. Practice w AI | 1.01*** | .81*** | .98*** | 1.00*** | 1.05*** | .84*** | .82*** | .65*** | .93*** | .73*** |
+| Practice w/o AI vs. Practice w/ AI | 1.01*** | .81*** | .98*** | 1.00*** | 1.05*** | .84*** | .82*** | .65*** | .93*** | .73*** |
 | (.056) | (.055) | (.055) | (.056) | (.061) | (.055) | (.055) | (.054) | (.056) | (.055) | |
@@
 Cheaters excludes the 3% of participants who admitted to cheating on the test phase.
-Followup is the subsample of non-attriting participants who returned to the one-day followup.
-LM to ER are disagregated scores for each of the five principles.
+Follow-up is the subsample of non-attriting participants who returned to the one-day follow-up.
+LM to ER are disaggregated scores for each of the five principles.
 LM = Less is More, ER = Easy Reading, EN = Easy Navigation, F = Formatting, ER = Easy Responding.
@@
 For participants assigned to practice with the AI tool, the quality of AI rewrites did not correlate with participants’ final submissions, r = .06, p = .25.
-| GPT-4o | Claude | Ex. Controls | Ex. Cheaters | Followup | LM | ER | EN | F | ER | |
+| GPT-4o | Claude | Ex. Controls | Ex. Cheaters | Follow-up | LM | ER | EN | F | ER | |
 | Means — (SE) | ||||||||||
@@
 | (.161) | (.072) | (.047) | (.048) | (.200) | (.155) | (.143) | (.192) | (.394) | (.202) | |
-| Practice wo AI | 4.53 | 6.74 | 4.51 | 4.52 | 4.63 | 3.84 | 6.03 | 5.56 | 2.76 | 4.45 |
+| Practice w/o AI | 4.53 | 6.74 | 4.51 | 4.52 | 4.63 | 3.84 | 6.03 | 5.56 | 2.76 | 4.45 |
 | (.160) | (.071) | (.048) | (.048) | (.199) | (.154) | (.142) | (.190) | (.392) | (.200) | |
-| Practice w AI | 5.01 | 6.90 | 4.98 | 4.99 | 5.11 | 4.12 | 6.21 | 6.17 | 3.86 | 4.66 |
+| Practice w/ AI | 5.01 | 6.90 | 4.98 | 4.99 | 5.11 | 4.12 | 6.21 | 6.17 | 3.86 | 4.66 |
 | (.161) | (.072) | (.049) | (.049) | (.202) | (.155) | (.143) | (.192) | (.394) | (.202) | |
 | Effect Sizes (d) — (SE) | ||||||||||
-| No practice vs. Practice wo AI | .09 | .09 | .09 | .10 | .09 | .10* | .11* | .01 | .09 | .01 |
+| No practice vs. Practice w/o AI | .09 | .09 | .09 | .10 | .09 | .10* | .11* | .01 | .09 | .01 |
 | (.053) | (.053) | (.052) | (.053) | (.056) | (.053) | (.053) | (.053) | (.053) | (.053) | |
-| No practice vs. Practice w AI | .47*** | .36*** | .46*** | .46*** | .48*** | .34*** | .28*** | .42*** | .46*** | .14** |
+| No practice vs. Practice w/ AI | .47*** | .36*** | .46*** | .46*** | .48*** | .34*** | .28*** | .42*** | .46*** | .14** |
 | (.054) | (.053) | (.053) | (.054) | (.057) | (.053) | (.053) | (.053) | (.054) | (.053) | |
-| Practice wo AI vs. Practice w AI | .38*** | .28*** | .36*** | .36*** | .39*** | .23*** | .17** | .41*** | .36*** | .13* |
+| Practice w/o AI vs. Practice w/ AI | .38*** | .28*** | .36*** | .36*** | .39*** | .23*** | .17** | .41*** | .36*** | .13* |
 | (.054) | (.053) | (.053) | (.054) | (.057) | (.054) | (.053) | (.054) | (.054) | (.053) | |
@@
 Cheaters excludes the 3% of participants who admitted to cheating on the test phase.
-Followup is the subsample of non-attriting participants who returned to the one-day followup.
-LM to ER are disagregated scores for each of the five principles.
+Follow-up is the subsample of non-attriting participants who returned to the one-day follow-up.
+LM to ER are disaggregated scores for each of the five principles.
 LM = Less is More, ER = Easy Reading, EN = Easy Navigation, F = Formatting, ER = Easy Responding.
@@
 | ||||||||||
-B3 AI practice was less effortful
+C3 AI practice was less effortful
 Table S7 shows OLS models predicting practice effort metrics from practice condition.
@@
 | Means — (SE) | |||||
-| Practice wo AI | 2.37 | 4.31 | 6.52 | 6.76 | 430.95 |
+| Practice w/o AI | 2.37 | 4.31 | 6.52 | 6.76 | 430.95 |
 | (.152) | (.322) | (.291) | (.913) | (57.349) | |
-| Practice w AI | 2.30 | 3.36 | 5.93 | 6.62 | 383.38 |
+| Practice w/ AI | 2.30 | 3.36 | 5.93 | 6.62 | 383.38 |
 | (.153) | (.325) | (.293) | (.919) | (57.887) | |
 | Effect Sizes (d) — (SE) | |||||
-| Practice wo AI vs. Practice w AI | -.07 | -.45*** | -.31*** | -.02 | -.13* |
+| Practice w/o AI vs. Practice w/ AI | -.07 | -.45*** | -.31*** | -.02 | -.13* |
 | (.053) | (.054) | (.053) | (.053) | (.053) | |
@@
 | (.069) | (.213) | (.266) | (.257) | (41.391) | |
-| Practice wo AI | 2.15 | 4.87 | 6.91 | 4.98 | 409.09 |
+| Practice w/o AI | 2.15 | 4.87 | 6.91 | 4.98 | 409.09 |
 | (.068) | (.211) | (.265) | (.255) | (41.143) | |
-| Practice w AI | 2.19 | 5.05 | 6.69 | 5.14 | 446.59 |
+| Practice w/ AI | 2.19 | 5.05 | 6.69 | 5.14 | 446.59 |
 | (.069) | (.213) | (.267) | (.257) | (41.408) | |
 | Effect Sizes (d) — (SE) | |||||
-| No practice vs. Practice wo AI | -.31*** | -.09 | .18*** | -.32*** | .03 |
+| No practice vs. Practice w/o AI | -.31*** | -.09 | .18*** | -.32*** | .03 |
 | (.053) | (.053) | (.053) | (.053) | (.053) | |
-| No practice vs. Practice w AI | -.24*** | .02 | .07 | -.24*** | .15** |
+| No practice vs. Practice w/ AI | -.24*** | .02 | .07 | -.24*** | .15** |
 | (.053) | (.053) | (.053) | (.053) | (.053) | |
-| Practice wo AI vs. Practice w AI | .07 | .11* | -.11* | .08 | .12* |
+| Practice w/o AI vs. Practice w/ AI | .07 | .11* | -.11* | .08 | .12* |
 | (.053) | (.053) | (.053) | (.053) | (.053) | |
@@
 Learning rate is defined as the difference between test and pretest, divided by the effort metric.
-It shows how many points (10 point scale) the participant improved per unit effort (e.g., per minute spent practicing).
+It shows how many points (10-point scale) the participant improved per unit effort (e.g., per minute spent practicing).
 Participants who practiced with AI improved their skill more efficiently.
@@
 | Means — (SE) | |||||
-| Practice wo AI | .28 | .12 | .27 | .32 | .43 |
+| Practice w/o AI | .28 | .12 | .27 | .32 | .43 |
 | (.062) | (.094) | (.038) | (.096) | (.073) | |
-| Practice w AI | .43 | .36 | .37 | .61 | .62 |
+| Practice w/ AI | .43 | .36 | .37 | .61 | .62 |
 | (.062) | (.094) | (.038) | (.097) | (.073) | |
 | Effect Sizes (d) — (SE) | |||||
-| Practice wo AI vs. Practice w AI | .39*** | .40*** | .41*** | .47*** | .40*** |
+| Practice w/o AI vs. Practice w/ AI | .39*** | .40*** | .41*** | .47*** | .40*** |
 | (.054) | (.054) | (.054) | (.054) | (.054) | |
@@
 A smaller proportion of participants even edited the cover letter email before passing it along to the AI tool.
-B4 Seeing an AI example did not discourage motivation for future learning
+C4 AI practice did not discourage motivation for future learning
 Table S10 presents differences in perceived learning, perceived writing skill, and the likelihood of asking for feedback across conditions, with effect sizes and means reported for each comparison.
@@
 | (.209) | (.208) | (.064) | |
-| Practice wo AI | 5.90 | 6.61 | .62 |
+| Practice w/o AI | 5.90 | 6.61 | .62 |
 | (.207) | (.206) | (.066) | |
-| Practice w AI | 6.03 | 6.56 | .58 |
+| Practice w/ AI | 6.03 | 6.56 | .58 |
 | (.209) | (.208) | (.068 ) | |
 | Effect Sizes (ds/odds ratios) | |||
-| No practice vs. Practice wo AI | -.00 | .13* | 1.10 |
+| No practice vs. Practice w/o AI | -.00 | .13* | 1.10 |
 | (.053) | (.053) | (.128) | |
-| No practice vs. Practice w AI | .08 | .10 | 1.30* |
+| No practice vs. Practice w/ AI | .08 | .10 | 1.30* |
 | (.053) | (.053) | (.149) | |
-| Practice wo AI vs. Practice w AI | .08 | -.03 | 1.17 |
+| Practice w/o AI vs. Practice w/ AI | .08 | -.03 | 1.17 |
 | (.053) | (.053) | (.134) | |
 | Note. *** p < .001, ** p < .01, * p < .05. | |||
-B5 The effects of practicing with AI persist
+C5 The benefits of practicing with AI were just as large a day later
 Table S11 shows means and standardized differences for measures of writing skill and related outcomes during the follow-up phase.
@@
 | (.212) | (.094) | (.054) | (.054) | (.201) | (.206) | (.235) | (.510) | (.267) | |
-| Practice wo AI | 4.79 | 6.78 | 4.84 | 4.86 | 4.44 | 6.45 | 5.52 | 2.59 | 4.96 |
+| Practice w/o AI | 4.79 | 6.78 | 4.84 | 4.86 | 4.44 | 6.45 | 5.52 | 2.59 | 4.96 |
 | (.211) | (.093) | (.053) | (.054) | (.200) | (.205) | (.234) | (.507) | (.266) | |
-| Practice w AI | 5.34 | 6.95 | 5.35 | 5.37 | 4.72 | 6.67 | 6.14 | 3.85 | 5.30 |
+| Practice w/ AI | 5.34 | 6.95 | 5.35 | 5.37 | 4.72 | 6.67 | 6.14 | 3.85 | 5.30 |
 | (.214) | (.094) | (.055) | (.055) | (.203) | (.208) | (.237) | (.515) | (.270) | |
 | Effect Sizes (d) — (SE) | |||||||||
-| No practice vs. Practice wo AI | .05 | .06 | .07 | .08 | .11 | .08 | -.02 | .05 | -.03 |
+| No practice vs. Practice w/o AI | .05 | .06 | .07 | .08 | .11 | .08 | -.02 | .05 | -.03 |
 | (.056) | (.054) | (.055) | (.056) | (.056) | (.056) | (.056) | (.056) | (.056) | |
-| No practice vs. Practice w AI | .46*** | .33*** | .44*** | .45*** | .32*** | .25*** | .40*** | .45*** | .17** |
+| No practice vs. Practice w/ AI | .46*** | .33*** | .44*** | .45*** | .32*** | .25*** | .40*** | .45*** | .17** |
 | (.057) | (.055) | (.056) | (.057) | (.056) | (.056) | (.057) | (.057) | (.056) | |
-| Practice wo AI vs. Practice w AI | .41*** | .28*** | .37*** | .37*** | .22*** | .17** | .42*** | .40*** | .20*** |
+| Practice w/o AI vs. Practice w/ AI | .41*** | .28*** | .37*** | .37*** | .22*** | .17** | .42*** | .40*** | .20*** |
 | (.057) | (.055) | (.056) | (.056) | (.056) | (.056) | (.057) | (.057) | (.056) | |
@@
 Cheaters excludes the 3% of participants who admitted to cheating on the test phase.
-LM to ER are disagregated scores for each of the five principles.
+LM to ER are disaggregated scores for each of the five principles.
 LM = Less is More, ER = Easy Reading, EN = Easy Navigation, F = Formatting, ER = Easy Responding.
@@
 | |||||||||
-B6 AI practice was equally effective across subgroups
+As mentioned in the main text the effect of practicing with AI and seeing an AI example did not become attenuated one day later. See Table S34.
+| Term | Estimate | SE | t | p-value |
+| Intercept | -11.768 | 3.394 | -3.467 | 0.001 |
+| Condition: No practice | -0.104 | 0.067 | -1.541 | 0.123 |
+| Condition: Practice w/ AI | 0.491 | 0.068 | 7.211 | 0.000 |
+| Condition: No practice Follow-up | 0.035 | 0.098 | 0.359 | 0.720 |
+| Condition: Practice w/ AI Follow-up | 0.039 | 0.099 | 0.399 | 0.690 |
+| Phase: Follow-up | 0.330 | 0.069 | 4.787 | 0.000 |
+| Pretest score | 0.394 | 0.026 | 15.220 | 0.000 |
+| Year of birth | 0.007 | 0.002 | 4.284 | 0.000 |
+| Gender: Male | 0.000 | 0.041 | -0.012 | 0.990 |
+| Gender: Other | -0.272 | 0.210 | -1.292 | 0.196 |
+| Race/ethnicity: White | 0.108 | 0.092 | 1.167 | 0.243 |
+| Race/ethnicity: Black | -0.198 | 0.097 | -2.048 | 0.041 |
+| Race/ethnicity: Asian | 0.239 | 0.109 | 2.193 | 0.028 |
+| Race/ethnicity: Latino | 0.070 | 0.095 | 0.736 | 0.462 |
+| Race/ethnicity: Other | 0.093 | 0.128 | 0.724 | 0.469 |
+| Education: High school graduate (high school diploma or equivalent including GED) | 0.775 | 0.274 | 2.824 | 0.005 |
+| Education: Some college but no degree | 0.699 | 0.271 | 2.579 | 0.010 |
+| Education: Associate degree in college (2-year) | 0.672 | 0.276 | 2.431 | 0.015 |
+| Education: Bachelor’s degree in college (4-year) | 0.682 | 0.269 | 2.536 | 0.011 |
+| Education: Master’s degree | 0.568 | 0.272 | 2.090 | 0.037 |
+| Education: Doctoral degree (PhD) | 1.122 | 0.308 | 3.643 | 0.000 |
+| Education: Non-PhD Professional degree (JD, MD) | 0.474 | 0.332 | 1.429 | 0.153 |
+| Writing skill | 0.014 | 0.013 | 1.045 | 0.296 |
+| Motivation: Hardly motivated | -0.362 | 0.194 | -1.865 | 0.062 |
+| Motivation: Somewhat motivated | -0.237 | 0.174 | -1.360 | 0.174 |
+| Motivation: Very motivated | -0.338 | 0.174 | -1.937 | 0.053 |
+| Motivation: Extremely motivated | -0.429 | 0.178 | -2.410 | 0.016 |
+| Experience: I have tried AI writing assistant(s) but hardly ever use them | 0.145 | 0.061 | 2.354 | 0.019 |
+| Experience: I use AI writing assistant(s) a few times per week | 0.029 | 0.071 | 0.410 | 0.682 |
+| Experience: I use AI writing assistant(s) about once a week | 0.040 | 0.072 | 0.549 | 0.583 |
+| Experience: I use AI writing assistant(s) every day | 0.099 | 0.095 | 1.045 | 0.296 |
+C6 AI practice was equally effective across subgroups
 To test for moderation effects of pre-treatment demographic variables, we ran separate linear in which writing skill during the test phase was regressed on condition, the pre-treatment moderator of interest, writing skill at baseline, and an interaction term between the moderator condition.
@@
 Asking to see feedback was a binary Yes/No variable, and was modelled with logistic regression.
-Models match the pre-registered main specification, and thus control for all other pre-treatment variables.
+Models match the pre-registered main specification, thereby controlling for all other pre-treatment variables.
 Per.
 = Perceived | |||||||||||||
-Appendix C Results Study 3
-C1 Randomization, Balance, and Missingness
+C7 Pairwise comparisons
+The relative likelihood of a cover letter receiving an invitation to an interview was correlated with the GPT-rated writing quality. See Figure S5.
+As shown in Table S14, participants who had practiced writing cover letters with AI were more likely to be invited to a hypothetical job interview.
+| Without controls | With controls | |
+| Means (SE) | ||
+| Practice w AI | .50 | .53 |
+| (.01) | (.01) | |
+| Practice wo AI | .48 | .48 |
+| (.01) | (.01) | |
+| See AI example | .51 | .49 |
+| (.01) | (.01) | |
+| Model coefficients (SE) | ||
+| Precision () | 16.17*** | 5.36*** |
+| (1.47) | (.30) | |
+| Symmetry (Log()) | .35*** | -.91*** |
+| (.08) | (.07) | |
+| Pairwise comparison (SE) | ||
+| Practice w AI - Practice wo AI | .08** | .21*** |
+| (.03) | (.04) | |
+| Practice w AI - See AI example | -.02 | .14** |
+| (.03) | (.04) | |
+| Practice wo AI - See AI example | -.10*** | -.06 |
+| (.03) | (.04) | |
+| Statistics | ||
+| N | 1934.00 | 2153.00 |
+| AIC | 2447.07 | 2329.29 |
+| BIC | 2474.91 | 2499.52 |
+| log(Likelihood) | -1218.54 | -1134.64 |
+Appendix D Results Study 3
+D1 Mediation
+Participants predicted they would learn more from the feedback of human editors than from AI (a path = 1.05, SE = 0.22, p <.001).
+In turn, higher predicted effectiveness predicted greater willingness to pay (b path = 0.06, SE = 0.01, p <.001).
+Participants’ greater willingness to pay for human feedback was largely explained by their prediction that they would learn more from experienced human editors than from AI (indirect effect = 0.06, SE = 0.01, p < .001), accounting for most of the overall tendency to pay more for human editors (total effect = 0.10, SE = 0.02, p < .001).
+Once these predictions were taken into account, the remaining preference for human feedback was small and only marginally significant (c’ path = 0.04, SE = 0.02, p = .054).
+See Figure S6 and Section D1 of Supplementary Information for details.
+Appendix E Results Study 4
+E1 Randomization, Balance, and Missingness
+Participants in Study 4 were recontacted to return two days after the initial intake section.
+While most people responded, 14% of recontacted participants did not respond.
+As shown in Table S15, missingness and attrition rates were low and did not differ by condition.
+| Condition | Retention Rate |
+| Practice w/ AI | 85.11% |
+| Google examples | 85.70% |
+| Editor feedback | 86.95% |
+| Overall | 85.92% |
+| 1.446 | |
+| p-value | .485 |
+Pre-treatment variables were balanced across experimental conditions, ensuring that random assignment was successful.
+To assess balance, we conducted a series of one-way ANOVAs for continuous variables and chi-square tests for categorical variables.
+Given the multiple comparisons, we applied the Benjamini-Hochberg (BH) procedure to control the false discovery rate.
+All statistical tests confirmed that none of the pre-treatment variables differed significantly across conditions.
+See Table S16.
+| Overall | Practice w/ AI | Google examples | Editor feedback | p | SMD | ||
+| n | 2997 | 1001 | 1000 | 996 | |||
+| Age (mean (SD)) | 41.47 (13.83) | 41.38 (13.78) | 41.11 (13.76) | 41.92 (13.95) | .666 | 0.039 | |
+| Gender (%) | .666 | 0.059 | |||||
+| Female | 1824 (60.9) | 596 (59.5) | 628 (62.8) | 600 (60.2) | |||
+| Male | 1124 (37.5) | 390 (39.0) | 352 (35.2) | 382 (38.4) | |||
+| Other | 49 ( 1.6) | 15 ( 1.5) | 20 ( 2.0) | 14 ( 1.4) | |||
+| Race/Ethnicity | |||||||
+| White (%) | 2209 (73.7) | 745 (74.4) | 736 (73.6) | 728 (73.1) | .950 | 0.020 | |
+| Black (%) | 473 (15.8) | 164 (16.4) | 145 (14.5) | 164 (16.5) | .666 | 0.036 | |
+| Asian (%) | 204 ( 6.8) | 61 ( 6.1) | 68 ( 6.8) | 75 ( 7.5) | .666 | 0.038 | |
+| Latino (%) | 253 ( 8.4) | 79 ( 7.9) | 95 ( 9.5) | 79 ( 7.9) | .666 | 0.038 | |
+| Other (%) | 102 ( 3.4) | 35 ( 3.5) | 40 ( 4.0) | 27 ( 2.7) | .666 | 0.048 | |
+| Education Level (%) | .964 | 0.080 | |||||
+| Less than high school degree | 10 ( 0.3) | 4 ( 0.4) | 2 ( 0.2) | 4 ( 0.4) | |||
+| High school graduate (high school diploma or equivalent including GED) | 271 ( 9.0) | 80 ( 8.0) | 94 ( 9.4) | 97 ( 9.7) | |||
+| Some college but no degree | 508 (17.0) | 178 (17.8) | 168 (16.8) | 162 (16.3) | |||
+| Associate degree in college (2-year) | 283 ( 9.4) | 91 ( 9.1) | 98 ( 9.8) | 94 ( 9.4) | |||
+| Bachelor’s degree in college (4-year) | 1165 (38.9) | 398 (39.8) | 394 (39.4) | 373 (37.4) | |||
+| Master’s degree | 596 (19.9) | 199 (19.9) | 193 (19.3) | 204 (20.5) | |||
+| Doctoral/Professional degree | 164 ( 5.5) | 51 ( 5.1) | 51 ( 5.1) | 62 ( 6.2) | |||
+| Perceived Writing Skill (mean (SD)) | 6.69 (1.67) | 6.67 (1.68) | 6.63 (1.69) | 6.76 (1.63) | .666 | 0.051 | |
+| Motivation to Improve Writing (%) | .748 | 0.083 | |||||
+| Not at all motivated | 24 ( 0.8) | 9 ( 0.9) | 6 ( 0.6) | 9 ( 0.9) | |||
+| Hardly motivated | 166 ( 5.5) | 59 ( 5.9) | 55 ( 5.5) | 52 ( 5.2) | |||
+| Somewhat motivated | 1036 (34.6) | 354 (35.4) | 330 (33.0) | 352 (35.3) | |||
+| Very motivated | 1135 (37.9) | 387 (38.7) | 390 (39.0) | 358 (35.9) | |||
+| Extremely motivated | 636 (21.2) | 192 (19.2) | 219 (21.9) | 225 (22.6) | |||
+| Experience with AI writing assistants (%) | .666 | 0.090 | |||||
+| I have never tried any AI writing assistant | 516 (17.2) | 188 (18.8) | 173 (17.3) | 155 (15.6) | |||
+| I have tried AI writing assistant(s) but hardly ever use them | 1211 (40.4) | 383 (38.3) | 411 (41.1) | 417 (41.9) | |||
+| I use AI writing assistant(s) a few times per week | 557 (18.6) | 197 (19.7) | 173 (17.3) | 187 (18.8) | |||
+| I use AI writing assistant(s) about once a week | 503 (16.8) | 158 (15.8) | 178 (17.8) | 167 (16.8) | |||
+| I use AI writing assistant(s) every day | 210 ( 7.0) | 75 ( 7.5) | 65 ( 6.5) | 70 ( 7.0) | |||
+| Pretest Writing Skill (mean (SD)) | 4.40 (0.67) | 4.41 (0.67) | 4.40 (0.68) | 4.40 (0.65) | .964 | 0.008 |
+E2 Details on professional editors
+We recruited 44 professional writers and editors through Journalist’s Resource, paying them $100 to edit 20 cover letters each.
+They were asked to spend about 7 minutes per cover letter.
+In total, they edited 1,227 cover letters.
+To determine if editors changed their behavior over time, we fit a series of multilevel models, with cover letter position as a predictor, random intercept for editors, and text quality (as rated by GPT-4o), time spent per cover letter, and number of insertions, deletions, or modifications (i.e., Levenstein distance).
+These analyses include editors who were able to edit at least 20 letters, and focus on the first 20 letters edited.
+The quality of the edited letters improved over time ( = 0.013, SE = 0.005, p = 0.013), even though editors spent less time as the task progressed ( = -0.387, SE = 0.028, p < .001).
+Interestingly, editors also were able to improve the emails using fewer edits as the task progressed ( = -2.572, SE = 0.926, p = 0.00561).
+Despite spending less time per letter, editors produced higher quality edits with fewer but more effective changes, demonstrating clear learning effects in professional editing tasks.
+E3 Details on participants’ search behavior
+We replicated the first five pages of a Google search for “cover letter examples” and made this available to participants through Qualtrics to monitor their search behavior.
+Across 635 participants assigned to use Google to search for cover letter examples and tips, we found search behavior consistent with prior research on web-browsing behavior (39).
+Most participants (97%) stayed on page 1, while only 3% ventured beyond the first page of results.
+Participants spent an average of 2.8 minutes (median: 1.3 minutes) browsing search results, with an average of 0.9 minutes between clicks.
+On average, participants made 2.9 clicks during their search session.
+E4 AI practice improved writing skill more than getting feedback from professional editors and looking for examples online
+The AI tool improved performance while participants used it.
+Table S17 shows means and standardized differences for different measures of writing skill during the practice phase.
+The robustness checks included after the main specification, show that results are similar when using a different language model (Column 2), when not including control variables (Column 3), when excluding participants who admitted to cheating in the test phase (Column 4), and for each of the 5 principles separately (Columns 5 - 9).
+| GPT-4o | Ex. Controls | Ex. Cheaters | LM | ER | EN | F | ER | |
+| Means | ||||||||
+| AI | 6.24 | 6.22 | 6.28 | 5.61 | 7.26 | 6.88 | 5.41 | 6.04 |
+| (.172) | (.048) | (.173) | (.167) | (.125) | (.190) | (.397) | (.214) | |
+| 4.81 | 4.80 | 4.84 | 4.23 | 6.33 | 5.85 | 2.71 | 4.91 | |
+| (.173) | (.048) | (.174) | (.167) | (.125) | (.191) | (.398) | (.215) | |
+| Editors | 5.18 | 5.17 | 5.21 | 4.55 | 6.56 | 6.24 | 3.44 | 5.09 |
+| (.172) | (.048) | (.173) | (.166) | (.124) | (.190) | (.396) | (.214) | |
+| Effect Sizes (d) | ||||||||
+| AI vs. Google | -1.04*** | -1.01*** | -1.03*** | -1.03*** | -.93*** | -.67*** | -.85*** | -.66*** |
+| (.051) | (.050) | (.052) | (.051) | (.051) | (.050) | (.051) | (.050) | |
+| AI vs. Editors | -.77*** | -.75*** | -.77*** | -.79*** | -.70*** | -.42*** | -.62*** | -.56*** |
+| (.050) | (.049) | (.050) | (.050) | (.050) | (.049) | (.050) | (.050) | |
+| Google vs. Editors | .27*** | .26*** | .27*** | .24*** | .23*** | .25*** | .23*** | .10* |
+| (.049) | (.048) | (.049) | (.049) | (.049) | (.049) | (.049) | (.049) | |
+| Note.
+GPT-4o is the main specification.
+Ex.
+Controls is the main specification, unadjusted for demographic and pretreatment variables, Ex.
+Cheaters excludes the 3% of participants who admitted to cheating on the test phase.
+LM to ER are disaggregated scores for each of the five principles.
+LM = Less is More, ER = Easy Reading, EN = Easy Navigation, F = Formatting, ER = Easy Responding.
+*** p < .001, ** p < .01, * p < .05.
+| ||||||||
+During the test phase, when participants had to rewrite a cover letter without the help of the AI tool, participants who had practiced with AI outperformed participants who searched for cover letter examples and tips and participants who received personalized feedback from professional editors.
+Again, the learning gains are robust to different specifications, subsamples, and measures or writing quality.
+See Table S18.
+| GPT-4o | Claude | Ex. Controls | Ex. Cheaters | LM | ER | EN | F | ER | |
+| Means | |||||||||
+| AI | 4.98 | 5.14 | 4.77 | 4.99 | 3.98 | 6.34 | 6.00 | 3.63 | 4.93 |
+| (.133) | (.103) | (.041) | (.134) | (.126) | (.133) | (.156) | (.333) | (.169) | |
+| 4.44 | 4.90 | 4.24 | 4.46 | 3.59 | 5.95 | 5.50 | 2.43 | 4.74 | |
+| (.133) | (.103) | (.040) | (.134) | (.126) | (.133) | (.156) | (.333) | (.169) | |
+| Editors | 4.75 | 5.03 | 4.55 | 4.76 | 3.70 | 6.10 | 5.95 | 3.18 | 4.80 |
+| (.133) | (.103) | (.040) | (.134) | (.126) | (.133) | (.155) | (.333) | (.169) | |
+| Effect Sizes (d) | |||||||||
+| AI vs. Google | -.46*** | -.27*** | -.44*** | -.46*** | -.36*** | -.34*** | -.37*** | -.41*** | -.12* |
+| (.050) | (.050) | (.049) | (.050) | (.049) | (.049) | (.049) | (.049) | (.049) | |
+| AI vs. Editors | -.20*** | -.12* | -.18*** | -.20*** | -.25*** | -.21*** | -.04 | -.16** | -.09 |
+| (.049) | (.049) | (.048) | (.049) | (.049) | (.049) | (.049) | (.049) | (.049) | |
+| Google vs. Editors | .26*** | .15** | .26*** | .26*** | .11* | .13** | .34*** | .26*** | .04 |
+| (.049) | (.049) | (.048) | (.049) | (.049) | (.049) | (.049) | (.049) | (.049) | |
+| Note.
+GPT-4o is the main specification.
+Ex.
+Controls is the main specification, unadjusted for demographic and pretreatment variables, Ex.
+Cheaters excludes the 3% of participants who admitted to cheating on the test phase.
+LM to ER are disaggregated scores for each of the five principles.
+LM = Less is More, ER = Easy Reading, EN = Easy Navigation, F = Formatting, ER = Easy Responding.
+*** p < .001, ** p < .01, * p < .05.
+| |||||||||
+E5 AI practice wasn’t any more effortful than getting feedback from professional editors and was less effortful than looking for examples online
+Table S19 shows OLS models predicting practice effort metrics from practice condition.
+Results show that participants practicing without AI expended more effort, measured subjectively or objectively, through keystrokes or practice time.
+As pre-registered, time is square-root-transformed, and keystrokes are log-transformed.
+Differences are slightly smaller when using untransformed variables.
+| sqrt(Time) | log(Keystrokes) | Subjective Rating (0 - 10) | Time | Keystrokes | Keystrokes/min | |
+| Means | ||||||
+| AI | 2.45 | 3.76 | 7.14 | 7.04 | 251.19 | 1.52 |
+| (.112) | (.251) | (.217) | (.702) | (41.527) | (.089) | |
+| 2.76 | 4.68 | 7.11 | 8.96 | 369.45 | 1.78 | |
+| (.112) | (.251) | (.217) | (.702) | (41.464) | (.089) | |
+| Editors | 2.54 | 4.50 | 7.27 | 7.50 | 281.33 | 1.83 |
+| (.112) | (.251) | (.216) | (.702) | (41.482) | (.089) | |
+| Effect Sizes (d) | ||||||
+| AI vs. Google | .32*** | .42*** | -.02 | .31*** | .32*** | .34*** |
+| (.049) | (.049) | (.049) | (.049) | (.049) | (.050) | |
+| AI vs. Editors | .09 | .34*** | .07 | .07 | .08 | .41*** |
+| (.049) | (.049) | (.049) | (.049) | (.049) | (.049) | |
+| Google vs. Editors | -.22*** | -.08 | .09 | -.24*** | -.24*** | .06 |
+| (.049) | (.049) | (.049) | (.049) | (.049) | (.049) | |
+| Note. *** p < .001, ** p < .01, * p < .05. | ||||||
+Table S20 shows OLS models predicting test effort metrics from practice condition.
+Results show some differences: participants who practiced with AI pressed more keys but reported less subjective effort.
+| sqrt(Time) | log(Keystrokes) | Subjective Rating (0 - 10) | Time | Keystrokes | Keystrokes/min | |
+| Means | ||||||
+| AI | 2.32 | 5.44 | 7.95 | 5.55 | 415.78 | 1.92 |
+| (.056) | (.177) | (.225) | (.218) | (39.760) | (.041) | |
+| 2.23 | 5.29 | 7.52 | 5.20 | 384.82 | 1.92 | |
+| (.056) | (.177) | (.225) | (.218) | (39.747) | (.041) | |
+| Editors | 2.29 | 5.39 | 8.00 | 5.41 | 381.98 | 1.95 |
+| (.056) | (.177) | (.225) | (.218) | (39.793) | (.041) | |
+| Effect Sizes (d) | ||||||
+| AI vs. Google | -.18*** | -.09 | -.22*** | -.18*** | -.09 | -.02 |
+| (.049) | (.049) | (.049) | (.049) | (.049) | (.049) | |
+| AI vs. Editors | -.07 | -.03 | .02 | -.08 | -.10* | .07 |
+| (.049) | (.049) | (.049) | (.049) | (.049) | (.049) | |
+| Google vs. Editors | .11* | .06 | .24*** | .11* | -.01 | .09 |
+| (.049) | (.049) | (.049) | (.049) | (.049) | (.049) | |
+| Note. *** p < .001, ** p < .01, * p < .05. | ||||||
+E6 AI practice did not create more of an illusion of mastery than getting feedback from professional editors or looking for examples online
+As reported in the main text, there were no differences between participants who practiced with AI and participants who practiced with editor feedback on how much they thought they learned, their perceived skill, and their likelihood to ask for feedback after the test.
+See Table S21
+| Perceived Learning | Perceived Writing Skill | Asked for Feedback | |
+| Means | |||
+| AI | 6.56 | 6.74 | .67 |
+| (.227) | (.157) | (.054) | |
+| 6.01 | 6.59 | .59 | |
+| (.227) | (.157) | (.058) | |
+| Editors | 6.41 | 6.76 | .66 |
+| (.227) | (.156) | (.054) | |
+| Effect Sizes (d) | |||
+| AI vs. Google | -.28*** | -.11* | 1.38** |
+| (.049) | (.049) | (.147) | |
+| AI vs. Editors | -.07 | .02 | 1.03 |
+| (.049) | (.049) | (.112) | |
+| Google vs. Editors | .20*** | .12* | .75** |
+| (.049) | (.049) | (.080) | |
+| Note. *** p < .001, ** p < .01, * p < .05. | |||
+E7 Results were not moderated by individual differences
+As in Study 2, we tested whether each of the pretreatment demographic variables moderated the effects of seeing an AI example.
+To do this, we ran separate linear in which writing skill during the test phase was regressed on condition, the pre-treatment moderator of interest, writing skill at baseline, and an interaction term between the moderator condition.
+After correcting the p-values for the interaction terms, none were significant at the .05 level, suggesting that seeing AI examples was equally effective across groups.
+See Table S22
+| Test | Time Practice | Keys Practice | Effort Practice | Per. Learning | Per. Skill | Want Feedback | ||||||||
+| GE | EF | GE | EF | GE | EF | GE | EF | GE | EF | GE | EF | GE | EF | |
+| Pretest | 0.805 | 0.950 | 0.990 | 0.877 | 0.990 | 0.990 | 0.805 | 0.997 | 0.919 | 0.930 | 0.990 | 0.990 | 0.990 | 0.690 |
+| Year of Birth | 0.996 | 0.402 | 0.990 | 0.748 | 0.877 | 0.894 | 0.798 | 0.852 | 0.765 | 0.708 | 0.936 | 0.695 | 0.990 | 0.789 |
+| Writing Skill | 0.971 | 0.991 | 0.889 | 0.805 | 0.877 | 0.894 | 0.990 | 0.985 | 0.878 | 0.572 | 0.789 | 0.957 | 0.877 | 0.572 |
+| Gender vs. Female | ||||||||||||||
+| Male | 0.695 | 0.957 | 0.877 | 0.805 | 0.899 | 0.894 | 0.891 | 0.950 | 0.981 | 0.889 | 0.990 | 0.923 | 0.314 | 0.402 |
+| Other | 0.790 | 0.889 | 0.885 | 0.990 | 0.916 | 0.990 | 0.899 | 0.798 | 0.950 | 0.861 | 0.877 | 0.990 | 0.798 | 0.419 |
+| Race/Etnicity | ||||||||||||||
+| White | 0.267 | 0.402 | 0.402 | 0.542 | 0.950 | 0.877 | 0.878 | 0.887 | 0.862 | 0.877 | 0.805 | 0.990 | 0.852 | 0.449 |
+| Black | 0.748 | 0.922 | 0.748 | 0.990 | 0.891 | 0.877 | 0.442 | 0.211 | 0.189 | 0.533 | 0.140 | 0.427 | 0.990 | 0.990 |
+| Asian | 0.150 | 0.432 | 0.990 | 0.789 | 0.894 | 0.605 | 0.519 | 0.805 | 0.937 | 0.899 | 0.990 | 0.805 | 0.805 | 0.378 |
+| Latino | 0.189 | 0.789 | 0.402 | 0.889 | 0.318 | 0.894 | 0.981 | 0.605 | 0.899 | 0.805 | 0.852 | 0.789 | 0.990 | 0.990 |
+| Other | 0.894 | 0.798 | 0.877 | 0.990 | 0.990 | 0.805 | 0.959 | 0.660 | 0.805 | 0.701 | 0.877 | 0.427 | 0.990 | 0.877 |
+| Education | ||||||||||||||
+| High school | 0.990 | 0.990 | 0.882 | 0.877 | 0.990 | 0.874 | 0.990 | 0.419 | 0.998 | 0.189 | 0.140 | 0.166 | 0.990 | 0.990 |
+| Some college | 0.990 | 0.990 | 0.877 | 0.899 | 0.959 | 0.877 | 0.990 | 0.333 | 0.990 | 0.140 | 0.143 | 0.186 | 0.990 | 0.990 |
+| Associate | 0.990 | 0.990 | 0.899 | 0.690 | 0.990 | 0.798 | 0.990 | 0.542 | 0.990 | 0.140 | 0.144 | 0.166 | 0.990 | 0.990 |
+| Bachelor’s | 0.990 | 0.990 | 0.877 | 0.878 | 0.899 | 0.877 | 0.990 | 0.438 | 0.990 | 0.143 | 0.140 | 0.172 | 0.990 | 0.990 |
+| Master’s | 0.990 | 0.981 | 0.889 | 0.877 | 0.957 | 0.852 | 0.973 | 0.605 | 0.990 | 0.143 | 0.140 | 0.189 | 0.990 | 0.990 |
+| Doctoral/Professional | 0.990 | 0.894 | 0.894 | 0.805 | 0.990 | 0.789 | 0.990 | 0.402 | 0.927 | 0.140 | 0.140 | 0.140 | 0.990 | 0.990 |
+| Motivation (vs. Not at all) | ||||||||||||||
+| Hardly motivated | 0.990 | 0.990 | 0.611 | 0.768 | 0.877 | 0.899 | 0.690 | 0.889 | 0.901 | 0.885 | 0.577 | 0.990 | 0.798 | 0.990 |
+| Somewhat motivated | 0.996 | 0.957 | 0.605 | 0.805 | 0.732 | 0.854 | 0.278 | 0.542 | 0.838 | 0.990 | 0.877 | 0.877 | 0.877 | 0.934 |
+| Very motivated | 0.990 | 0.990 | 0.419 | 0.789 | 0.790 | 0.852 | 0.419 | 0.690 | 0.936 | 0.964 | 0.763 | 0.899 | 0.878 | 0.919 |
+| Extremely motivated | 0.990 | 0.936 | 0.547 | 0.789 | 0.690 | 0.789 | 0.448 | 0.608 | 0.894 | 0.990 | 0.660 | 0.894 | 0.789 | 0.874 |
+| Experience with AI writing assistants (vs. None) | ||||||||||||||
+| Hardly ever | 0.990 | 0.427 | 0.789 | 0.990 | 0.267 | 0.990 | 0.990 | 0.990 | 0.990 | 0.971 | 0.990 | 0.990 | 0.789 | 0.877 |
+| A few times per week | 0.805 | 0.877 | 0.789 | 0.990 | 0.660 | 0.990 | 0.930 | 0.990 | 0.889 | 0.444 | 0.970 | 0.990 | 0.899 | 0.877 |
+| About once a week | 0.805 | 0.798 | 0.278 | 0.790 | 0.189 | 0.990 | 0.990 | 0.990 | 0.981 | 0.990 | 0.990 | 0.985 | 0.981 | 0.990 |
+| Every day | 0.899 | 0.150 | 0.957 | 0.894 | 0.990 | 0.832 | 0.790 | 0.695 | 0.790 | 0.990 | 0.852 | 0.790 | 0.790 | 0.852 |
+| Note.
+Models for test performance, square-root practice time, log keystrokes, subjective effort, perceived learning and perceived writing skill or OLS models.
+Asking to see feedback was a binary Yes/No variable, and was modelled with logistic regression.
+Models match the pre-registered main specification, thereby controlling for all other pre-treatment variables.
+Per.
+= Perceived | ||||||||||||||
+E8 Pairwise comparisons
+The relative likelihood of a cover letter receiving an invitation to an interview was correlated with the GPT-rated writing quality. See Figure S9.
+Table S23 shows the beta regressions for the relative likelihood by condition.
+| As preregistered | All data | |||
+| Without controls | With controls | Without controls | With controls | |
+| Means (SE) | ||||
+| AI | .50 | .51 | .50 | .50 |
+| (.00) | (.01) | (.01) | (.01) | |
+| .49 | .49 | .49 | .49 | |
+| (.00) | (.01) | (.01) | (.01) | |
+| Editors | .51 | .51 | .51 | .51 |
+| (.00) | (.01) | (.01) | (.01) | |
+| Model coefficients (SE) | ||||
+| Precision () | 16.78*** | 7.17*** | 9.94*** | 9.94*** |
+| (1.28) | (.41) | (.74) | (.74) | |
+| Symmetry (log()) | .60*** | -.01 | -1.05*** | -1.05*** |
+| (.07) | (.06) | (.09) | (.09) | |
+| Pairwise comparison (SE) | ||||
+| AI - Google | .06* | .08* | .05 | .05 |
+| (.03) | (.04) | (.03) | (.03) | |
+| AI - Editors | -.01 | -.00 | -.04 | -.04 |
+| (.03) | (.04) | (.03) | (.03) | |
+| Google - Editors | -.06* | -.08* | -.09** | -.09** |
+| (.03) | (.04) | (.03) | (.03) | |
+| Statistics | ||||
+| N | 2575.00 | 2513.00 | 2575.00 | 2575.00 |
+| AIC | 3697.77 | 3750.64 | 867.58 | 867.58 |
+| BIC | 3727.04 | 3832.25 | 896.84 | 896.84 |
+| log(Likelihood) | -1843.89 | -1861.32 | -428.79 | -428.79 |
+Appendix F Results Study 5
+F1 Randomization, Balance, and Missingness
 As in Study 2, technical issues caused small amounts of missing data.
@@
 As shown in Table S24, missingness and attrition rates were low for the main and follow-up samples and did not differ by condition.
-| Condition | Main Sample | Followup Sample |
-| Practice wo AI | ||
-| Practice w AI | ||
+| Condition | Main Sample | Follow-up Sample |
+| Practice w/o AI | ||
+| Practice w/ AI | ||
 | See AI example | ||
@@
 See Table S25.
-| Overall | Practice wo AI | Practice w AI | See AI example | p | SMD | |
+| Overall | Practice w/o AI | Practice w/ AI | See AI example | p | SMD | |
 | n | 2003 | 672 | 652 | 679 | ||
@@
 | Pretest Writing Skill (mean (SD)) | 4.21 (0.88) | 4.23 (0.90) | 4.24 (0.90) | 4.17 (0.84) | .655 | 0.051 |
-C2 AI examples improve writing skill
+F2 AI examples improve writing skill
 The AI tool improved performance while participants used it.
@@
 For participants assigned to practice with the AI tool, The quality of AI rewrites did not correlate with participants’ final submissions, r = .06, p = .25.
-| GPT-4o | Claude | Ex. Controls | Ex. Cheaters | Followup | LM | ER | EN | F | ER | |
+| GPT-4o | Claude | Ex. Controls | Ex. Cheaters | Follow-up | LM | ER | EN | F | ER | |
 | Means — (SE) | ||||||||||
-| Practice wo AI | 4.72 | 4.96 | 4.62 | 4.61 | 4.64 | 4.33 | 6.61 | 5.43 | 2.93 | 4.31 |
+| Practice w/o AI | 4.72 | 4.96 | 4.62 | 4.61 | 4.64 | 4.33 | 6.61 | 5.43 | 2.93 | 4.31 |
 | (.374) | (.379) | (.048) | (.048) | (.651) | (.396) | (.265) | (.438) | (.795) | (.466) | |
-| Practice w AI | 6.19 | 6.36 | 6.08 | 6.08 | 5.98 | 5.53 | 7.44 | 6.60 | 5.64 | 5.74 |
+| Practice w/ AI | 6.19 | 6.36 | 6.08 | 6.08 | 5.98 | 5.53 | 7.44 | 6.60 | 5.64 | 5.74 |
 | (.373) | (.379) | (.048) | (.049) | (.656) | (.395) | (.265) | (.438) | (.794) | (.466) | |
@@
 | Effect Sizes (d) — (SE) | ||||||||||
-| Practice wo AI vs. Practice w AI | 1.22*** | 1.15*** | 1.19*** | 1.21*** | 1.14*** | .95*** | .96*** | .83*** | 1.06*** | .95*** |
+| Practice w/o AI vs. Practice w/ AI | 1.22*** | 1.15*** | 1.19*** | 1.21*** | 1.14*** | .95*** | .96*** | .83*** | 1.06*** | .95*** |
 | (.060) | (.059) | (.059) | (.060) | (.112) | (.059) | (.059) | (.058) | (.059) | (.059) | |
-| Practice wo AI vs. See AI example | 2.58*** | 2.52*** | 2.54*** | 2.55*** | 2.50*** | 2.28*** | 2.07*** | 1.95*** | 2.12*** | 1.81*** |
+| Practice w/o AI vs. See AI example | 2.58*** | 2.52*** | 2.54*** | 2.55*** | 2.50*** | 2.28*** | 2.07*** | 1.95*** | 2.12*** | 1.81*** |
 | (.070) | (.070) | (.069) | (.070) | (.132) | (.067) | (.066) | (.065) | (.066) | (.063) | |
-| Practice w AI vs. See AI example | 1.36*** | 1.37*** | 1.35*** | 1.35*** | 1.36*** | 1.33*** | 1.10*** | 1.12*** | 1.06*** | .86*** |
+| Practice w/ AI vs. See AI example | 1.36*** | 1.37*** | 1.35*** | 1.35*** | 1.36*** | 1.33*** | 1.10*** | 1.12*** | 1.06*** | .86*** |
 | (.061) | (.061) | (.060) | (.061) | (.113) | (.060) | (.059) | (.059) | (.059) | (.058) | |
@@
 Cheaters excludes the 3% of participants who admitted to cheating on the test phase.
-LM to ER are disagregated scores for each of the five principles.
+LM to ER are disaggregated scores for each of the five principles.
 LM = Less is More, ER = Easy Reading, EN = Easy Navigation, F = Formatting, ER = Easy Responding.
@@
 | ||||||||||
-| GPT-4o | Claude | Ex. Controls | Ex. Cheaters | Followup | LM | ER | EN | F | ER | |
+| GPT-4o | Claude | Ex. Controls | Ex. Cheaters | Follow-up | LM | ER | EN | F | ER | |
 | Means — (SE) | ||||||||||
-| Practice wo AI | 5.39 | 5.28 | 4.55 | 4.55 | 4.71 | 4.39 | 7.01 | 6.19 | 4.26 | 5.13 |
+| Practice w/o AI | 5.39 | 5.28 | 4.55 | 4.55 | 4.71 | 4.39 | 7.01 | 6.19 | 4.26 | 5.13 |
 | (.411) | (.426) | (.055) | (.055) | (.717) | (.391) | (.367) | (.479) | (.982) | (.496) | |
-| Practice w AI | 5.82 | 5.84 | 5.00 | 5.00 | 5.00 | 4.69 | 7.11 | 6.61 | 5.29 | 5.38 |
+| Practice w/ AI | 5.82 | 5.84 | 5.00 | 5.00 | 5.00 | 4.69 | 7.11 | 6.61 | 5.29 | 5.38 |
 | (.410) | (.426) | (.056) | (.056) | (.722) | (.390) | (.366) | (.478) | (.981) | (.495) | |
@@
 | Effect Sizes (d) — (SE) | ||||||||||
-| Practice wo AI vs. Practice w AI | .32*** | .41*** | .32*** | .33*** | .22* | .24*** | .09 | .28*** | .33*** | .16** |
+| Practice w/o AI vs. Practice w/ AI | .32*** | .41*** | .32*** | .33*** | .22* | .24*** | .09 | .28*** | .33*** | .16** |
 | (.057) | (.057) | (.056) | (.057) | (.106) | (.057) | (.057) | (.057) | (.057) | (.057) | |
-| Practice wo AI vs. See AI example | .36*** | .49*** | .34*** | .35*** | .29** | .22*** | .06 | .39*** | .38*** | .14** |
+| Practice w/o AI vs. See AI example | .36*** | .49*** | .34*** | .35*** | .29** | .22*** | .06 | .39*** | .38*** | .14** |
 | (.056) | (.056) | (.056) | (.057) | (.106) | (.056) | (.056) | (.056) | (.056) | (.056) | |
-| Practice w AI vs. See AI example | .04 | .08 | .01 | .02 | .06 | -.03 | -.02 | .11* | .06 | -.01 |
+| Practice w/ AI vs. See AI example | .04 | .08 | .01 | .02 | .06 | -.03 | -.02 | .11* | .06 | -.01 |
 | (.056) | (.056) | (.056) | (.057) | (.104) | (.056) | (.056) | (.056) | (.056) | (.056) | |
@@
 Cheaters excludes the 3% of participants who admitted to cheating on the test phase.
-LM to ER are disagregated scores for each of the five principles.
+LM to ER are disaggregated scores for each of the five principles.
 LM = Less is More, ER = Easy Reading, EN = Easy Navigation, F = Formatting, ER = Easy Responding.
@@
 | ||||||||||
-C3 Seeing AI examples was less effortful
+F3 Seeing AI examples was less effortful
 Table S28 shows OLS models predicting practice effort metrics from practice condition.
@@
 Differences are slightly smaller when using untransformed variables.
-| sqrt(Time) | log(Keystrokes) | Subjective Rating (0 - 10) | Time | Keystrokes | |
-| Means — (SE) | |||||
-| Practice wo AI | 2.83 | 5.01 | 6.17 | 9.00 | 259.34 |
-| (.270) | (.541) | (.642) | (1.495) | (99.161) | |
-| Practice w AI | 2.71 | 4.05 | 5.89 | 8.65 | 228.45 |
-| (.270) | (.540) | (.641) | (1.493) | (99.050) | |
-| See AI example | 1.85 | .81 | 5.52 | 4.99 | 24.98 |
-| (.271) | (.542) | (.643) | (1.499) | (99.392) | |
-| Effect Sizes (d) — (SE) | |||||
-| Practice wo AI vs. Practice w AI | -.14* | -.55*** | -.14* | -.07 | -.10 |
-| (.056) | (.056) | (.057) | (.056) | (.056) | |
-| Practice wo AI vs. See AI example | -1.13*** | -2.41*** | -.32*** | -.83*** | -.73*** |
-| (.059) | (.067) | (.056) | (.057) | (.056) | |
-| Practice w AI vs. See AI example | -.99*** | -1.86*** | -.18** | -.76*** | -.64*** |
-| (.058) | (.063) | (.056) | (.057) | (.056) | |
-| Note. *** p < .001, ** p < .01, * p < .05. | |||||
+| sqrt(Time) | log(Keystrokes) | Subjective Rating (0 - 10) | Time | Keystrokes | Keystrokes/min | |
+| Means | ||||||
+| Practice wo AI | 2.83 | 5.01 | 6.17 | 9.00 | 259.34 | 27.62 |
+| (.270) | (.541) | (.642) | (1.495) | (99.161) | (13.213) | |
+| Practice w AI | 2.71 | 4.05 | 5.89 | 8.65 | 228.45 | 21.20 |
+| (.270) | (.540) | (.641) | (1.493) | (99.050) | (13.200) | |
+| See AI example | 1.85 | .81 | 5.52 | 4.99 | 24.98 | -8.17 |
+| (.271) | (.542) | (.643) | (1.499) | (99.392) | (13.246) | |
+| Effect Sizes (d) | ||||||
+| Practice wo AI vs. Practice w AI | -.14* | -.55*** | -.14* | -.07 | -.10 | -.15** |
+| (.056) | (.056) | (.057) | (.056) | (.056) | (.056) | |
+| Practice wo AI vs. See AI example | -1.13*** | -2.41*** | -.32*** | -.83*** | -.73*** | -.84*** |
+| (.059) | (.067) | (.056) | (.057) | (.056) | (.056) | |
+| Practice w AI vs. See AI example | -.99*** | -1.86*** | -.18** | -.76*** | -.64*** | -.69*** |
+| (.058) | (.063) | (.056) | (.057) | (.056) | (.056) | |
+| Note. *** p < .001, ** p < .01, * p < .05. | ||||||
 Table S29 shows OLS models predicting test effort metrics from practice condition.
 Results show some differences: participants who had seen the AI example write for longer duing the test, and pressed more keys, however their subjective experience of effort was not different from those who practice with or without the AI tool.
-| sqrt(Time) | log(Keystrokes) | Subjective Rating (0 - 10) | Time | Keystrokes | |
-| Means — (SE) | |||||
-| Practice wo AI | 2.45 | 5.40 | 7.05 | 6.14 | 432.01 |
-| (.170) | (.562) | (.623) | (.655) | (109.856) | |
-| Practice w AI | 2.50 | 5.52 | 7.19 | 6.35 | 466.90 |
-| (.169) | (.561) | (.622) | (.654) | (109.733) | |
-| See AI example | 2.57 | 5.86 | 7.21 | 6.63 | 517.26 |
-| (.170) | (.563) | (.625) | (.656) | (110.112) | |
-| Effect Sizes (d) — (SE) | |||||
-| Practice wo AI vs. Practice w AI | .09 | .06 | .07 | .10 | .10 |
-| (.057) | (.056) | (.057) | (.057) | (.056) | |
-| Practice wo AI vs. See AI example | .22*** | .25*** | .08 | .23*** | .24*** |
-| (.056) | (.055) | (.056) | (.056) | (.055) | |
-| Practice w AI vs. See AI example | .13* | .19*** | .01 | .13* | .14* |
-| (.056) | (.056) | (.056) | (.056) | (.055) | |
-| Note. *** p < .001, ** p < .01, * p < .05. | |||||
+| sqrt(Time) | log(Keystrokes) | Subjective Rating (0 - 10) | Time | Keystrokes | Keystrokes/min | |
+| Means | ||||||
+| Practice wo AI | 2.45 | 5.40 | 7.05 | 6.14 | 432.01 | 68.46 |
+| (.170) | (.562) | (.623) | (.655) | (109.856) | (19.117) | |
+| Practice w AI | 2.50 | 5.52 | 7.19 | 6.35 | 466.90 | 72.19 |
+| (.169) | (.561) | (.622) | (.654) | (109.733) | (19.090) | |
+| See AI example | 2.57 | 5.86 | 7.21 | 6.63 | 517.26 | 77.84 |
+| (.170) | (.563) | (.625) | (.656) | (110.112) | (19.159) | |
+| Effect Sizes (d) | ||||||
+| Practice wo AI vs. Practice w AI | .09 | .06 | .07 | .10 | .10 | .06 |
+| (.057) | (.056) | (.057) | (.057) | (.056) | (.057) | |
+| Practice wo AI vs. See AI example | .22*** | .25*** | .08 | .23*** | .24*** | .15** |
+| (.056) | (.055) | (.056) | (.056) | (.055) | (.056) | |
+| Practice w AI vs. See AI example | .13* | .19*** | .01 | .13* | .14* | .09 |
+| (.056) | (.056) | (.056) | (.056) | (.055) | (.056) | |
+| Note. *** p < .001, ** p < .01, * p < .05. | ||||||
 Table S30 shows OLS models predicting learning rate metrics from practice condition.
@@
 Participants who had seen an AI example improved their skill more efficiently.
-| sqrt(Time) | log(Keystrokes) | Subjective Rating (0 - 10) | Time | Keystrokes | |
-| Means — (SE) | |||||
-| Practice wo AI | .20 | .15 | .18 | .31 | .34 |
-| (.155) | (.272) | (.085) | (.278) | (.167) | |
-| Practice w AI | .30 | .22 | .25 | .45 | .49 |
-| (.155) | (.272) | (.085) | (.277) | (.167) | |
-| See AI example | .51 | .97 | .28 | 1.08 | .62 |
-| (.155) | (.273) | (.085) | (.278) | (.168) | |
-| Effect Sizes (d) — (SE) | |||||
-| Practice wo AI vs. Practice w AI | .21*** | .08 | .25*** | .15** | .27*** |
-| (.057) | (.057) | (.057) | (.057) | (.057) | |
-| Practice wo AI vs. See AI example | .62*** | .94*** | .38*** | .86*** | .53*** |
-| (.057) | (.058) | (.056) | (.058) | (.057) | |
-| Practice w AI vs. See AI example | .41*** | .86*** | .13* | .71*** | .25*** |
-| (.057) | (.058) | (.056) | (.057) | (.056) | |
-| Note. *** p < .001, ** p < .01, * p < .05. | |||||
-C4 Seeing an AI example did not discourage motivation for future learning
+| sqrt(Time) | log(Keystrokes) | Subjective Rating (0 - 10) | Time | Keystrokes | Keystrokes/min | |
+| Means | ||||||
+| Practice wo AI | .20 | .15 | .18 | .31 | .34 | .20 |
+| (.155) | (.272) | (.085) | (.278) | (.167) | (.277) | |
+| Practice w AI | .30 | .22 | .25 | .45 | .49 | .30 |
+| (.155) | (.272) | (.085) | (.277) | (.167) | (.277) | |
+| See AI example | .51 | .97 | .28 | 1.08 | .62 | 1.00 |
+| (.155) | (.273) | (.085) | (.278) | (.168) | (.278) | |
+| Effect Sizes (d) | ||||||
+| Practice wo AI vs. Practice w AI | .21*** | .08 | .25*** | .15** | .27*** | .12* |
+| (.057) | (.057) | (.057) | (.057) | (.057) | (.057) | |
+| Practice wo AI vs. See AI example | .62*** | .94*** | .38*** | .86*** | .53*** | .90*** |
+| (.057) | (.058) | (.056) | (.058) | (.057) | (.058) | |
+| Practice w AI vs. See AI example | .41*** | .86*** | .13* | .71*** | .25*** | .78*** |
+| (.057) | (.058) | (.056) | (.057) | (.056) | (.058) | |
+| Note. *** p < .001, ** p < .01, * p < .05. | ||||||
+F4 Seeing an AI example did not discourage motivation for future learning
 Table S31 presents differences in perceived learning, perceived writing skill, and the likelihood of asking for feedback across conditions, with effect sizes and means reported for each comparison.
@@
 | Means — (SE) | |||
-| Practice wo AI | 5.26 | 6.33 | .64 |
+| Practice w/o AI | 5.26 | 6.33 | .64 |
 | (.549) | (.517) | (.670) | |
-| Practice w AI | 5.25 | 6.36 | .46 |
+| Practice w/ AI | 5.25 | 6.36 | .46 |
 | (.549) | (.516) | (.669) | |
@@
 | (.551) | (.518) | (.671) | |
-| extbfEffect Sizes (d) | |||
-| Practice wo AI vs. Practice w AI | -.01 | .02 | 1.19 |
+| Effect Sizes (d) | |||
+| Practice w/o AI vs. Practice w/ AI | -.01 | .02 | 1.19 |
 | (.057) | (.057) | (.122) | |
-| Practice wo AI vs. See AI example | .09 | -.06 | 1.10 |
+| Practice w/o AI vs. See AI example | .09 | -.06 | 1.10 |
 | (.056) | (.056) | (.121) | |
-| Practice w AI vs. See AI example | .09 | -.08 | 0.921 |
+| Practice w/ AI vs. See AI example | .09 | -.08 | 0.921 |
 | (.056) | (.056) | (.120) | |
 | Note. *** p < .001, ** p < .01, * p < .05. | |||
-C5 The effects of seeing an AI example persist
+F5 The benefits of seeing an AI example were just as large a day later
 Table S32 shows means and standardized differences for measures of writing skill and related outcomes during the follow-up phase.
@@
 | Means — (SE) | |||||||||
-| Practice wo AI | 4.95 | 5.10 | 4.87 | 4.88 | 5.32 | 6.83 | 5.40 | 1.98 | 5.23 |
+| Practice w/o AI | 4.95 | 5.10 | 4.87 | 4.88 | 5.32 | 6.83 | 5.40 | 1.98 | 5.23 |
 | (.776) | (.798) | (.109) | (.110) | (.750) | (.730) | (.847) | (1.829) | (.905) | |
-| Practice w AI | 5.37 | 5.67 | 5.34 | 5.38 | 5.57 | 6.98 | 5.76 | 2.91 | 5.61 |
+| Practice w/ AI | 5.37 | 5.67 | 5.34 | 5.38 | 5.57 | 6.98 | 5.76 | 2.91 | 5.61 |
 | (.781) | (.804) | (.103) | (.105) | (.756) | (.735) | (.853) | (1.842) | (.912) | |
@@
 | Effect Sizes (d) — (SE) | |||||||||
-| Practice wo AI vs. Practice w AI | .29** | .40*** | .32** | .34** | .18 | .11 | .24* | .28** | .23* |
+| Practice w/o AI vs. Practice w/ AI | .29** | .40*** | .32** | .34** | .18 | .11 | .24* | .28** | .23* |
 | (.106) | (.107) | (.104) | (.106) | (.106) | (.106) | (.106) | (.106) | (.106) | |
-| Practice wo AI vs. See AI example | .32** | .43*** | .35*** | .33** | .16 | .11 | .31** | .35*** | .14 |
+| Practice w/o AI vs. See AI example | .32** | .43*** | .35*** | .33** | .16 | .11 | .31** | .35*** | .14 |
 | (.106) | (.107) | (.104) | (.106) | (.106) | (.106) | (.106) | (.107) | (.106) | |
-| Practice w AI vs. See AI example | .02 | .03 | .02 | -.01 | -.02 | -.00 | .07 | .07 | -.09 |
+| Practice w/ AI vs. See AI example | .02 | .03 | .02 | -.01 | -.02 | -.00 | .07 | .07 | -.09 |
 | (.104) | (.104) | (.101) | (.103) | (.104) | (.104) | (.104) | (.104) | (.104) | |
@@
 Cheaters excludes the 3% of participants who admitted to cheating on the test phase.
-LM to ER are disagregated scores for each of the five principles.
+LM to ER are disaggregated scores for each of the five principles.
 LM = Less is More, ER = Easy Reading, EN = Easy Navigation, F = Formatting, ER = Easy Responding.
@@
 | Means — (SE) | ||||
-| Practice wo AI | 4.95 | 6.02 | 5.47 | 4.85 |
+| Practice w/o AI | 4.95 | 6.02 | 5.47 | 4.85 |
 | (.776) | (1.539) | (.492) | (.811) | |
-| Practice w AI | 5.37 | 6.64 | 5.57 | 5.41 |
+| Practice w/ AI | 5.37 | 6.64 | 5.57 | 5.41 |
 | (.781) | (1.446) | (.483) | (.817) | |
@@
 | Effect Sizes (d) — (SE) | ||||
-| Practice wo AI vs. Practice w AI | .29** | .43 | .07 | .40** |
+| Practice w/o AI vs. Practice w/ AI | .29** | .43 | .07 | .40** |
 | (.106) | (.387) | (.187) | (.147) | |
-| Practice wo AI vs. See AI example | .32** | .34 | .29 | .34* |
+| Practice w/o AI vs. See AI example | .32** | .34 | .29 | .34* |
 | (.106) | (.354) | (.190) | (.149) | |
-| Practice w AI vs. See AI example | .02 | -.09 | .22 | -.06 |
+| Practice w/ AI vs. See AI example | .02 | -.09 | .22 | -.06 |
 | (.104) | (.344) | (.177) | (.145) | |
@@
 Cheaters excludes the 3% of participants who admitted to cheating on the test phase.
-LM to ER are disagregated scores for each of the five principles.
+LM to ER are disaggregated scores for each of the five principles.
 LM = Less is More, ER = Easy Reading, EN = Easy Navigation, F = Formatting, ER = Easy Responding.
@@
 | ||||
-C6 Seeing AI examples was equally effective across subgroups
+As mentioned in the main text the effect of practicing with AI and seeing an AI example did not become attenuated one day later. See Table S34.
+| Term | Estimate | SE | t | p-value |
+| Intercept | -8.469 | 4.502 | -1.881 | 0.060 |
+| Condition: Practice w AI | 0.421 | 0.076 | 5.551 | 0.000 |
+| Condition: See AI example | 0.467 | 0.075 | 6.253 | 0.000 |
+| Condition: Practice w/ AI Follow-up | -0.007 | 0.158 | -0.044 | 0.965 |
+| Condition: See AI example Follow-up | -0.006 | 0.157 | -0.036 | 0.971 |
+| Phase: Follow-Up | 0.356 | 0.114 | 3.139 | 0.002 |
+| Pretest score | 0.394 | 0.031 | 12.877 | 0.000 |
+| Year of birth | 0.006 | 0.002 | 2.524 | 0.012 |
+| Gender: Male | -0.096 | 0.055 | -1.724 | 0.085 |
+| Gender: Other | 0.117 | 0.249 | 0.470 | 0.639 |
+| Race/ethnicity: White | 0.215 | 0.117 | 1.846 | 0.065 |
+| Race/ethnicity: Black | -0.140 | 0.125 | -1.125 | 0.261 |
+| Race/ethnicity: Asian | 0.575 | 0.148 | 3.880 | 0.000 |
+| Race/ethnicity: Latino | 0.216 | 0.123 | 1.756 | 0.079 |
+| Race/ethnicity: Other | -0.180 | 0.250 | -0.719 | 0.472 |
+| Education: High school graduate (high school diploma or equivalent including GED) | -0.184 | 0.397 | -0.464 | 0.642 |
+| Education: Some college but no degree | -0.187 | 0.394 | -0.475 | 0.635 |
+| Education: Associate degree in college (2-year) | -0.207 | 0.399 | -0.518 | 0.605 |
+| Education: Bachelor’s degree in college (4-year) | -0.121 | 0.390 | -0.310 | 0.756 |
+| Education: Master’s degree | -0.033 | 0.393 | -0.085 | 0.933 |
+| Education: Doctoral degree (PhD) | -0.227 | 0.439 | -0.518 | 0.604 |
+| Education: Non-PhD Professional degree (JD, MD) | 0.099 | 0.456 | 0.218 | 0.828 |
+| Writing skill | 0.045 | 0.017 | 2.578 | 0.010 |
+| Motivation: Hardly motivated | -0.196 | 0.258 | -0.760 | 0.447 |
+| Motivation: Somewhat motivated | -0.142 | 0.243 | -0.582 | 0.561 |
+| Motivation: Very motivated | -0.338 | 0.244 | -1.388 | 0.165 |
+| Motivation: Extremely motivated | -0.430 | 0.249 | -1.731 | 0.084 |
+| Experience: I have tried AI writing assistant(s) but hardly ever use them | 0.132 | 0.079 | 1.678 | 0.093 |
+| Experience: I use AI writing assistant(s) a few times per week | 0.003 | 0.094 | 0.034 | 0.973 |
+| Experience: I use AI writing assistant(s) about once a week | 0.070 | 0.095 | 0.730 | 0.465 |
+| Experience: I use AI writing assistant(s) every day | -0.063 | 0.131 | -0.480 | 0.631 |
+F6 Seeing AI examples was equally effective across subgroups
 As in Study 2, we tested whether each of the pretreatment demographic variables moderated the effects of seeing an AI example.
@@
 | ||||||||||||||||
+F7 Pairwise Comparisons
+The relative likelihood of a cover letter receiving an invitation to an interview was correlated with the GPT-rated writing quality. See Figure S11.
+As shown in Table S36, participants who had practiced writing cover letters with AI were more likely to be invited to a hypothetical job interview.
+| Without controls | With controls | |
+| Means (SE) | ||
+| Practice w AI | .50 | .51 |
+| (.01) | (.01) | |
+| Practice wo AI | .48 | .48 |
+| (.01) | (.01) | |
+| See AI example | .51 | .51 |
+| (.01) | (.01) | |
+| Model coefficients (SE) | ||
+| Precision () | 16.17*** | 7.77*** |
+| (1.47) | (.51) | |
+| Symmetry (Log()) | .35*** | -.23*** |
+| (.08) | (.07) | |
+| Pairwise comparison (SE) | ||
+| Practice w AI - Practice wo AI | .08** | .09* |
+| (.03) | (.04) | |
+| Practice w AI - See AI example | -.02 | -.03 |
+| (.03) | (.04) | |
+| Practice wo AI - See AI example | -.10*** | -.12** |
+| (.03) | (.04) | |
+| Statistics | ||
+| N | 1934.00 | 1916.00 |
+| AIC | 2447.07 | 2518.06 |
+| BIC | 2474.91 | 2684.80 |
+| log(Likelihood) | -1218.54 | -1229.03 |

### v4

@@
-Coach not crutch: AI assistance can enhance rather than hinder skill development
+Coach not crutch: Evidence that AI can improve writing skill despite reducing effort
 Abstract
-Does using AI make you stupid?
-Contrary to predictions by lay forecasters, participants randomly assigned to practice writing with access to an AI tool improved more on a writing test one day later compared to writers assigned to practice without AI—despite exerting less effort.
-Likewise, whereas forecasters predicted that AI would be less helpful, and less worth paying for, than personalized feedback from experienced human editors, we found the converse: practicing with AI increased skill more than feedback from human editors (or simply googling examples and tips).
-The results of a third experiment suggest that AI may teach by example: writers merely shown an AI-generated cover letter (without the opportunity to practice) performed as well as writers who practiced writing with the original AI tool.
-Collectively, the findings of these pre-registered studies constitute an existence proof that by providing personalized, high-quality examples, AI can improve, rather than undermine, learning.
+In a series of highly-powered empirical studies, we examine the intuition that by sparing effort, using AI inevitably hinders learning.
+First, in a nationally representative survey of young adults, the majority expressed the view that using AI makes people lazier and less capable.
+Next, in a random-assignment experiment, we gave participants a tutorial on best practices in professional writing, then provided one group with access to an AI writing tool and asked another to practice writing on their own.
+Those who practiced with AI indeed exerted less effort while practicing—yet wrote better cover letters in no-AI writing tests.
+In a second experiment with more rigorous control conditions, access to AI improved writing more than either googling cover letter examples and tips or receiving personalized feedback on their practice letters from experienced human editors.
+A third experiment explained these learning gains by showing that AI can teach by example: participants who viewed a cover letter that had been revised by the AI tool but did no further practice improved their writing as much as those who practiced writing with the original AI tool.
+Collectively, these pre-registered experiments suggest that AI can exert opposing effects on effort and learning rate—making it possible in at least some cases to work less and learn more.
 keywords:
-[…]
-Generative AI (henceforth AI) tools are increasingly powerful and prevalent (1), and there is mounting evidence that they can dramatically boost performance.
-For example, working side-by-side with AI as a copilot has been shown to increase both quality and speed in a variety of professional writing tasks (e.g., emails, memos, short reports) (2, 3, 4).
-Nevertheless, there is growing concern that AI tools will be used as a crutch, providing immediate gains in performance at the expense of long-run development of human capital (5, 6, 7).
-For instance, in a 2024 poll, 62% of surveyed adults predicted that Generative AI would “lead to humans becoming less intelligent” (8).
-In January 2023, New York City public schools banned ChatGPT, citing “concerns about negative impacts on student learning (9).” When this ban was lifted three months later, it was not because of AI’s potential to scaffold learning, but instead because of the “reality that students are participating in and will work in a world where understanding Generative AI is crucial (10).” However, teachers remain worried: in a nationally representative poll in May 2024, four times as many K-12 educators judged the use of AI tools as net harmful (24%) than net beneficial (6%) (11).
-Concerns that using AI tools hinders learning (while increasing short-term performance) are justified for at least three reasons.
-First, AI systems based on large language models like GPT-4 have been shown to confidently assert erroneous facts (i.e., hallucinations (12)), make reasoning and arithmetic errors (13), and complete other tasks with varying degrees of accuracy.
-Second, the fluent and instantaneous solutions AI tools generate may contribute to an illusion of mastery.
-To the extent users conflate the skills of an AI tool with their own, they may be less likely to seek feedback and improve.
-Prior research has found that searching for information on the Internet, for example, creates an illusion whereby people conflate knowledge outside their heads with what they personally know themselves (14).
-Third, technological tools reduce the learner’s need to be cognitively engaged with the task at hand.
-For instance, knowing that one can search for a fact on a computer has been shown to reduce memory for that fact, instead encouraging recall of how to search for it (15).
-And drivers who use GPS tend to have worse hippocampal-dependent spatial memory, both cross-sectionally and longitudinally (16, 17).
-When AI-powered tools instantaneously produce turn-key solutions for complex cognitive tasks, they may be especially detrimental to learning.
-It is, after all, tempting to copy and paste the output of an AI tool without even laying eyes on it.
-On the other hand, using AI tools to assist us in our work may, in fact, help us develop our own skills.
-In particular, the current generation of AI tools may teach by example, offering high-quality and personally tailored demonstrations of abstract principles that are otherwise difficult to grasp.
-Classic research shows that worked examples of math problems (i.e., not just answers but the step-by-step process by which problems are solved) scaffold learning more effectively than explanations alone (18, 19).
-It seems likely that, compared to textbooks, conventional computer tutoring programs (e.g., (20)), and human teachers, AI tools may provide higher quality, just-in-time examples exquisitely tailored to a learner’s specific needs.
-Indeed, it’s possible that in both quality and didactic utility, AI examples (e.g., of an excellent cover letter) might surpass even those provided by domain experts (e.g., professional writers and editors).
-Thus, if the upside of exposure to excellent examples tailored to the learners’ needs outweighs the downside of diminished engagement, AI tools may support skill development.
-The scant research on how AI tools influence skill development has yielded mixed results.
-Some studies report that interacting with AI tools improves skill on subsequent tests in which AI tools are not available (21, 22), while others have shown null or even negative effects (23, 24, 22).
-Notably, these studies examine AI tutors, chatbots, or explanations explicitly designed to support learning, rather than simply providing solutions, as is typical in real-world use.
-Further, they focused exclusively on mathematics and computer programming.
-In this investigation, we ask whether AI tools can support skill development, above and beyond merely improving performance while using the AI tool.
-We focus on writing—the most common use of AI at work, as ranked in a nationally representative survey of American adults in August 2024 (26).
-In our experimental paradigm all participants are given a baseline writing test (i.e., revising a poorly written cover letter), followed by a short lesson on five evidence-based strategies for professional writing (25, 27).
-Next, participants are randomly assigned to different types of practice (e.g., practice with access to an AI tool, practice without access to an AI tool, practice with feedback from professional editors, and practice with Google Search to find examples and tips).
-Finally, all participants complete an incentivized test in which they rewrite cover letters without access to AI.
-In Study 1, forecasters presented with this design were twice as likely to predict that practicing writing with the assistance of AI would impair learning compared to practicing without it.
-In Study 2, however, participants who practiced with AI actually learned more (i.e., wrote better cover letters during the test phase) than participants who practiced without AI—an advantage that persisted in a one-day follow-up test.
-In Study 3, forecasters predicted that feedback from AI would be less helpful than feedback from professional editors and, accordingly, were less willing to pay for it.
-However, in Study 4, participants who practiced with AI learned more than participants who practiced with feedback from professional editors (or those who practiced with Google Search to find examples and tips).
-Finally, in Study 5, we explored the mechanism for these learning gains by introducing an example-only condition.
-Participants who saw an AI-generated example (but did not have an opportunity to practice) improved their writing as much as participants who practiced with AI; benefits again persisted in a one-day follow-up test.
-Study 1: Forecasting the effect of AI on learning
-We showed N = 150 participants screenshots of a random assignment study with three conditions.
-We asked them to rank-order these conditions according to how much they predicted future participants would learn in each.
-Lay forecasters predicted that practicing with AI would hinder learning
-Confirming our pre-registered hypothesis, nearly twice as many forecasters (64.7%) ranked practicing alone above practicing writing with access to an AI tool as the converse (35.3%, (1) = 12.9, p < .001), see Figure 3.
-Participants made this prediction regardless of self-reported experience with AI (OR = 0.72, p = .096) or any other measured demographic characteristic (ps > .05).
-See Table S2
-In open-ended responses, forecasters who were pessimistic about the effect of the AI tool on learning speculated that it would crowd out effort (e.g., “Practicing alone would force more recall and problem-solving skills, while AI essentially just gives them the answer.”, “I think oftentimes using AI impedes the learning process because it’s the ‘easy way.”’).
-Those with positive views, on the other hand, cited the possibility of AI providing insights or examples that would be otherwise unavailable (“As much as I hate AI, I do not believe you can improve in any manner if you do not have examples or other ways of learning, and AI can provide this.”)
+Human Capital Development | Generative AI | Skill AcquisitionThis manuscript was compiled on February 18, 2026
+Generative AI (henceforth AI) tools are increasingly powerful and prevalent (bubeck2023), and there is mounting evidence that they can dramatically boost performance.
+For example, working side-by-side with AI as a copilot has been shown to increase both quality and speed in a variety of professional writing tasks (e.g., emails, memos, short reports) (noy2023, dellacqua2023, wiles2023algorithmic).
+Nevertheless, advances in AI in recent years have been accompanied by concerns that immediate gains in performance are offset by the subsequent erosion of human capital (hofman2023sports, puntoni2021, borges2024).
+For example, in a 2024 poll of American adults, 62% predicted that generative AI would “lead to humans becoming less intelligent” (hawkins2024between).
+In a 2024 survey of K–12 educators, four times as many judged AI to be net harmful (24%) as net beneficial (6%) (lin2024a).
+Concerns that AI hinders learning are justified for at least three reasons.
+First, AI sometimes confidently asserts erroneous facts (i.e., hallucinations (ji2023survey, yuan2023well)).
+Second, using AI may create an illusion of mastery (fisher2015), particularly if users conflate its skilled responses with their own.
+Third and most important, using AI may reduce human effort because the more AI solves your problems, the less you have to think about them yourself (kosmyna2025your, bastani2024).
+All things being equal, less effort results in less learning (bjork2011making, james1983talks).
+Analogous deskilling effects have been documented with prior technologies.
+In cross-sectional and longitudinal studies, drivers who rely on GPS exhibit poorer hippocampal-dependent spatial memory (maguire2000, griesbauer2022london).
+Likewise, knowing that information can be easily retrieved from the internet reduces the likelihood that one will commit it to memory.
+(sparrow2011).
+AI may be uniquely harmful because, unlike search engines or GPS, it can replace the process of thinking itself—not just the storage or retrieval of facts.
+The assumption that diminished effort inevitably reduces learning is shaping the design of real-world AI learning tools.
+For example, when Khan Academy released its AI tutor Khanmigo, it announced on its website that it would prevent students from seeing full solutions: “Khanmigo challenges you to think critically and solve problems without giving you direct answers111emphasis added” (khanmigo2025).
+Skill acquisition, however, depends on more than effort; learning rate also matters (duckworth2015mechanics).
+A wide array of factors have been shown to influence how quickly students learn in conventional educational settings (wisniewski_power_2020, bloom_2_1984, sana_interleaving_2022).
+One feature of optimal learning environments with special relevance to the affordances of generative AI is the availability of specific examples that illustrate abstract principles that are otherwise difficult for the learner to grasp and apply.
+Separate from its influence on effort, therefore, AI might change the quality of the learning environment.
+See Figure 1.
+The net effect of effort and learning environment on skill acquisition remains an open empirical question.
+Here, we focus on writing—the most common use of AI at work.
+In a nationally representative sample of American adults in 2024, writing tasks were the most frequent workplace application (bick2024rapid).
+This pattern is corroborated by large-scale usage data: In an analysis of over 1 million ChatGPT conversations writing constituted 24% of all messages and 40% of work-related messages, making it the dominant occupational use case (chatterji_how_nodate).
+Notably, approximately two-thirds of writing-related queries involved modifying existing text rather than generating novel content.
+We know of only one study (kosmyna2025your) that has experimentally manipulated access to AI (vs.
+no access) during a writing task.
+Fifty-four undergraduates at four selective universities were randomly assigned to condition and then asked to write essays: 18 participants wrote with access to AI, 18 wrote without access to AI, and 18 wrote with access to Google Search.
+Although a no-AI test of writing skill was not administered, compared to the other two conditions, participants given access to AI were less likely to recall a verbatim sentence from their essays immediately after writing them.
+We developed an experimental paradigm to determine whether using AI as a productivity tool for writing cover letters enhances or erodes subsequent writing skill.
+To begin, all participants complete a baseline questionnaire and writing pretest (i.e., revising a poorly written cover letter).
+Next, as shown in Figure 4, all participants are taught five evidence-based principles for professional writing (i.e., Less is more, Make reading easy, Design for easy navigation, Use enough formatting but no more, Make responding easy) (rogers2023, shulman2024) and are shown a domain-general example for each principle.
+After this lesson, participants are randomly assigned to different types of practice (e.g., practice with access to an AI tool, practice without access to an AI tool, practice with feedback from professional editors, and practice with Google Search to find examples and tips).
+Finally, all participants rewrite a new cover letter without access to AI (and with copy-pasting functionality disabled).
+Each participant’s submitted cover letter is rated on adherence to each of the five writing principles, respectively.
+To gauge the ecological validity of the resulting composite index, we recruited a separate sample of raters (blind to condition) who read randomly chosen pairs of cover letters from the main experiment and were asked to “choose the cover letter that would make you more likely to offer an interview to the candidate.” See Methods for details.
+In this investigation, we examine contemporary assumptions about learning and AI and empirically test the accuracy of these beliefs.
+In Study 1, a nationally representative survey of young adults in the U.S.
+in October 2025, the majority endorsed the view that using AI makes people lazier, less smart, and less educated.
+In Study 2, participants who practiced writing cover letters with AI actually learned more than participants who practiced without AI—an advantage that persisted in a one-day follow-up test.
+A separate sample of lay forecasters predicted the opposite result.
+In Study 3, again against forecaster predictions, participants who practiced with AI learned more than participants who practiced with feedback from professional editors (or those who practiced with Google Search to find examples and tips).
+Finally, in Study 4, we explored the mechanism for these learning gains by introducing an example-only condition.
+Participants who viewed an AI-generated example (but did not have an opportunity to edit it) improved their writing as much as participants who practiced with AI; benefits again persisted in a one-day follow-up test.
+Study 1: Public perceptions of AI and learning
+In partnership with the Gallup Organization, we asked a nationally representative sample of American adults aged 18 to 28; (N = 2,637) in October 2025 about AI use in the past month and attitudes about AI in general.
+Consistent with 2024 polls, writing is the most common use case for AI (45% of all adults surveyed) after using AI as a replacement for Google (64%) and ahead of generating images (35%) or help with math (23%).
+In general, AI is used more commonly for work (53%) than for personal reasons like life advice (28%) or companionship (17%).
+Participants expressed concern that “AI makes people lazier” (79%, One sample d = 0.66, p < .001) and, to a more moderate degree, that “AI makes people learn less” (62%, One sample d = 0.34, p < .001) and “AI makes people less smart” (63%, One sample d = 0.30, p < .001).
+See Figure 5.
+See Section B in the Supplementary Information for details.
 Study 2: Effect of practicing with AI on learning
-Study 2 tested whether the predictions of Study 1 forecasters were accurate.
-Specifically, N = 2,238 participants completed a baseline questionnaire and pretest (rewriting a poorly written cover letter), followed by a lesson introducing five principles of effective writing (i.e., Less is more, Make reading easy, Design for easy navigation, Use enough formatting but no more, Make responding easy) (25).
-Next, participants were randomly assigned to one of three practice conditions: (1) rewriting a new cover letter with an AI writing tool that instantly revised text based on these principles, (2) rewriting the new cover letter without the AI tool, or (3) a no-practice control.
-At the end of the session, all participants completed a test of writing skill (rewriting a new cover letter without access to the AI writing tool) and an exit questionnaire.
-One day later, all participants were invited to complete a similar incentivized test of writing skill.
+In Study 2 (N = 2,238), we preregistered an experiment in which participants completed a baseline questionnaire and pretest, followed by a lesson on effective writing.
+They were then randomly assigned to one of three practice conditions: (1) rewriting a new cover letter with an AI writing tool that instantly revised text based on these principles, (2) rewriting the new cover letter without the AI tool, or (3) a no-practice control.
+Because forgetting occurs rapidly—with substantial decay even within 24 hours (e.g., (roediger_test-enhanced_2006, murre2015replication)), we invited all participants to complete a similar incentivized test of writing skill one day later.
 See Figure 3.
-We assessed writing skill in two ways.
-First, we used GPT-4o to rate cover letters on each of the principles of effective writing and then averaged these ratings to produce a summary score (Cronbach’s = .81).
-This operationalization demonstrated convergent validity with ratings by human research assistants who rated a random subsample of n = 100 cover letters on the same principles (r = .70, p < .001).
-Second, we recruited participants naive to this investigation to read two letters and select the one more likely to secure its writer a job interview.
-These two operationalizations converged: preferred cover letters obtained higher AI ratings of writing skill (rs = .29).
 AI practice improved writing skill
-Consistent with other studies demonstrating the productivity benefits of AI tools (3, 2), participants given access to the AI writing tool produced cover letters during the practice phase that were dramatically higher in quality than participants without access (d = 1.01, p <.001).
-The learning advantage of having practiced with AI was evident in the test phase: Consistent with our preregistered hypothesis, participants who had practiced with the AI tool produced higher-quality writing than did participants who either had practiced without the AI tool (d = .38, p <.001) or who had not practiced at all (d = .47, p <.001).
+Consistent with our preregistered hypothesis, participants who had practiced with the AI tool wrote better cover letters than participants who either had practiced without the AI tool (d = .38, p <.001) or who had not practiced at all (d = .47, p <.001).
 See Figure S3.
@@
 Compared to participants who practiced alone, participants who practiced with the AI tool spent 0.44 fewer minutes during the practice phase (3.73 vs.
-4.17; d = .12, p = .025, quantity of effort), logged roughly a quarter as many total keystrokes (26% d = .44, p <.001) and logged 65% fewer keystrokes per minute (d = .22, p <.001, intensity or quality of effort), and self-reported expending less effort during practice (d = .31, p <.001).
+4.17; d = .12, p = .025), logged roughly a quarter as many total keystrokes (26% d = .44, p <.001) and logged 65% fewer keystrokes per minute (d = .22, p <.001), and self-reported expending less effort during practice (d = .31, p <.001).
 Nevertheless, it would be inaccurate to label writers practicing with AI as entirely disengaged.
-Rather than copying and pasting the AI tool’s output, the majority of participants practicing with AI chose to interact with the task for over 3 minutes, and 95% made at least one edit to the AI-generated output before final submission.
+Rather than simply copying and pasting the AI tool’s output, the majority of participants practicing with AI chose to interact with the task for over 3 minutes, and 95% made at least one edit to the AI-generated output before final submission.
 See Figure S4 in Supplementary Information for details.
@@
 When asked if they wanted to see additional feedback, they were about as likely to say yes as participants who had practiced without AI (.64 vs.
-.60, p = .167), and slightly less likely than those who had not practice at all (.65 vs.
+.60, p = .167), and slightly less likely than those who had not practiced at all (.65 vs.
 .60, p = .039).
 See Section C4 in the Supplementary Information for details.
-The benefits of practicing with AI were just as large a day later
+The benefits of practicing with AI were just as large one day later
 To examine whether the treatment effects persisted over time, we re-contacted all participants one day later.
@@
 See Table S14 in the Supplementary Information for details.
-Study 3: Willingness to pay for AI vs. human editors
-Given unwarranted pessimism about the pedagogical power of AI, in Study 3 we examined whether people underinvest in AI as a learning resource.
-Specifically, we asked N = 150 people to imagine they were participants in Study 2 and asked them how much they would be willing to pay for feedback from the AI tool vs.
-feedback from experienced human editors.
-Participants were willing to pay more for feedback from expert human editors than for AI feedback
-Out of a maximum of a dollar (the total amount of the hypothetical bonus participants could earn), participants on average were willing to pay 28 cents for the AI tool vs.
-38 cents for feedback from an experienced human editor (t(149) = 4.83, p <.001, d = 0.39).
-Participants predicted that editorial feedback would be more beneficial
-Participants predicted that experienced human editors (M = 6.68 on a 0—10 Likert scale) would be more effective in helping them improve their writing, compared to feedback from an AI tool (M = 5.64; t(148) = 4.77, p <.001, d = 0.39).
-Predicted effectiveness mediated underinvestment in AI feedback
-People’s beliefs about how effective AI and human editor feedback would be predicted willingness to pay (b = 0.06, SE = 0.01, p < .001) and accounted for most of the preference for human editors (indirect effect = 0.06, p < .001).
-Once these predictions were accounted for, the remaining preference for human feedback was reduced by 60% (c’ = 0.04, SE = 0.02, p = .054).
-See Figure 5 and Section D1 of Supplementary Information for details.
-Study 4: Practicing with AI vs. with expert feedback (vs. with Google Search)
-Participants (N = 2,997) completed a baseline questionnaire and pretest, followed by the same lesson on effective writing as in Study 2, and then practiced rewriting a new cover letter on their own.
+Forecasters predicted AI would reduce effort and harm learning
+We showed N = 150 participants screenshots of this design and asked them to rank-order the study conditions according to how much they predicted participants would learn in each.
+Confirming our pre-registered hypothesis, nearly twice as many forecasters (64.7%) ranked practicing alone above practicing writing with access to an AI tool as the converse (35.3%, (1) = 12.9, p < .001), see Figure S6.
+Participants made this prediction regardless of self-reported experience with AI (OR = 0.72, p = .096) or any other measured demographic characteristic (ps > .05).
+See Table S16
+Forecasters who were pessimistic about the effect of the AI tool on learning speculated that it would crowd out effort (e.g., “Practicing alone would force more recall and problem-solving skills, while AI essentially just gives them the answer.”, “I think oftentimes using AI impedes the learning process because it’s the ‘easy way.”’).
+Those with positive views, on the other hand, cited the possibility of AI providing insights or examples that would be otherwise unavailable (“As much as I hate AI, I do not believe you can improve in any manner if you do not have examples or other ways of learning, and AI can provide this.”)
+Study 3: Practicing with AI vs. with expert feedback (vs. with Google Search)
+In Study 3 (N = 2,997), we preregistered an experiment testing whether AI was more helpful than two ecologically valid comparison conditions.
+Participants (N = 2,997) completed a baseline questionnaire and pretest, followed by the lesson on effective writing, and then practiced rewriting a new cover letter on their own.
 Next, participants were randomly assigned to one of three conditions: (1) practice with AI, (2) practice with feedback from professional human editors, (3) practice with Google Search.
@@
 During the practice phase, participants working with AI outperformed participants who received personalized feedback from professional editors (d = 0.76, p <.001) and those assigned to Google cover letter examples and tips (d = 1.03, p <.001).
-During the test phase, in which participants were asked to write a new cover letter without access to any outside resources, participants who had practiced with the AI tool produced higher-quality cover letters than those who had practiced with Google Search (d = 0.46, p <.001), and even outperformed those who had practiced with personalized feedback from professional editors (d = 0.20, p <.001).
+During the test phase, in which participants were asked to write a new cover letter without access to any outside resources, participants who had practiced with the AI tool wrote higher-quality cover letters than those who had practiced with Google Search (d = 0.46, p <.001), and even outperformed those who had practiced with personalized feedback from professional editors (d = 0.20, p <.001).
 See Figure 7A.
@@
 Thus, the AI tool’s advantage over professional editors cannot be attributed to a decline in editorial performance.
-See Section E2 of Supplementary Information for details.
+See Section D2 of Supplementary Information for details.
+The quality of edits provided by AI was higher than that of edits from professional human editors.
+We used the AI tool to rewrite each of the texts assigned to the editor pool, creating a counterfactual of how AI would have edited those letters.
+Relative to the counterfactual, letters rewritten with the AI tool were rated higher for writing quality (d = 1.66, p < .001), and when shown to independent raters, they were considerably more likely to secure a hypothetical job interview (d = 0.75, p < .001).
+Thus, AI users encountered examples that more closely adhered to the writing principles than those provided by human editors, helping explain the learning advantage associated with AI use.
 Practicing with AI was not more effortful than getting feedback from professional editors and was less effortful than looking for examples online
@@
 Nevertheless, participants across all three groups reported similar levels of subjective effort (ps > .05).
-Likewise, participants who practiced with AI logged about as many keystroles as those who practiced with editor feedback (68% fewer, d = 0.08, p > .05) and 76% fewer keystrokes than those who practiced with Google Search (d = 0.32, p < .001).
+Likewise, participants who practiced with AI logged about as many keystrokes as those who practiced with editor feedback (68% fewer, d = 0.08, p > .05) and 76% fewer keystrokes than those who practiced with Google Search (d = 0.32, p < .001).
 They also typed at a slower rate than participants who practiced with editor feedback (d = 0.41, p < .001) or those who practiced with Google Search (d = 0.34, p < .001).
@@
 Both groups were more likely to request feedback than those who had practiced with Google search (64%, p = .003 and .007, respectively).
-See Section E6 in the Supplementary Information for details.
+See Section D6 in the Supplementary Information for details.
 Results were not moderated by individual differences
 As in Study 2, the above findings were consistent across all measured individual differences. See Table S24 in the Supplementary Information for details.
-Study 5: Examples as the mechanism for AI learning gains
-To better understand what drives the benefit of AI practice, in Study 5 (N = 2,003), we preregistered a replication and extension in which we replaced the no-practice condition of Study 2 with an example-only condition.
+Forecasters predicted AI would be less helpful than human editors and were less willing to pay for it
+We asked N = 150 people to imagine they were participants in our experiment and asked them how much they would be willing to pay for feedback from the AI tool vs.
+feedback from experienced human editors.
+Out of a maximum of a dollar (the total amount of the hypothetical bonus participants could earn), participants on average were willing to pay 28 cents for the AI tool vs.
+38 cents for feedback from an experienced human editor (t(149) = 4.83, p < .001, d = 0.39).
+Participants predicted that experienced human editors (M = 6.68 on a 0—10 Likert scale) would be more effective in helping them improve their writing, compared to feedback from an AI tool (M = 5.64; t(148) = 4.77, p < .001, d = 0.39).
+People’s beliefs about the effectiveness of AI and human editor feedback predicted willingness to pay (b = 0.06, SE = 0.01, p < .001) and accounted for most of the preference for human editors (indirect effect = 0.06, p < .001).
+Once these predictions were accounted for, the remaining preference for human feedback was reduced by 60% (c’ = 0.04, SE = 0.02, p = .054).
+See Figure S11 and Section D9 of Supplementary Information for details.
+Study 4: Examples as the mechanism for AI learning gains
+To better understand how using AI improves writing skill, in Study 4 (N = 2,003), we preregistered a replication and extension in which we replaced the no-practice condition of Study 2 with an example-only condition.
 In this condition, we showed participants an AI-generated writing example that they could not edit.
@@
 .52, p = .561).
-See Figures S10 and 7B.
+See Figures S12 and 8B.
 Seeing an AI example was even less effortful than practicing with AI
@@
 Moreover, all participants requested feedback at similar rates (proportions ranged from 63% to 67%).
-See Section F4 in the Supplementary Information for details.
+See Section E4 in the Supplementary Information for details.
 The benefits of seeing an AI example were just as large a day later
@@
 The effect remained robust after 24 hours.
-Participants who had practiced with the AI-tool (d = .29, p = .006) and participants who had simply seen an AI example (d = .32, p = .003), both continued to outperform those who had practiced without the tool.
+Participants who had practiced with the AI tool (d = .29, p = .006) and participants who had simply seen an AI example (d = .32, p = .003), both continued to outperform those who had practiced without the tool.
 Participants who had merely seen an AI example performed as well as those who had practiced with AI (d = .02, p = .830).
 As in Study 2, the magnitude of the follow-up effect size was not significantly different from the immediate test phase effect size (condition time interactions were non-significant (ps > .96).
-See Figure 7A and Section F5 of Supplementary Information for details.
+See Figure 8A and Section E5 of Supplementary Information for details.
 Results were not moderated by individual differences
@@
 Discussion
-Contrary to lay forecasters’ expectations (Studies 1 and 3), participants who practiced writing cover letters with an AI tool learned more than those who practiced on their own (Studies 2 and 5), more than those who had access to Google cover letter examples, and even more than those who received personalized feedback from professional human editors (Study 4).
-Specifically, participants who had practiced with AI, subsequently wrote cover letters (without AI assistance) that were rated higher in writing quality and were more likely to secure a hypothetical interview—both immediately after practice and one day later.
-Learning gains were not the result of greater effort; in fact, participants who had practiced with AI expended less effort during practice than those who had practiced alone.
-Instead, these gains were at least in part explained by exposure to high-quality, just-in-time personalized examples: participants who merely viewed an AI-generated cover letter (without editing it) improved their writing skill as much as those given the option to practice editing the cover letter (Study 5).
-Study 1 forecasters who believed AI would harm learning cited reduced effort as the mechanism—and prior research supports this concern (28, 29).
-All things equal, the more effort a learner invests in a given practice task, the more they learn.
-So why did practicing with AI improve writing skill more than practicing alone, even while reducing both the quantity and quality of effort?
-Likewise, why did writers who googled how to compose a cover letter work longer and harder than writers who practiced with AI, yet learn less?
-We propose a third mechanism that offsets decrements in human effort: the utility of the practice task itself.
-See Figure 8.
-Perhaps the excellent personalized examples provided by AI are analogous to demonstrations of an Olympic coach.
-The higher the utility of the practice task, the faster the learning.
-While not immediately apparent to forecasters, teachers (11), and the general population (8), the pedagogical benefits of viewing AI-generated examples are consistent with prior research on learning.
-Research has shown that humans are especially adept at observing, imitating, and learning from others (30, 31, 32).
-Our findings also align with the expert performance literature: the most successful learners engage in deliberate practice, which (in addition to concentration, feedback, and repetition) depends upon detailed mental representations of excellent performance (33).
-Future Directions
+Although many believe that using AI undermines effort and learning, we found in our experiments that participants assigned to practice writing cover letters with AI learned more than those assigned to practice on their own.
+Access to AI also led to more learning than did access to Google search for cover letter tips and examples, or even personalized feedback from professional human editors.
+Lay beliefs about AI and effort were more accurate.
+Participants given access to an AI tool expended less effort while practicing than those who practiced alone or who practiced with access to Google tips and examples.
+Not surprisingly, practicing with AI was about as effortful as practicing with feedback from a professional human editor.
+What accounts for the learning gains from practicing with AI?
+We argue that AI improves the learning environment by providing high-quality, just-in-time, personalized examples that illustrate otherwise abstract principles.
+As shown in Figure 1, even if AI reduces the effort people expend during practice, it may nevertheless improve learning by enhancing the learning environment.
+This mechanism is consistent with a large literature showing that humans are especially adept at learning by observation (bandura1971, meltzoff2005imitation, lyons2007hidden).
+Relatedly, individualized tutoring is a much more effective learning environment than whole group instruction (bloom_2_1984)—and experts in diverse domains engage in deliberate practice, acquiring increasingly more sophisticated mental representations provided by their teachers and coaches (anders2008deliberate).
+Compared to textbooks, conventional computer-tutoring programs (e.g., (anderson1995cognitive)), and even human instructors, AI systems are especially adept at instantly generating specific, high-quality examples tailored to a learner’s momentary needs.
+Indeed, we found that AI-generated examples of cover letters during the practice phase were better written and more likely to elicit interviews than cover letters produced by professional writers and editors.
 Three promising directions for future research are worth highlighting.
-First, it is not clear whether the observed benefits of AI examples would be as large in domains other than professional writing.
-At a glance, a single AI-generated example visually communicates the elements of effective professional writing (e.g., short sentences and boldface formatting).
-In other domains, however, merely observing a solution may be less informative.
-For instance, the final answer to a math or computer programming problem does not instantly reveal the procedure that produced it.
-Second, additional research is needed to explore moderators of learning from AI tools.
-Certain metacognitive strategies may enhance the benefits of interacting with AI.
-For example, experimental research suggests that learners benefit more from AI explanations for math problems if they first try to solve them on their own (21).
-Similarly, correlational evidence suggests that asking AI for explanations as opposed to answers is associated with more learning in mathematics (23) and computer programming (22).
-Conversely, other factors may minimize learning benefits.
-In our experimental paradigm, participants practiced for as long as they wanted, with the foreknowledge that their skills would subsequently be tested (and rewarded monetarily) without access to AI.
-During practice, therefore, participants were incentivized to prioritize gains in acquired skill over performance in the moment.
-In real-world settings, there is often time pressure and competing incentives around performance and learning, which we speculate would reduce the learning gains associated with practicing with AI.
-More broadly, we believe concerns about AI as a crutch are justified when AI displaces effort without increasing task utility.
-Third, in our experimental paradigm, participants interacted with the AI tool only once.
-It is common, however, to use AI tools repeatedly.
+First, it is not clear whether the observed benefits of AI examples would generalize to other domains like math (kumar2023, bastani2024) and computer programming (nie2024gpt).
+At a glance, a single AI-generated example illustrates the principles of effective professional writing.
+In other domains, merely observing an example may be less informative.
+For instance, the final answer to a math or computer programming problem does not often reveal the process that produced it.
+Second, additional research is needed to explore moderators.
+There may be metacognitive strategies that enhance the benefits of writing with AI.
+For instance, trying on one’s own before using AI has been shown to improve learning more than the converse (cf.
+(kumar2023)).
+Conversely, time pressure and competing priorities around performance and learning may minimize the learning benefits.
+More broadly, we believe concerns about AI as a crutch are justified when AI displaces effort without improving the learning experience.
+Third, in our experimental paradigm, participants interacted with AI only once.
+It is common, however, to use AI repeatedly.
 When do repeated interactions lead to diminishing or even negative returns, and in what scenarios might skill development continue over time?
 Consider, for instance, the game of Go.
-The introduction of superhuman AI has been associated with an increase in the novelty and quality of decisions made by human Go players over time, with elite players reporting that they have been inspired by AI solutions they’d never seen before (34).
-Future research, ideally in field settings, is needed to establish the long-term benefits and costs of using AI tools.
-Conclusion
-Our findings should temper widespread concern that AI tools invariably boost momentary productivity at the expense of long-term skill development.
-Although the AI writing tool reduced the effort users invested in practicing, it nevertheless accelerated their skill development.
-The underappreciated efficacy of timely and tailored examples carries practical implications: Many AI tools designed to support learning are explicitly programmed not to “give away” answers.
-However, it may be that in addition to hints, leading questions, and explanations, learners benefit from demonstrations of the principles they are attempting to master.
-Decades before the advent of generative AI, the legendary UCLA basketball coach John Wooden declared that the four laws of learning are explanation, demonstration, imitation, and repetition (35).
-Few learners have access to the best human teachers, coaches, and mentors, but generative AI now makes it possible to learn from personalized, just-in-time demonstrations tailored to any domain.
-In doing so, AI has the potential not only to boost productivity but also to democratize opportunities for building human capital at scale.
+The introduction of superhuman AI has been associated with an increase in the novelty and quality of decisions made by human Go players over time, with elite players reporting that they have been inspired by AI solutions they’d never seen before (shin2023).
+Future research, ideally in field settings, is needed to establish the long-term benefits and costs of using AI.
+Taken together, our findings challenge the widely held view that productivity boosts from AI inevitably hamper the development of human skill.
+By illustrating otherwise abstract principles with examples personalized to the precise needs of the learner at the moment, AI can, at least in some use cases, improve the quality of the learning environment.
+Rather than viewing AI as a universal threat to human capital, our results provide an existence proof that AI can support the development of skills at scale.
 Methods
@@
 All studies were preregistered on https://aspredicted.org. Accordingly, the analyses presented in the main text were also pre-registered, with exceptions noted below.
-Study 1 was preregistered (#191800). The moderation analyses were not pre-registered.
+Study 1 was a descriptive survey and was not preregistered.
 Study 2 was pre-registered (#201140).
@@
 After preregistration, we considered a keystrokes-per-minute metric of effort intensity rather than duration; thus, this analysis was not preregistered.
-Study 3 was pre-registered (#236487). Analyses on predicted effectiveness of AI vs. human feedback and the mediation analysis were not preregistered.
-Study 4 was pre-registered (#239157). The pairwise comparisons analysis was pre-registered separately (#243115). The keystrokes-per-minute metric of effort was not pre-registered.
-Study 5 was pre-registered (#197704).
+The forecasting data was preregistered (#191800), but not the moderation analysis within it.
+Study 3 was pre-registered (#239157).
+The pairwise comparisons analysis (#243115) and the comparison between editor rewrites and counterfactual AI rewrites (#255792) were pre-registered separately.
+The keystrokes-per-minute metric of effort was not pre-registered.
+The willingness to pay data was preregistered (#236487), but not the analyses on the predicted effectiveness of AI vs.
+human feedback and the mediation analysis.
+Study 4 was pre-registered (#197704).
 The one-day follow-up was collected in three batches: a pilot, followed by second and third batches.
-We pre-registered batch 2 (#199451) but report pooled results in the main text, with separate batch details available in Supplementary Information Section F5.
+We pre-registered batch 2 (#199451), but report pooled results in the main text, with separate batch details available in Supplementary Information Section E5.
 The pairwise comparisons analysis was pre-registered separately (#205315).
 The keystrokes-per-minute metric of effort was not pre-registered.
+See all preregistrations here.
 Participants
-We sampled participants from Prolific from all our studies. We excluded all Prolific users who participated in one of the earlier studies from participation in subsequent studies.
-Participants in Study 1 were predominantly female (n = 93, 62%), with ages ranging from 21 to 81 (M = 38.4, SD = 12.2).
-They were predominantly white (75%).
-A small proportion were students (13%), and most were employed (62%).
-In Study 2, the sample was more evenly split between men (46%) and women (52%), ranging in age from 18 to 82 (M = 36.0, SD = 12.5).
+Participants in Study 1 were a nationally representative sample of young adults aged 18-28 (M = 25.3, SD = 2.4) recruited through the Gallup organization.
+The sample comprised 38% men, 57% women, and 5% nonbinary participants.
+The majority (64%) were White, with the remainder comprising Hispanic (15%), Asian (10%), Black (10%), and other race (1%) participants.
+We used sample weights to ensure national representativeness of the data for all analyses.
+We sampled participants from Prolific for all our experiments. We excluded all Prolific users who participated in one of the earlier studies from participation in subsequent studies.
+In Study 2, the sample was somewhat evenly split between men (46%) and women (52%), ranging in age from 18 to 82 (M = 36.0, SD = 12.5).
 Over half of the sample (58%) was White, with the rest comprising Black (33%), Latino (6%), and Asian (5%) participants.
@@
 Most participants (93%) were at least somewhat motivated to improve their writing, and had varying levels of experience with AI writing assistants (36% had tried them, but hardly ever used them, 47% used them at least a few times per week, and 17% had never used AI assistants before).
-Participants in Study 3 were evenly split between men (52%) and women (48%), and their ages ranged from 18 to 75 (M = 42.4, SD = 14.4).
-They were predominantly white (67%).
-A small proportion were students (19%), and most were employed (56%).
-In Study 4, the sample consisted of relatively fewer men (38%) than women (61%), and ranged in age from 19 to 92 (M = 41.5, SD = 13.8).
+Participants forecasting the results of Study 2 were predominantly female (n = 93, 62%), with ages ranging from 21 to 81 (M = 38.4, SD = 12.2).
+They were predominantly white (75%).
+A small proportion were students (13%), and most were employed (62%).
+In Study 3, the sample consisted of relatively fewer men (38%) than women (61%), and ranged in age from 19 to 92 (M = 41.5, SD = 13.8).
 Over half of the sample (74%) was White, with the rest comprising Black (16%), Latino (8%), and Asian (7%) participants.
@@
 Most participants (94%) were at least somewhat motivated to improve their writing, and had varying levels of experience with AI writing assistants: 17% had never used AI assistants, 40% had tried them but hardly ever used them, and 43% used them at least weekly.
-Study 5 had similar proportions of men (46%) and women (53%), and participants ranged in age from 18 to 95 (M = 37.9, SD = 12.6).
+Participants forecasting Study 3 were evenly split between men (52%) and women (48%), and their ages ranged from 18 to 75 (M = 42.4, SD = 14.4).
+They were predominantly white (67%).
+A small proportion were students (19%), and most were employed (56%).
+Study 4 had similar proportions of men (46%) and women (53%), and participants ranged in age from 18 to 95 (M = 37.9, SD = 12.6).
 Over half of the sample (64%) was White, with the remainder comprising Black (24%), Latino (8%), and Asian (6%) participants.
@@
 Procedure
-All studies followed the same general structure with study-specific variations.
+In Study 1, participants completed a brief survey about their AI usage in the past month and their general attitudes toward AI and learning.
+Participants rated their agreement with statements about AI’s impact on learning, intelligence, and work habits using 6-point Likert-type scales (1 = Strongly disagree, 6 = Strongly agree).
+All experiments (Studies 2–4) followed the same general structure with study-specific variations.
 Participants first saw an introductory screen and completed a brief questionnaire reporting demographics, prior experience with AI writing tools, motivation to improve their writing, and perceived writing skill.
 They then completed a short pre-test (2 minutes) in which they edited a poorly written cover letter, followed by a brief lesson on the five principles of effective writing.
-In Studies 2 and 5, participants were then randomized to a practice condition (or skipped practice in the no-practice control).
+In Studies 2 and 4, participants were then randomized to a practice condition (or skipped practice in the no-practice control).
 During practice, participants rewrote a new cover letter or, in the example-only condition, viewed an AI-rewritten version of the letter.
@@
 Finally, participants could request optional feedback and were asked whether they had used outside resources during the test.
-A small percentage of participants (2.9%) admitted to doing so; as per our preregistration, these participants are included in analyses, though results excluding them are consistent with our main interpretation (see Tables S6, S18, and S27).
-In Study 4, we adapted the procedure to compare AI practice with two ecologically valid human alternatives: feedback from professional editors and guidance from Google Search.
+A small percentage of participants (2.9%) admitted to doing so; as per our preregistration, these participants are included in analyses, though results excluding them are consistent with our main interpretation (see Tables S7, S20, and S29).
+In Study 3, we adapted the procedure to compare AI practice with two ecologically valid human alternatives: feedback from professional editors and guidance from Google Search.
 Participants completed the same baseline questionnaire, pretest, and writing lesson.
 They then edited a new cover letter without any assistance and were randomly assigned to one of three conditions: (1) practice with AI, (2) practice with feedback from professional editors, (3) practice with a Google Search for cover letter examples and tips.
-In the second condition condition, participants’ cover letters were forwarded to editors recruited in partnership with The Journalists’ Resource at Harvard Kennedy School, who provided both written feedback and a rewritten version of the letter.
+In the second condition, participants’ cover letters were forwarded to editors recruited in partnership with The Journalists’ Resource at Harvard Kennedy School, who provided both written feedback and a rewritten version of the letter.
 The following day, participants were asked to revise their original cover letter using the resources from their assigned condition.
@@
 Measurement
-As per our pre-registration, we used OpenAI’s GPT-4o to rate text samples for writing quality.
+For Studies 2-4, as per our pre-registration, we used OpenAI’s GPT-4o to rate text samples for writing quality.
 To do this, we independently rated each cover letter and each writing principle.
-Research has demonstrated that large language models can provide ratings of writing quality that align closely with human judgments, offering reliability and consistency across various evaluation contexts (36, 37).
+Research has demonstrated that large language models can provide ratings of writing quality that align closely with human judgments, offering reliability and consistency across various evaluation contexts (rathje2024, hackl2023).
 See our prompts in Table S1.
 We then took the unweighted average of these 5 scores as our main dependent variable.
-See disaggregated analyses by each writing principle and additional outcomes on Tables S6, S18, and S27.
+See disaggregated analyses by each writing principle and additional outcomes on Tables S7, S20, and S29.
 To validate these ratings, the first author and a trained research assistant took a sample of n = 100 cover letters from Study 2, and rated them on the 5 principles.
-The average of these two ratings correlated about as highly with the computer ratings (r = .70, p < .001) as they did with the average interrater reliability (r = .74, p < .001).
-These ratings had high internal consistency (Cronbach’s s = 0.81, 0.78, and 0.83 for studies 2, 4, 5, respectively.)
+The average of these two ratings correlated about as highly with the computer ratings (r = .70, p < .001) as they did with the average interrater reliability (r = .74, p < .001, difference in rs, p = .56,).
+These ratings had high internal consistency (Cronbach’s s = 0.81, 0.78, and 0.83 for studies 2, 3, and 4, respectively).
 To address concerns that particular LLMs might be biased in favor of their own output, we also used Claude to rate the cover letters.
-We find that GPT ratings correlate strongly with Claude ratings (r = .71, p < .001), and that the effects are not attenuated by using different models (See Tables S6, S18, and S27), suggesting that our effects are not explainable by same-model bias.
+We find that GPT ratings correlate strongly with Claude ratings (r = .71, p < .001), and that the effects are not attenuated by using different models (See Tables S7, S20, and S29), suggesting that our effects are not explainable by same-model bias.
 Additionally, we asked a separate sample of participants to read pairs of test-phase cover letters randomly selected from different conditions, and to indicate which letter would be more likely to secure a job interview.
-Cover letters more likely to secure an interview obtained higher AI ratings of writing skill (rs = .29, .35, and .28, ps < .001, for Study 2, 4 and 5, respectively).
+Cover letters more likely to secure an interview obtained higher AI ratings of writing skill (rs = .29, .35, and .28, ps < .001, for Study 2, 3, and 4, respectively).
 Statistical analysis
-As per our pre-registrations, we fit ANCOVA models predicting outcomes from condition indicators, controlling for pretest score and baseline characteristics (age, gender, race/ethnicity, primary language, education level, motivation to improve writing skills, self-rated writing skill, experience with AI writing assistants, and baseline writing effectiveness).
+For Study 1, we conducted one-sample t-tests comparing mean responses to the scale midpoint.
+For Studies 2–4, as per our pre-registrations, we fit ANCOVA models predicting outcomes from condition indicators, controlling for pretest score and baseline characteristics (age, gender, race/ethnicity, primary language, education level, motivation to improve writing skills, self-rated writing skill, experience with AI writing assistants, and baseline writing effectiveness).
 We used logistic regression to predict whether participants chose to see optional test feedback from condition, controlling for pretest score and baseline characteristics.
 Our analyses of the hypothetical hiring situation use beta regression, because the relative likelihood of a cover letter being preferred is bounded between 0 and 1.
-When correcting for multiple comparisons in exploratory moderation analyses, we used the Benjamini-Hochberg correction (38).
+When correcting for multiple comparisons in exploratory moderation analyses, we used the Benjamini-Hochberg correction (benjamini1995controlling).
 C Results Study 2
@@
 - C7 Pairwise comparisons
+- C8 Forecasting
 D Results Study 3
+- D1 Randomization, Balance, and Missingness
+- D2 Details on professional editors
+- D3 Details on participants’ search behavior
+- D4 AI practice improved writing skill more than getting feedback from professional editors and looking for examples online
+- D5 AI practice wasn’t any more effortful than getting feedback from professional editors and was less effortful than looking for examples online
+- D6 AI practice did not create more of an illusion of mastery than getting feedback from professional editors or looking for examples online
+- D7 Results were not moderated by individual differences
+- D8 Pairwise comparisons
+- D9 Forecasting and Willingness to Pay
 E Results Study 4
 - E1 Randomization, Balance, and Missingness
-- E2 Details on professional editors
-- E3 Details on participants’ search behavior
-- E4 AI practice improved writing skill more than getting feedback from professional editors and looking for examples online
-- E5 AI practice wasn’t any more effortful than getting feedback from professional editors and was less effortful than looking for examples online
-- E6 AI practice did not create more of an illusion of mastery than getting feedback from professional editors or looking for examples online
-- E7 Results were not moderated by individual differences
-- E8 Pairwise comparisons
-F Results Study 5
-- F1 Randomization, Balance, and Missingness
-- F2 AI examples improve writing skill
-- F3 Seeing AI examples was less effortful
-- F4 Seeing an AI example did not discourage motivation for future learning
-- F5 The benefits of seeing an AI example were just as large a day later
-- F6 Seeing AI examples was equally effective across subgroups
-- F7 Pairwise Comparisons
+- E2 AI examples improve writing skill
+- E3 Seeing AI examples was less effortful
+- E4 Seeing an AI example did not discourage motivation for future learning
+- E5 The benefits of seeing an AI example were just as large a day later
+- E6 Seeing AI examples was equally effective across subgroups
+- E7 Pairwise Comparisons
 Appendix A Additional methods
@@
 The human raters correlated with the AI-generated ratings satisfactorily (r = .70, p < .001).
+The correlation indicating AI-human agreement was not smaller than the interrater agreement correlation (difference in rs = .04, p = .56), suggesting that AI ratings achieved similar levels of consistency as human raters.
 See Figure S1.
@@
 Appendix B Results Study 1
-Table S2 reports results from a logistic regression predicting the probability of answering correctly as a function of demographic and background characteristics.
-Odds ratios (ORs) are displayed along with their standard errors, Wald z statistics, and p-values.
-The model includes age, sex, ethnicity, student status, employment, and prior AI experience as predictors.
-| Predictor | OR | SE | z | p-value |
-| Prior AI experience | 0.718 | 0.199 | -1.664 | 0.096 |
-| Age (years) | 1.014 | 0.021 | 0.675 | 0.500 |
-| Male (vs. Female) | 1.035 | 0.467 | 0.073 | 0.942 |
-| White (vs. Other) | 0.966 | 0.509 | -0.068 | 0.946 |
-| Student (vs. Not) | 1.120 | 0.618 | 0.183 | 0.855 |
-| Employment status | ||||
-| Other | 0.770 | 0.687 | -0.382 | 0.703 |
-| Part-time | 0.663 | 0.601 | -0.684 | 0.494 |
-| Unemployed (seeking) | 0.645 | 0.593 | -0.738 | 0.461 |
+We partnered with the Gallup organization to conduct a nationally representative survey of young adults (18-28 years old, N = 2,472).
+All analyses use survey weights to ensure representativeness of the U.S.
+population.
+B1 Usage
+We queried participants about the frequency with which they used AI for different purposes. Table S2 shows the percentage of participants who engaged in these uses at least once in the past month.
+| Use Case | Percentage |
+| Google replacement | 64.7 |
+| Work | 51.5 |
+| Writing | 45.9 |
+| Image generation | 39.3 |
+| Life advice | 32.2 |
+| Math | 29.8 |
+| Friend/companionship | 22.9 |
+| Cheating | 16.4 |
+| Romantic advice | 9.9 |
+B2 Attitudes
+Participants answered four items assessing their beliefs about AI and its effects on human capabilities: “AI makes people less smart”, “AI reduces people’s ability to learn”, “AI makes people more smart” (reverse-coded), and “AI makes people lazier”.
+The four items showed good internal consistency (Cronbach’s = 0.84; weighted = 0.80).
+All four items showed significantly more negative attitudes than the scale midpoint (all ps < .001).
+Specifically, 63% agreed that AI makes people less smart, 62% agreed that AI reduces the ability to learn, and 79% agreed that AI makes people lazier.
+Overall, 73% of respondents scored above the midpoint on a composite variable composed of these 4 items, indicating predominantly negative attitudes.
+See Table S3
+| Item | M | SE | df | Cohen’s | ||
+| “AI makes people less smart” | 3.93 | 0.04 | 10.3 | 2636 | 0.30 | |
+| “AI reduces ability to learn” | 3.99 | 0.04 | 11.8 | 2636 | 0.34 | |
+| “AI makes people more smart” (rev.) | 2.80 | 0.04 | -18.9 | 2636 | -0.55 | |
+| “AI makes people lazier” | 4.37 | 0.04 | 22.6 | 2636 | 0.66 | |
+| AI Impact Score (composite) | 4.12 | 0.03 | 20.4 | 2636 | 0.57 |
+There was no significant relationship between age and the composite measure of AI Impact Score ( = 0.005, p = 0.507).
 Appendix C Results Study 2
@@
 C6 AI practice was equally effective across subgroups
-To test for moderation effects of pre-treatment demographic variables, we ran separate linear in which writing skill during the test phase was regressed on condition, the pre-treatment moderator of interest, writing skill at baseline, and an interaction term between the moderator condition.
+To test for moderation effects of pre-treatment demographic variables, we ran separate linear regressions in which writing skill during the test phase was regressed on condition, the pre-treatment moderator of interest, writing skill at baseline, and an interaction term between the moderator condition.
 After correcting the p-values for the interaction terms, none were significant at the .05 level, suggesting that practicing with AI was equally effective across groups.
@@
 | log(Likelihood) | -1218.54 | -1134.64 |
+C8 Forecasting
+Table S16 reports results from a logistic regression predicting the probability of answering correctly as a function of demographic and background characteristics.
+Odds ratios (ORs) are displayed along with their standard errors, Wald z statistics, and p-values.
+The model includes age, sex, ethnicity, student status, employment, and prior AI experience as predictors.
+| Predictor | OR | SE | z | p-value |
+| Prior AI experience | 0.718 | 0.199 | -1.664 | 0.096 |
+| Age (years) | 1.014 | 0.021 | 0.675 | 0.500 |
+| Male (vs. Female) | 1.035 | 0.467 | 0.073 | 0.942 |
+| White (vs. Other) | 0.966 | 0.509 | -0.068 | 0.946 |
+| Student (vs. Not) | 1.120 | 0.618 | 0.183 | 0.855 |
+| Employment status | ||||
+| Other | 0.770 | 0.687 | -0.382 | 0.703 |
+| Part-time | 0.663 | 0.601 | -0.684 | 0.494 |
+| Unemployed (seeking) | 0.645 | 0.593 | -0.738 | 0.461 |
 Appendix D Results Study 3
-D1 Mediation
-Participants predicted they would learn more from the feedback of human editors than from AI (a path = 1.05, SE = 0.22, p <.001).
-In turn, higher predicted effectiveness predicted greater willingness to pay (b path = 0.06, SE = 0.01, p <.001).
-Participants’ greater willingness to pay for human feedback was largely explained by their prediction that they would learn more from experienced human editors than from AI (indirect effect = 0.06, SE = 0.01, p < .001), accounting for most of the overall tendency to pay more for human editors (total effect = 0.10, SE = 0.02, p < .001).
-Once these predictions were taken into account, the remaining preference for human feedback was small and only marginally significant (c’ path = 0.04, SE = 0.02, p = .054).
-See Figure S6 and Section D1 of Supplementary Information for details.
-Appendix E Results Study 4
-E1 Randomization, Balance, and Missingness
-Participants in Study 4 were recontacted to return two days after the initial intake section.
+D1 Randomization, Balance, and Missingness
+Participants in Study 3 were recontacted to return two days after the initial intake section.
 While most people responded, 14% of recontacted participants did not respond.
@@
 | Pretest Writing Skill (mean (SD)) | 4.40 (0.67) | 4.41 (0.67) | 4.40 (0.68) | 4.40 (0.65) | .964 | 0.008 |
-E2 Details on professional editors
+D2 Details on professional editors
 We recruited 44 professional writers and editors through Journalist’s Resource, paying them $100 to edit 20 cover letters each.
@@
 In total, they edited 1,227 cover letters.
-To determine if editors changed their behavior over time, we fit a series of multilevel models, with cover letter position as a predictor, random intercept for editors, and text quality (as rated by GPT-4o), time spent per cover letter, and number of insertions, deletions, or modifications (i.e., Levenstein distance).
+To determine if editors changed their behavior over time, we fit a series of multilevel models, with cover letter position as a predictor, random intercept for editors, and text quality (as rated by GPT-4o), time spent per cover letter, and number of insertions, deletions, or modifications (i.e., Levenshtein distance).
 These analyses include editors who were able to edit at least 20 letters, and focus on the first 20 letters edited.
@@
 Despite spending less time per letter, editors produced higher quality edits with fewer but more effective changes, demonstrating clear learning effects in professional editing tasks.
-E3 Details on participants’ search behavior
+Why did participants learn more from AI than they did from professional human editors.
+One possibility is that the quality of the edits provided by AI were of higher quality than those of human editors.
+To test this possibility, we used our AI tool to rewrite the same letters that were assigned to professional human editors.
+That way, for each participant in the “practice with editor feedback condition” we have the actual letter they received (as edited by the human expert), and the counterfactual letter they would have received had they been assigned to practice with AI.
+As shown in Figure 8(a), a sample of naive human evaluators were twice as likely to prefer hte cover letter edited by AI than the one participants received, edited by a professional human editor (paired d = 0.62, p < .001) Figure 8(b) shows the densities of writing quality as evaluated by our AI rating procedure.
+Cover letters edited by AI were rated as significantly higher quality than those submitted by professional human editors (paired d = 1.66, p < .001)
+D3 Details on participants’ search behavior
 We replicated the first five pages of a Google search for “cover letter examples” and made this available to participants through Qualtrics to monitor their search behavior.
-Across 635 participants assigned to use Google to search for cover letter examples and tips, we found search behavior consistent with prior research on web-browsing behavior (39).
+Across 635 participants assigned to use Google to search for cover letter examples and tips, we observed search behavior consistent with prior research on web browsing (pan2007google).
 Most participants (97%) stayed on page 1, while only 3% ventured beyond the first page of results.
@@
 On average, participants made 2.9 clicks during their search session.
-E4 AI practice improved writing skill more than getting feedback from professional editors and looking for examples online
+D4 AI practice improved writing skill more than getting feedback from professional editors and looking for examples online
 The AI tool improved performance while participants used it.
@@
 | |||||||||
-E5 AI practice wasn’t any more effortful than getting feedback from professional editors and was less effortful than looking for examples online
+D5 AI practice wasn’t any more effortful than getting feedback from professional editors and was less effortful than looking for examples online
 Table S21 shows OLS models predicting practice effort metrics from practice condition.
@@
 | Note. *** p < .001, ** p < .01, * p < .05. | ||||||
-E6 AI practice did not create more of an illusion of mastery than getting feedback from professional editors or looking for examples online
+D6 AI practice did not create more of an illusion of mastery than getting feedback from professional editors or looking for examples online
 As reported in the main text, there were no differences between participants who practiced with AI and participants who practiced with editor feedback on how much they thought they learned, their perceived skill, and their likelihood to ask for feedback after the test.
@@
 | Note. *** p < .001, ** p < .01, * p < .05. | |||
-E7 Results were not moderated by individual differences
+D7 Results were not moderated by individual differences
 As in Study 2, we tested whether each of the pretreatment demographic variables moderated the effects of seeing an AI example.
-To do this, we ran separate linear in which writing skill during the test phase was regressed on condition, the pre-treatment moderator of interest, writing skill at baseline, and an interaction term between the moderator condition.
+To do this, we ran separate linear regressions in which writing skill during the test phase was regressed on condition, the pre-treatment moderator of interest, writing skill at baseline, and an interaction term between the moderator condition.
 After correcting the p-values for the interaction terms, none were significant at the .05 level, suggesting that seeing AI examples was equally effective across groups.
@@
 = Perceived | ||||||||||||||
-E8 Pairwise comparisons
+D8 Pairwise comparisons
 The relative likelihood of a cover letter receiving an invitation to an interview was correlated with the GPT-rated writing quality. See Figure S10.
@@
 | log(Likelihood) | -1843.89 | -1861.32 | -428.79 | -428.79 |
-Appendix F Results Study 5
-F1 Randomization, Balance, and Missingness
+D9 Forecasting and Willingness to Pay
+Participants predicted they would learn more from the feedback of human editors than from AI (a path = 1.05, SE = 0.22, p <.001).
+In turn, higher predicted effectiveness predicted greater willingness to pay (b path = 0.06, SE = 0.01, p <.001).
+Participants’ greater willingness to pay for human feedback was largely explained by their prediction that they would learn more from experienced human editors than from AI (indirect effect = 0.06, SE = 0.01, p < .001), accounting for most of the overall tendency to pay more for human editors (total effect = 0.10, SE = 0.02, p < .001).
+Once these predictions were taken into account, the remaining preference for human feedback was small and only marginally significant (c’ path = 0.04, SE = 0.02, p = .054).
+See Figure S11 and Section D9 of Supplementary Information for details.
+Appendix E Results Study 4
+E1 Randomization, Balance, and Missingness
 As in Study 2, technical issues caused small amounts of missing data.
@@
 | Pretest Writing Skill (mean (SD)) | 4.21 (0.88) | 4.23 (0.90) | 4.24 (0.90) | 4.17 (0.84) | .655 | 0.051 |
-F2 AI examples improve writing skill
+E2 AI examples improve writing skill
 The AI tool improved performance while participants used it.
@@
 | ||||||||||
-F3 Seeing AI examples was less effortful
+E3 Seeing AI examples was less effortful
 Table S30 shows OLS models predicting practice effort metrics from practice condition.
@@
 Table S31 shows OLS models predicting test effort metrics from practice condition.
-Results show some differences: participants who had seen the AI example write for longer duing the test, and pressed more keys, however their subjective experience of effort was not different from those who practice with or without the AI tool.
+Results show some differences: participants who had seen the AI example wrote for longer during the test, and pressed more keys; however, their subjective experience of effort was not different from those who practiced with or without the AI tool.
 | sqrt(Time) | log(Keystrokes) | Subjective Rating (0 - 10) | Time | Keystrokes | Keystrokes/min | |
@@
 | Note. *** p < .001, ** p < .01, * p < .05. | ||||||
-F4 Seeing an AI example did not discourage motivation for future learning
+E4 Seeing an AI example did not discourage motivation for future learning
 Table S33 presents differences in perceived learning, perceived writing skill, and the likelihood of asking for feedback across conditions, with effect sizes and means reported for each comparison.
@@
 | Note. *** p < .001, ** p < .01, * p < .05. | |||
-F5 The benefits of seeing an AI example were just as large a day later
+E5 The benefits of seeing an AI example were just as large a day later
 Table S34 shows means and standardized differences for measures of writing skill and related outcomes during the follow-up phase.
@@
 | Experience: I use AI writing assistant(s) every day | -0.063 | 0.131 | -0.480 | 0.631 |
-F6 Seeing AI examples was equally effective across subgroups
+E6 Seeing AI examples was equally effective across subgroups
 As in Study 2, we tested whether each of the pretreatment demographic variables moderated the effects of seeing an AI example.
-To do this, we ran separate linear in which writing skill during the test phase was regressed on condition, the pre-treatment moderator of interest, writing skill at baseline, and an interaction term between the moderator condition.
+To do this, we ran separate linear regressions in which writing skill during the test phase was regressed on condition, the pre-treatment moderator of interest, writing skill at baseline, and an interaction term between the moderator condition.
 After correcting the p-values for the interaction terms, none were significant at the .05 level, suggesting that seeing AI examples was equally effective across groups.
@@
 | ||||||||||||||||
-F7 Pairwise Comparisons
+E7 Pairwise Comparisons
 The relative likelihood of a cover letter receiving an invitation to an interview was correlated with the GPT-rated writing quality. See Figure S13.